### v2

# The solving degrees for computing Gröbner bases of affine semi-regular polynomial sequences

Abstract
Determining the complexity of computing Gröbner bases is an important problem both in theory and in practice, and for that the solving degree plays a key role. In this paper, we study the solving degrees of affine semi-regular sequences and their homogenized sequences. Some of our results are considered to give mathematically rigorous proofs of the correctness of methods for computing Gröbner bases of the ideal generated by an affine semi-regular sequence. This paper is a sequel of the authors’ previous work [30] and gives additional results on the solving degrees and important behaviors of Gröbner basis computation.
We also define the generalized degree of regularity for a sequence of homogeneous polynomials. For the homogenization of an affine semi-regular sequence, we relate its generalized degree of regularity with its maximal Gröbner basis degree (i.e., the solving degree of the homogenized sequence). The definition of a generalized (cryptographic) semi-regular sequence is also given, and it derives a new cryptographic assumption to estimate the security of cryptosystems and signature schemes. From our experimental observation, we raise a conjecture and some questions related to this generalized semi-regularity. These new definitions and our results provide a theoretical formulation of (somehow heuristic) discussions done so far in the cryptographic community.
1 Introduction
Let $K$ be a field, and let $\overline{K}$ denote its algebraic closure. We denote by $\mathbb{A}_{K}^{n}$ (resp. $\mathbb{P}_{K}^{n}$) the $n$-dimensional affine (resp. projective) space over $K$. Let $R=K[x_{1},\ldots,x_{n}]$ be the polynomial ring in $n$ variables over $K$. For a given monomial ordering $\prec$ on the set of monomials in $R$, let $\mathrm{LM}(f)$ denote the leading monomial of $f\in R\smallsetminus\{0\}$ with respect to it. For a non-empty subset $F\subset R\smallsetminus\{0\}$, put $\mathrm{LM}(F):=\{\mathrm{LM}(f):f\in F\}$. A set $F$ (resp. a sequence $\bm{F}$) of polynomials in $R$ is said to be homogeneous if the elements of $F$ (resp. $\bm{F}$) are all homogeneous, and otherwise $F$ is said to be affine. We denote by $\langle F\rangle_{R}$ (or $\langle F\rangle$ simply) the ideal generated by a non-empty subset $F$ of $R$. For a polynomial $f$ in $R\smallsetminus\{0\}$, let $f^{\rm top}$ denote its maximal total degree part which we call the top part of $f$, and let $f^{h}$ denote its homogenization in $R^{\prime}=R[y]$ by an extra variable $y$, see Subsection A.2 below for details. For a sequence $\bm{F}=(f_{1},\ldots,f_{m})\in(R\smallsetminus\{0\})^{m}$, we also set $\bm{F}^{\rm top}:=(f_{1}^{\rm top},\ldots,f_{m}^{\rm top})$ and $\bm{F}^{h}:=(f_{1}^{h},\ldots,f_{m}^{h})$. For a finitely generated graded $R$-(or $R^{\prime}$-)module $M$, we also denote by ${\rm HF}_{M}$ and ${\rm HS}_{M}$ its Hilbert function and its Hilbert–Poincaré series, respectively.
A Gröbner basis of an ideal $I$ in $R$ is defined as a special kind of generating set for $I$, and it gives a computational tool to determine many properties of $I$. A typical application of computing Gröbner bases is solving the multivariate polynomial (MP) problem: Given a sequence $\bm{F}=(f_{1},\ldots,f_{m})$ of $m$ polynomials $f_{1},\ldots,f_{m}$ in $R\smallsetminus\{0\}$, find $(a_{1},\ldots,a_{n})\in K^{n}$ such that $f_{i}(a_{1},\ldots,a_{n})=0$ for all $i$ with $1\leq i\leq m$. A particular case where $f_{1},\ldots,f_{m}$ are all quadratic is called the MQ problem, and its hardness is applied to constructing public-key cryptosystems and digital signature schemes that are expected to be quantum resistant. Therefore, analyzing the complexity of computing Gröbner bases is one of the most important problems both in theory and in practice.
An algorithm for computing Gröbner bases was proposed first by Buchberger [6], and so far a number of its improvements such as the $F_{4}$ [18] and $F_{5}$ [19] algorithms have been proposed. In determining the complexity of computing Gröbner bases, as we will see in the first paragraph of Subsection 2.2 below, one of the most important cases is the case where the input system is zero-dimensional and where the monomial ordering is graded (i.e., degree-compatible), and we focus on that case in the rest of this paper. Namely, we suppose that the input sequence $\bm{F}=(f_{1},\ldots,f_{m})$ admits a finite number of zeros in $\mathbb{A}^{n}_{\overline{K}}$ (resp. $\mathbb{P}^{n-1}_{\overline{K}}$) if $\bm{F}$ is affine (resp. homogeneous), and we consider a monomial ordering $\prec$ on $R$ that compares monomials first by their total degrees, e.g., a degree reverse lexicographical (DRL) ordering. Then, the complexity of the Gröbner basis computation for $F=\{f_{1},\ldots,f_{m}\}$ is estimated as a function of the solving degree(s): To the authors’ best knowledge, there are three (in fact four) kinds of definitions of solving degree, and they will be rigorously described in Subsection 2.2 below. In the first definition, the solving degree is defined as the highest degree of the polynomials involved during the Gröbner basis computation. Since this solving degree depends on an algorithm $\mathcal{A}$ that one adopts, we denote it by $\mathrm{sd}_{\prec}^{\mathcal{A}}(F)$. On the other hand, in the second and the third definitions, which were originally provided in a series of Gorla et al.’s studies (cf. [8], [4], [24], [9], [23]), we can see that the solving degrees do not depend on an algorithm, but only on $F$ and $\prec$. The solving degree in the second (resp. third) definition is defined by using Macaulay matrices (resp. those with mutants), and it is denoted by $\mathrm{sd}_{\prec}^{\rm mac}(F)$ (resp. $\mathrm{sd}_{\prec}^{\rm mut}(F)$) in this paper, where the subscripts “mac” and “mut” stand for Macaulay matrices and mutants respectively. Note that, when $F$ is homogeneous, these three solving degrees coincide with one another (for $\mathcal{A}$ with suitable setting) and we call them the solving degree simply; they are equal to the maximal Gröbner basis degree $\mathrm{max.GB.deg}_{\prec}(F)$ of $F$ with respect to $\prec$. In this case, we can apply a well-known bound [31, Theorem 2] by Lazard. In the following, we mainly treat with the case where $F$ is affine.
In their celebrated works (cf. [8], [4], [24], [9], [23]), Gorla et al. have studied well the relations between the solving degrees $\mathrm{sd}_{\prec}^{\rm mac}(F)$ and $\mathrm{sd}_{\prec}^{\rm mut}(F)$ and other invariants such as the degree of regularity and the Castelnuovo–Mumford regularity. Their results provide a mathematically rigorous framework for estimating the complexity of computing Gröbner bases. In particular, Caminata-Gorla [8] proved the following upper-bound on $\mathrm{sd}_{\prec}^{\rm mac}(F)$ by using Lazard’s bound:
  • ([8, Theorem 11]) When $K=\mathbb{F}_{q}$, the solving degree $\mathrm{sd}_{\prec}^{\rm mac}(F)$ for a DRL ordering $\prec$ can be upper-bounded by the Macaulay bound $d_{1}+\cdots+d_{\ell}-\ell+1$ with $d_{1}\geq d_{2}\geq\cdots\geq d_{m}$ and $\ell=\min\{n+1,\ell\}$, if $F$ contains the field equations $x_{i}^{q}-x_{i}$ for all $1\leq i\leq n$.
As for upper-bounds on the solving degrees $\mathrm{sd}_{\prec}^{\mathcal{A}}(F)$ and $\mathrm{sd}_{\prec}^{\rm mut}(F)$, we know the following:
  • Semaev-Tenti [40] (see also Tenti’s PhD thesis [41]) constructed a Buchberger-like algorithm $\mathcal{A}$ for the case $K=\mathbb{F}_{q}$ such that $\mathrm{sd}_{\prec}^{\mathcal{A}}(F)\leq 2D-2$ with $D:=d_{\rm reg}(\langle F^{\rm top}\rangle)$ for a DRL ordering $\prec$, assuming that $\{x_{i}^{q}-x_{i}:1\leq i\leq n\}\subset F$ and $\max\{q,\deg(f_{1}),\ldots,\deg(f_{m})\}\leq D$. Here $d_{\rm reg}(\langle F^{\rm top}\rangle)$ is the degree of regularity for $F^{\rm top}$, i.e., the smallest integer $d$ with $R_{d}=\langle F^{\rm top}\rangle_{d}$, where $R_{d}$ denotes the homogeneous part (component) of degree $d$ and where we set $I_{d}=I\cap R_{d}$ for a homogeneous ideal $I$ of $R$.
  • Caminata-Gorla proved in [9, Theorem 3.1] that $\mathrm{sd}_{\prec}^{\rm mut}(F)=\max\{d_{F},{\rm max.GB.deg}_{{\prec}}(F)\}$ for any graded monomial ordering $\prec$, where $d_{F}$ denotes the last fall degree of $F$ defined in [9, Definition 1.5] (originally in [27], [26]). Recently, Salizzoni [39] also proved $\mathrm{sd}_{\prec}^{\rm mut}(F)\leq D+1$, in the case where $\max\{\deg(f_{1}),\ldots,\deg(f_{m})\}\leq D<\infty$.
In this paper, by a mathematically rigorous way following Gorla et al.’s works, we study the solving degrees and related Gröbner bases of affine semi-regular polynomial sequences, where a sequence $\bm{F}=(f_{1},\ldots,f_{m})\in(R\smallsetminus K)^{m}$ of (not necessarily homogeneous) polynomials is said to be affine semi-regular (resp. affine cryptographic semi-regular) if $\bm{F}^{\rm top}=(f_{1}^{\rm top},\ldots,f_{m}^{\rm top})$ is semi-regular (resp. cryptographic semi-regular), see Definitions 2.1.3, 2.1.9, and 2.1.12 for details. Note that homogeneous semi-regular sequences are conjectured by Pardue [36, Conjecture B] to be generic sequences of polynomials (see e.g., [36] for the definition of genericness), and affine (cryptographic) semi-regular sequences are often appearing in the construction of multivariate public key cryptosystems and digital signature schemes. As a sequel of the authors’ previous work [30], we investigate further results on the solving degrees and on behaviors of the computation of Gröbner bases.
As the first main result in this paper, we revisit the result in our previous paper [30] with some additional remarks, which shall give an explicit characterization (Theorem 1 below) of the Hilbert function and the Hilbert-Poincaré series associated to the homogenization $F^{h}$. This characterization is useful to analyze the Gröbner basis computation for both $F$ and $F^{h}$.
Theorem 1 (Theorem 3.1.1, Remark 3.1.2, Remark 3.1.3 and Corollary 3.1.5).
With notation as above, assume that $\bm{F}$ is affine cryptographic semi-regular, and put $D:=d_{\rm reg}(\langle F^{\rm top}\rangle)$. Then, we have the following:
  1. (1)
    For each $d$ with $d<D$, we have ${\rm HF}_{R^{\prime}/\langle F^{h}\rangle}(d)={\rm HF}_{R/\langle F^{\rm top}\rangle}(d)+{\rm HF}_{{R^{\prime}}/\langle F^{h}\rangle}(d-1)$, and hence ${\rm HF}_{R^{\prime}/\langle F^{h}\rangle}(d)=\sum_{i=0}^{d}{\rm HF}_{R/\langle F^{\rm top}\rangle}(i)$.
  2. (2)
    The Hilbert function ${\rm HF}_{R^{\prime}/\langle F^{h}\rangle}$ is unimodal and its highest value is attained at $d=D-1$. In more detail, the multiplication map by $y$ from $(R^{\prime}/\langle F^{h}\rangle)_{d-1}$ to $(R^{\prime}/\langle F^{h}\rangle)_{d}$ is injective for $d<D$ and surjective for $d\geq D$.
  3. (3)
    There exists $d_{0}$ such that ${\rm HF}_{R^{\prime}/\langle F^{h}\rangle}(d_{0})={\rm HF}_{R^{\prime}/\langle F^{h}\rangle}(d)$ for all $d$ with $d\geq d_{0}$, namely the number of projective zeros for $F^{h}$ is finite at most. Here, $d_{0}$ gives an upper-bound on the solving degree of $F^{h}$ (or equivalently the maximal degree of the Gröbner basis of $\langle F^{h}\rangle$).
  4. (4)
    ${\rm HS}_{R^{\prime}/\langle F^{h}\rangle}(z)\equiv\prod_{i=1}^{m}(1-z^{d_{i}})/(1-z)^{n+1}\pmod{z^{D}}$, so that $F^{h}$ is $D$-regular, equivalently $\mathrm{syz}(F^{h})_{<D}=\mathrm{tsyz}(F^{h})_{<D}$. Here we denote by $\mathrm{syz}(F^{h})$ and $\mathrm{tsyz}(F^{h})$ the module of syzygies of $F^{h}$ and that of trivial syzygies of $F^{h}$, respectively (see Appendix A.1 for the definition of $\mathrm{syz}(F^{h})$ and $\mathrm{tsyz}(F^{h})$).
As for (3) of Theorem 1, it follows from the proof of Lazard’s bound [31, Theorem 2] that $\mathrm{max.GB.deg}_{\prec^{h}}(F^{h})\leq d_{0}$ for a DRL ordering $\prec$ (we give an explicit proof for this in Lemma 2.2.2 below), where $\prec^{h}$ is the homogenization of $\prec$. As in [8, Theorem 11] recalled above, we can apply Lazard’s bound to obtaining $d_{0}\leq d_{1}+\cdots+d_{\ell}-\ell+1$ with $d_{i}=\deg(f_{i})$ and $\ell=\min\{n+1,\ell\}$, assuming $d_{1}\geq\cdots\geq d_{m}$ in descending order. As an additional result in this paper, we also obtain the following upper-bound on the solving degree of $F^{h}$:
Theorem 2 (Theorem 3.2.3 and Proposition 3.2.5).
  1. (1)
    Suppose that that $d_{1}\leq d_{2}\leq\cdots\leq d_{m}$ (in ascending order) and $m>n$. If $\bm{F}^{\rm top}$ satisfies a stronger condition that it is semi-regular, then the solving degree of $F^{h}$ is upper-bounded by $d_{1}+d_{2}+\cdots+d_{n}+d_{m}-n$. Moreover, if $d_{m}\leq D$, the solving degree of $F^{h}$ is upper-bounded by $d_{1}+d_{2}+\cdots+d_{n}+d_{n+1}-n$.
  2. (2)
    Let $S_{0}$ be the saturation exponent of $(\langle F^{h}\rangle:\langle y^{\infty}\rangle)$, that is, the minimum integer $s$ such that $(\langle F^{h}\rangle:\langle y^{s}\rangle)=(\langle F^{h}\rangle:\langle y^{\infty}\rangle)$. Then the solving degree of $F^{h}$ is upper-bounded by $D+s_{0}$.
Based on Theorem 1, we can explore the computations of reduced Gröbner bases of $\langle F\rangle$, $\langle F^{h}\rangle$, and $\langle F^{\rm top}\rangle$ in Section 4 below, dividing the cases into the degree less than $D$ or not. More precisely, denoting by $G$, $G_{\rm hom}$, and $G_{\rm top}$ the reduced Gröbner bases of $\langle F\rangle$, $\langle F^{h}\rangle$, and $\langle F^{\rm top}\rangle$ respectively, where their monomial orderings are DRL $\prec$ or its extension $\prec^{h}$, we revisit [30, Section 5] and obtain more precise results:
Theorem 3 (Section 4; cf. [30, Section 5]).
With notation as above, assume that $\bm{F}$ is affine cryptographic semi-regular, and that $D:=d_{\rm reg}(\langle F^{\rm top}\rangle)<\infty$.
  1. (1)
    ${\rm LM}(G_{\rm hom})_{d}={\rm LM}(G_{\rm top})_{d}$ for each degree $d<D$. This implies that the Gröbner basis computation process for $\langle F^{h}\rangle$ corresponds to that for $\langle F\rangle$, for each degree less than $D$.
  2. (2)
    $\langle{\rm LM}((G_{\rm hom})_{\leq D})\rangle_{R[y]}\cap R_{D}=R_{D}$. Moreover, for each element $g$ in $(G_{\rm hom})_{D}$ with $g^{\rm top}:=g(x_{1},\ldots,x_{n},0)\not=0$, the top-part $g^{\rm top}$ consists of one term, that is, $g^{\rm top}={\rm LT}(g)$.
  3. (3)
    There is a strong correspondence between the computation of $G_{\rm hom}$ and that of $G$ at early stages, namely, at the step degrees not greater than $D$.
  4. (4)
    If $D\geq{\rm max}\{\deg(f):f\in F\}$, then the maximal Gröbner basis degree with respect to a DRL ordering $\prec$ is upper-bounded by $D$. Moreover, there exists a Buchberger-like algorithm $\mathcal{A}$ whose solving degree $\mathrm{sd}_{\prec}^{\mathcal{A}}(F)$ is upper-bounded by $12D-$ 1, and by $22D-$ 2 in the strict sense (see (I) in Subsection 2.2 for details on the definition of the terminology ‘strict sense’).
Note that (2) and the first half of (4) hold not necessarily assuming the affine cryptographic semi-regularity of $\bm{F}$.
In particular, we rigorously prove some existing results, which are often used for analyzing the complexity of computing Gröbner bases, and moreover extend them to our case.
Furthermore, based on Lemma 2.2.2 below, for homogeneous ideals of projective dimension 0, we naturally extend the notion of degree of regularity. The new notion of generalized degree of regularity plays a very important role on the computation of Gröbner bases for such ideals, see Subsections 2.3 and 4.3. The following proposition summarizes theoretical results on this generalized degree of regularity:
Proposition 1 (Proposition 2.3.3).
With notation as above, assume that $R/\langle F^{\rm top}\rangle$ is Artinian, and that $\bm{F}^{\rm top}$ is cryptographic semi-regular. Then we have the following:
  1. 1.
    $\widetilde{d}_{\rm reg}(\langle F^{h}\rangle)\geq d_{\rm reg}(\langle F^{\rm top}\rangle)-1$.
  2. 2.
    $\mathrm{max.GB.deg}_{\prec^{h}}(F^{h})\leq\max\{{d}_{\rm reg}(\langle F^{\rm top}\rangle),\widetilde{d}_{\rm reg}(\langle F^{h}\rangle)\}$ for any graded monomial order $\prec$.
Moreover, the equality holds in the second inequality if $\langle\mathrm{LM}(\langle F^{h}\rangle)\rangle$ is weakly reverse lexicographic.
Here, a weakly reverse lexicographic ideal is a monomial ideal $J$ such that if $x^{\alpha}$ is one of the minimal generators of $J$ then every monomial of the same degree which preceeds $x^{\alpha}$ must belong to $J$ as well (see [36, Section 4] for the original definition).
Finally, in Subsection 4.3, we give some observation on the behavior of Gröbner basis computation based on our experiments, from which we arrive at a conjecture on the Hilbert–Poincaré series for affine polynomial sequences without constant terms. For this conjecture, we generalize the notion of cryptographic semi-regular in Subsection 2.3.
Notation
  • $R=K[x_{1},\ldots,x_{n}]$: The polynomial ring of $n$ variables over a field $K$.
  • $\mathrm{deg}(f)$: The total degree of $f\in R$.
  • $f^{\rm top}$: The maximal total degree part of $f\in R$, namely, $f^{\rm top}$ is the sum of all terms of $f$ whose total degree equals to $\deg(f)$.
  • $f^{h}$: The homogenization of $f\in R\smallsetminus\{0\}$ by an extra variable $y$, say $f^{h}:=y^{\deg(f)}f(x_{1}/y,\ldots,x_{n}/y)$.
  • ${\rm HF}_{M}$: The Hilbert function of a finitely generated graded $R$-module $M=\bigoplus_{d\in\mathbb{Z}}M_{d}$, say ${\rm HF}_{M}(d)=\mathrm{dim}_{K}M_{d}$ for each $d\in{\color[rgb]{0,0,0}\mathbb{Z}_{\geq 0}}$.
  • ${\rm HS}_{M}$: The Hilbert–Poincaré series of a finitely generated graded $R$-module $M=\bigoplus_{d\in\mathbb{Z}}M_{d}$, say ${\rm HS}_{M}(z)=\sum_{d=0}^{\infty}{\rm HF}_{M}(d)z^{d}\in\mathbb{Z}\llbracket z\rrbracket$.
  • $K_{\bullet}(f_{1},\ldots,f_{m})$: The Koszul complex on a sequence $(f_{1},\ldots,f_{m})$ of homogeneous polynomials in $R$.
  • $H_{i}(K_{\bullet}(f_{1},\ldots,f_{m}))$: The $i$-th homology group of the Koszul complex $K_{\bullet}(f_{1},\ldots,f_{m})$.
As for the definition of Koszul complex and homogenization, see Appendix A for details.
2 Preliminaries
In this section, we recall definitions of semi-regular sequences and solving degrees, and collect some known facts related to them.
2.1 Semi-regular sequences
We first review the notion of semi-regular sequence defined by Pardue [36].
Definition 2.1.1 (Semi-regular sequences, [36, Definition 1]).
Let $I$ be a homogeneous ideal of $R$. A degree-$d$ homogeneous element $f\in R$ is said to be semi-regular on $I$ if the multiplication map $(R/I)_{t-d}\longrightarrow(R/I)_{d}\ ;\ g\longmapsto gf$ is injective or surjective, for every $t$ with $t\geq d$. A sequence $(f_{1},\ldots,f_{m})\in(R\smallsetminus\{0\})^{m}$ of homogeneous polynomials is said to be semi-regular on $I$ if $f_{i}$ is semi-regular on $I+\langle f_{1},\ldots,f_{i-1}\rangle_{R}$, for every $i$ with $1\leq i\leq m$.
Throughout the rest of this subsection, let $f_{1},\ldots,f_{m}\in R\smallsetminus K$ be homogeneous elements of degree $d_{1},\ldots,d_{m}$ respectively, unless otherwise noted, and put $I=\langle f_{1},\ldots,f_{m}\rangle_{R}$, $I^{(0)}:=\{0\}$, and $A^{(0)}:=R/I^{(0)}=R$. For each $i$ with $1\leq i\leq m$, we also set $I^{(i)}:=\langle f_{1},\ldots,f_{i}\rangle_{R}$ and $A^{(i)}:=R/I^{(i)}$. The degree-$d$ homogeneous part $A_{d}^{(i)}$ of each $A^{(i)}$ is given by $A_{d}^{(i)}=R_{d}/I_{d}^{(i)}$, where $I_{d}^{(i)}=I^{(i)}\cap R_{d}$. We denote by $\psi_{f_{i}}$ the multiplication map
$A^{(i-1)}\ni g\longmapsto gf_{i}\in A^{(i-1)},$
which is a graded homomorphism of degree $d_{i}$. For every $t$ with $t\geq d_{i}$, the restriction map
$\psi_{f_{i}}|_{A^{(i-1)}_{t-d_{i}}}:A^{(i-1)}_{t-d_{i}}\longrightarrow A^{(i-1)}_{t}$
is a $K$-linear map.
The semi-regularity is characterized by equivalent conditions in Proposition 2.1.2 below. In particular, the fourth condition enables us to compute the Hilbert–Poincaré series of each $A^{(i)}$.
Proposition 2.1.2 (cf. [36, Proposition 1]).
With notation as above, the following are equivalent:
  1. 1.
    The sequence $(f_{1},\ldots,f_{m})$ is semi-regular.
  2. 2.
    For each $1\leq i\leq m$ and for each $t\geq d_{i}$, the multiplication map $\psi_{f_{i}}|_{A^{(i-1)}_{t-d_{i}}}$ is injective or surjective, namely $\dim_{K}A^{(i)}_{t}=\max\{0,\dim_{K}A^{(i-1)}_{t}-\dim_{K}A^{(i-1)}_{t-d_{i}}\}$.
  3. 3.
    For each $i$ with $1\leq i\leq m$, we have ${\rm HS}_{A^{(i)}}(z)=[{\rm HS}_{A^{(i-1)}}(z)(1-z^{d_{i}})]$, where []delimited-[][$\cdot$][ ⋅ ] means truncating a formal power series over $\mathbb{Z}$ after the last consecutive positive coefficient.
  4. 4.
    For each $i$ with $1\leq i\leq m$, we have ${\rm HS}_{A^{(i)}}(z)=\left[\frac{\prod_{j=1}^{i}(1-z^{d_{j}})}{(1-z)^{n}}\right]$.
When $K$ is an infinite field, Pardue also conjectured in [36, Conjecture B] that generic polynomial sequences are semi-regular.
We next review the notion of cryptographic semi-regular sequence, which is defined by a condition weaker than one for semi-regular sequence. The notion of cryptographic semi-regular sequence is introduced first by Bardet et al. (e.g., [2], [3]) motivated to analyze the complexity of computing Gröbner bases. Diem [14] also formulated cryptographic semi-regular sequences, in terms of commutative and homological algebra. The terminology “cryptographic” was named by Bigdeli et al. in their recent work [4], in order to distinguish such a sequence from a semi-regular one defined by Pardue (see Definition 2.1.1).
Definition 2.1.3 ([2, Definition 3]; see also [14, Definition 1]).
Let $f_{1},\ldots,f_{m}\in R$ be homogeneous polynomials of positive degrees $d_{1},\ldots,d_{m}$ respectively, and put $I=\langle f_{1},\ldots,f_{m}\rangle_{R}$. For each integer $d$ with $d\geq\mathrm{max}\{d_{i}:1\leq i\leq m\}$, we say that a sequence $(f_{1},\ldots,f_{m})$ is $d$-regular if it satisfies the following condition:
  • For each $i$ with $1\leq i\leq m$, if a homogeneous polynomial $g\in R$ satisfies $gf_{i}\in\langle f_{1},\ldots,f_{i-1}\rangle_{R}$ and $\mathrm{deg}(gf_{i})<d$, then we have $g\in\langle f_{1},\ldots,f_{i-1}\rangle_{R}$. In other word, the multiplication map $A^{(i-1)}_{t-d_{i}}\longrightarrow A^{(i-1)}_{t}\ ;\ g\mapsto gf_{i}$ is injective for every $t$ with $d_{i}\leq t<d$.
Diem [14] determined the (truncated) Hilbert series of $d$-regular sequences as in the following proposition:
Theorem 2.1.4 (cf. [14, Theorem 1]).
With the same notation as in Definition 2.1.3, the following are equivalent for each $d$ with $d\geq\mathrm{max}\{d_{i}:1\leq i\leq m\}$:
  1. 1.
    The sequence $(f_{1},\ldots,f_{m})$ is $d$-regular. Namely, for each $(i,t)$ with $1\leq i\leq m$ and $d_{i}\leq t<d$, the equality $\dim_{K}A^{(i)}_{t}=\dim_{K}A^{(i-1)}_{t}-\dim_{K}A^{(i-1)}_{t-d_{i}}$ holds.
  2. 2.
    We have
    ${\rm HS}_{A^{(m)}}(z)\equiv\frac{\prod_{j=1}^{m}(1-z^{d_{j}})}{(1-z)^{n}}\pmod{z^{d}}.$ (2.1.1)
  3. 3.
    $H_{1}(K_{\bullet}(f_{1},\ldots,f_{m}))_{\leq d-1}=0$.
Proposition 2.1.5 ([14, Proposition 2 (a)]).
With the same notation as in Definition 2.1.3, let $D$ and $i$ be natural numbers. Assume that $H_{i}(K(f_{1},\ldots,f_{m}))_{\leq D}=0$. Then, for each $j$ with $1\leq j<m$, we have $H_{i}(K(f_{1},\ldots,f_{j}))_{\leq D}=0$.
Definition 2.1.6.
A finitely generated graded $R$-module $M$ is said to be Artinian if there exists a sufficiently large $D\in\mathbb{Z}$ such that $M_{d}=0$ for all $d\geq D$.
Definition 2.1.7 ([2, Definition 4], [3, Definition 4]).
For a homogeneous ideal $I$ of $R$, we define its degree of regularity $d_{\rm reg}(I)$ as follows: If the finitely generated graded $R$-module $R/I$ is Artinian, we set $d_{\rm reg}(I):=\mathrm{min}\{d:R_{d}=I_{d}\}$, and otherwise we set $d_{\rm reg}(I):=\infty$.
As for an upper-bound on the degree of regularity, we refer to [23, Theorem 21].
Remark 2.1.8.
In Definition 2.1.7, since $R/I$ is Noetherian, it is Artinian if and only if it is of finite length. In this case, the degree of regularity $d_{\rm reg}(I)$ is equal to the Castelnuovo-Mumford regularity $\mathrm{reg}(I)$ of $I$ (see e.g., [16, §20.5] for the definition), whence $d_{\rm reg}(I)=\mathrm{reg}(I)=\mathrm{reg}(R/I)+1$.
Definition 2.1.9 ([2, Definition 5], [3, Definition 5]; see also [14, Section 2]).
A sequence $(f_{1},\ldots,f_{m})\in(R\smallsetminus K)^{m}$ of homogeneous polynomials is said to be cryptographic semi-regular if it is $d_{\rm reg}(I)$-regular, where we set $I=\langle f_{1},\ldots,f_{m}\rangle_{R}$.
The cryptographic semi-regularity is characterized by equivalent conditions in Proposition 2.1.10 below.
Proposition 2.1.10 ([14, Proposition 1 (d)]; see also [3, Proposition 6]).
With the same notation as in Definition 2.1.3, we put $D=d_{\rm reg}(I)$. Then, the following are equivalent:
  1. 1.
    $(f_{1},\ldots,f_{m})\in(R\smallsetminus K)^{m}$ is cryptographic semi-regular.
  2. 2.
    We have
    ${\rm HS}_{R/I}(z)=\left[\frac{\prod_{j=1}^{m}(1-z^{d_{j}})}{(1-z)^{n}}\right].$ (2.1.2)
  3. 3.
    $H_{1}(K_{\bullet}(f_{1},\ldots,f_{m}))_{\leq D-1}=0$.
Remark 2.1.11.
By the definition of degree of regularity, if $(f_{1},\ldots,f_{m})$ is cryptographic semi-regular, then $d_{\rm reg}(I)$ coincides with $\deg({\rm HS}_{R/I}(z))+1$, where we set $I=\langle f_{1},\ldots,f_{m}\rangle_{R}$.
In 1985, Fröberg had already conjectured in [21] that, when $K$ is an infinite field, a generic sequence of homogeneous polynomials $f_{1},\ldots,f_{m}\in R$ of degrees $d_{1},\ldots,d_{m}$ generates an ideal $I$ with the Hilbert–Poincaré series of the form (2.1.2), namely $(f_{1},\ldots,f_{m})$ is cryptographic semi-regular. It can be proved (cf. [36]) that Fröberg’s conjecture is equivalent to Pardue’s one [36, Conjecture B]. We also note that Moreno-Socías conjecture [35] is stronger than the above two conjectures, see [36, Theorem 2] for a proof.
It follows from the fourth condition of Proposition 2.1.2 together with the second condition of Proposition 2.1.10 that the semi-regularity implies the cryptographic semi-regularity. Note that, when $m\leq n$, both ‘semi-regular’ and ‘cryptographic semi-regular’ are equivalent to ‘regular’.
Finally, we define an affine semi-regular sequence.
Definition 2.1.12 (Affine semi-regular sequences).
A sequence $\bm{F}=(f_{1},\ldots,f_{m})\in(R\smallsetminus K)^{m}$ of not necessarily homogeneous polynomials $f_{1},\ldots,f_{m}$ is said to be semi-regular (resp. cryptographic semi-regular) if $\bm{F}^{\rm top}=(f_{1}^{\rm top},\ldots,f_{m}^{\rm top})$ is semi-regular (resp. cryptographic semi-regular). In this case, we call $F$ an affine semi-regular (resp. affine cryptographic semi-regular) sequence.
2.2 Solving degrees of Gröbner basis computation
In general, determining precisely the complexity of computing a Gröbner basis is very difficult; in the worst-case, the complexity is doubly exponential in the number of variables, see e.g., [10], [33], [37] for surveys. However, it is experimentally well-known that a Gröbner basis with respect to a graded monomial ordering, in particular degree reverse lexicographical (DRL) ordering, can be computed quite more efficiently than ones with respect to other orderings in general. Moreover, in the case where the input set $F=\{f_{1},\ldots,f_{m}\}$ of polynomials generate a zero-dimensional inhomogeneous ideal, once a Gröbner basis $G$ with respect to an efficient monomial ordering $\prec$ is computed, a Gröbner basis $G^{\prime}$ with respect to any other ordering $\prec^{\prime}$ can be computed easily by the FGLM basis conversion [20]. Even when $F$ is homogeneous, one can efficiently convert $G$ to $G^{\prime}$ by Gröbner walk [12] (or Hilbert driven [42] if both $\prec$ and $\prec^{\prime}$ are graded). From this, we focus on the case where the monomial ordering is graded, and if necessary we also assume that the ideal generated by the input polynomials is zero-dimensional.
Definitions of solving degrees
In the case where the chosen monomial ordering is graded, the complexity of computing a Gröbner basis is often estimated with the so-called solving degree. To the best of the authors’ knowledge, there are three (in fact four) kinds of definitions of solving degree, and we here review them. The first definition is explicitly provided first by Ding and Schmidt in [15], and it depends on algorithms or their implementations:
  • (I)
    As the first definition, we define the solving degree of an algorithm to compute a Gröbner basis as the highest degree of the polynomials involved during the execution of the algorithm, see [15, p. 36]. For example, applying Buchberger’s algorithm or its variants such as $F_{4}$ with the normal strategy, we collect critical S-pairs with the lowest degree and then reduce the corresponding S-polynomials in each iteration of the main loop of reductions. The lowest degree of each iteration is called the step degree. Then the solving degree is defined as the highest step degree. Instead, we may adopt the highest degree of $S$-polynomials appearing in the whole computation as in [41] and [40] by Semaeve-Tenti, and in this case we use the terminology “the solving degree in the strict sense”.
  • (I)’
    Their is a variant of the above first definition, where the solving degree is defined as a value depending not only on an algorithm but also on its implementation. More precisely, in [15, Section 2.1], the authors use the term solving degree for the step degree at which it takes the most amount of time among all iterations. In the cryptographic literature, the term solving degree often means this solving degree. Although this solving degree is estimated based on experiments, it is practically a quite important ingredient for analyzing the security of multivariate cryptosystems. The degree of regularity $d_{\rm reg}(\langle F^{\rm top}\rangle)$ can be often used as a proxy for this solving degree.
We do not consider the solving degree in (I)’, since this paper focuses on theoretical aspects on computing Gröbner bases, but not on aspects in practical implementation. For a graded monomial ordering $\prec$ on $R$ and an input set $F$ of non-zero polynomials in $R$, we denote by $\mathrm{sd}_{\prec}^{\mathcal{A}}(F)$ the solving degree in (I) of an algorithm $\mathcal{A}$ to compute a Gröbner basis of $F$ with respect to $\prec$.
On the other hand, Caminata and Gorla [8] defined the solving degree of an input system, so that it does not depend on an algorithm, by using Macaulay matrices. Here, a Macaulay matrix is defined as follows: For a (fixed) graded monomial ordering $\prec$ and a finite sequence $H=(h_{1},\ldots,h_{k})\in(R\smallsetminus\{0\})^{k}$ with $d:=\max\{\deg(h_{i}):1\leq i\leq k\}$, writing each $h_{i}$ as $h_{i}=\sum_{j=1}^{\ell}c_{i,j}t_{j}$, where $\mathcal{T}_{\leq d}=\{t_{1},\ldots,t_{\ell-1},t_{\ell}=1\}$ is the set of monomials in $R$ of degree $\leq d$ with $t_{1}\succ\cdots\succ t_{\ell}$, the Macaulay matrix of $H$, denoted by $\mathrm{Mac}_{\prec}(H)$ is defined to be the $k\times\ell$ matrix $(a_{i,j})_{i,j}$ over $K$ (we let $\mathrm{Mac}_{\prec}(H)$ be the $1\times 1$ zero-matrix if $H$ is empty). Moreover, for each non-negative integer $d$, the degree-$d$ Macaulay matrix of $F$, denoted by $M_{\leq d}(F)$ when $\prec$ is fixed, is defined as $M_{\leq d}(F):={\rm Mac}_{\prec}(\mathcal{S}_{\leq d}(F))$, where $\mathcal{S}_{\leq d}(F)$ is a sequence of the multiples $tf$ for $f\in F$ with $\deg(f)\leq d$ and $t\in\mathcal{T}_{\leq d-\deg(f)}$. Namely, the rows of $M_{\leq d}(F)$ correspond to $tf$’s above, and the columns are indexed by the monomials of degree at most $d$ in descending order with respect to $\prec$. Note that the order of elements in $\mathcal{S}_{\leq d}(F)$ can be arbitrary.
  • (II)
    We define the solving degree of $F$ with respect to a fixed (graded) monomial ordering as the lowest degree $d$ at which the reduced row echelon form (RREF) of $M_{\leq d}(F)$ produces a Gröbner basis of $F$.
Note that the computation of the RREF of $M_{\leq d}(F)$ corresponds to the standard XL algorithm [11], which is based on an idea of Lazard [31].
The third definition is given in Gorla et al.’s works (cf. [4], [24], [9], [23]), see also [39]. More precisely, for each non-negative integer $d\in\mathbb{Z}_{\geq 0}$, let $V_{F,d}$ be the smallest $K$-vector space such that $\{f\in F:\mathrm{deg}(f)\leq d\}\subset V_{F,d}$ and $\{tf:f\in V_{F,d},\ t\in\mathcal{T}_{\leq d-\mathrm{deg}(f)}\}\subset V_{F,d}$, where $\mathcal{T}_{\leq d}$ denotes the set of all monomials in $R$ of degree at most $d$. Then the third definition is as follows:
  • (III)
    The solving degree of $F$ is defined as the smallest $d$ for which $V_{F,d}$ contains a Gröbner basis of $F$ with respect to a fixed monomial ordering.
We can also describe the solving degree in (III) with Macaulay matrices. Specifically, we consider to compute a Gröbner basis of $F$ by the following mutant strategy:
  • Initialize $d$ as $d=\mathrm{max}\{\deg(f):f\in F\}$. Compute the RREF of $M_{\leq d}(F)$. If the RREF contains a polynomial $f$ with $\deg(f)<d$ whose leading monomial is not equal to that of any row of $M_{\leq d}(F)$, add to the RREF the new rows corresponding to $tf$ for all $t\in\mathcal{T}_{\leq d-\deg(f)}$ such that $tf$ does not belong to the linear space spanned by the rows of the RREF. Repeat the computation of the RREF and the operation of adding new rows, until there are no new rows to add. If the resulting matrix produces a Gröbner basis of $F$, then we stop, and otherwise we proceed to the next degree, $d+1$.
This strategy computes a basis of $V_{F,d}$ for each $d$, and therefore the smallest $d$ for which the mutant strategy terminates is equal to the solving degree of $F$ in terms of (III), see [24, Theorem 1]. As in [23], we refer to the algorithms such as Mutant-XL [7] and MXL2 [34] that employ this strategy as mutant algorithms. In the following, we denote the solving degree in (II) and that in (III) respectively by $\mathrm{sd}_{\prec}^{\rm mac}(F)$ and $\mathrm{sd}_{\prec}^{\rm mut}(F)$. By definitions, it is clear that $\mathrm{sd}_{\prec}^{\rm mut}(F)\leq\mathrm{sd}_{\prec}^{\rm mac}(F)$ for any graded monomial oredering $\prec$, and the equality holds if the elements in $F$ are all homogeneous.
In a series of their celebrated works (cf. [8], [4], [24], [9], [23]), Gorla et al. provided a mathematical formulation for the relations between the solving degrees $\mathrm{sd}_{\prec}^{\rm mac}(F)$ and $\mathrm{sd}_{\prec}^{\rm mut}(F)$ and algebraic invariants coming from $F$, such as the maximal Gröbner basis degree, the degree of regularity, the Castelnuovo–Mumford regularity, the first and last degrees, and so on. Here, the maximal Gröbner basis degree of the ideal $\langle F\rangle_{R}$ is the maximal degree of elements in the reduced Gröbner basis of $\langle F\rangle_{R}$ with respect to a fixed monomial ordering $\prec$, and is denoted by $\mathrm{max.GB.deg}_{\prec}(F)$. For any graded monomial oredering $\prec$, it is straightforward that
$\mathrm{max.GB.deg}_{\prec}(F)\leq\mathrm{sd}_{\prec}^{\rm mut}(F)\leq\mathrm{sd}_{\prec}^{\rm mac}(F).$ (2.2.1)
Upper bounds on solving degree
If $F$ consists of homogeneous elements, then one has $\mathrm{sd}_{\prec}^{\rm mac}(F)=\mathrm{sd}_{\prec}^{\rm mut}(F)$, and moreover these solving degrees are equal to $\mathrm{sd}_{\prec}^{\mathcal{A}}(F)$ if the algorithm $\mathcal{A}$ incrementally computes the reduced $d$-Gröbner basis for each $d$ in increasing the degree $d$. For example, Buchberger algorithm, $F_{4}$, $F_{5}$, matrix-$F_{5}$, and Hilbert driven algorithm are the cases. Furthermore, the equalities in (2.2.1) hold, and hence we can use a bound on $\mathrm{max.GB.deg}_{\prec}(F)$. Since we are now considering the zero-dimensional case, we can apply Lazard’s upper-bound below.
In the non-homogeneous case, i.e., $F$ contains at least one non-homogeneous element, the euqalities in (2.2.1) do not hold in general, and it is not so easy to estimate any of the solving degrees. A straightforward way of bounding the solving degrees in the non-homogeneous case is to apply the homogenization as follows. We set $\prec$ as the DRL ordering on $R$ with $x_{n}\prec\cdots\prec x_{1}$, and fix it throughout the rest of this subsection. Let $y$ be an extra variable for homogenization as in Subsection A.2, and $\prec^{h}$ the homogenization of $\prec$, so that $y\prec x_{i}$ for any $i$ with $1\leq i\leq n$. Then, we have
${\rm max.GB.deg}_{\prec}(F)\leq\mathrm{\rm sd}^{\rm mac}_{\prec}(F)=\mathrm{\rm sd}^{\rm mac}_{\prec^{h}}(F^{h})={\rm max.GB.deg}_{\prec^{h}}(F^{h}),$
see [8] for a proof. Here, we also recall Lazard’s bound for the maximal Gröbner basis degree of $\langle F^{h}\rangle_{R^{\prime}}$ with $R^{\prime}=R[y]$:
Theorem 2.2.1 (Lazard; [31, Theorem 2], [32, Théorèm 3.3]).
With notation as above, we assume that the number of projective zeros of $F^{h}$ is finite (and therefore $m\geq n$), and that $f_{1}^{h}=\cdots=f_{m}^{h}=0$ has no non-trivial solution over the algebraic closure $\overline{K}$ with $y=0$, i.e., $F^{\rm top}$ has no solution in $\overline{K}^{n}$ other than $(0,\ldots,0)$. Then, supposing also that $d_{1}\geq\cdots\geq d_{m}$, we have
$\mathrm{max.GB.deg}_{\prec^{h}}(F^{h})\leq d_{1}+\cdots+d_{\ell}-\ell+1$ (2.2.2)
with $\ell:=\mathrm{min}\{m,n+1\}$.
One of the most essential parts for the proof of Theorem 2.2.1 is an argument stated in the following lemma (we here write a proof for readers’ convenience):
Lemma 2.2.2.
With notation as above, let $d_{0}$ be a positive integer satisfying the following two properties:
  1. 1.
    The multiplication-by-$y$ map $(R^{\prime}/\langle F^{h}\rangle)_{d_{0}-1}\longrightarrow(R^{\prime}/\langle F^{h}\rangle)_{d_{0}}$ is surjective.
  2. 2.
    For any $d\in\mathbb{Z}$ with $d\geq d_{0}$, the multiplication-by-$y$ map $(R^{\prime}/\langle F^{h}\rangle)_{d}\longrightarrow(R^{\prime}/\langle F^{h}\rangle)_{d+1}$ is injective.
Then we have $\mathrm{max.GB.deg}_{\prec^{h}}(F^{h})\leq d_{0}$.
Proof.
Let $G$ be a Gröbner basis of $\langle F^{h}\rangle$ with respect to $\prec^{h}$. Clearly, we may suppose that each element of $G$ is homogeneous. It suffices to prove that $G_{\leq d_{0}}:=\{g\in G:\mathrm{deg}(g)\leq d_{0}\}$ is a Gröbner basis of $\langle F^{h}\rangle$ with respect to $\prec^{h}$. Indeed, the maximal degree of the reduced Gröbner basis of $\langle F^{h}\rangle$ with respect to $\prec^{h}$ is not greater than that of any Gröbner basis of $\langle F^{h}\rangle$ with respect to $\prec^{h}$.
Let $f\in\langle F^{h}\rangle$, and $d:=\mathrm{deg}(f)$. We show that there exists $g\in G_{\leq d_{0}}$ with $\mathrm{LM}(g)\mid\mathrm{LM}(f)$, by the induction on $d$. It suffices to consider the case where $f$ is homogeneous, since $\langle F^{h}\rangle$ is homogeneous. The case where $d\leq d_{0}$ is clear, and so we assume $d>d_{0}$.
First, if $\mathrm{LM}(f)\in R=K[x_{1},\ldots,x_{n}]$ (namely $y\nmid\mathrm{LM}(f)$), we choose an arbitrary monomial $t\in R$ of degree $d_{0}$ with $t\mid\mathrm{LM}(f)$. Since the multiplication map $(R^{\prime}/\langle F^{h}\rangle)_{d_{0}-1}\longrightarrow(R^{\prime}/\langle F^{h}\rangle)_{d_{0}}$ by $y$ is surjective, there exists a homogeneous polynomial $h\in(R^{\prime})_{d_{0}-1}$ such that $h_{1}:=t-yh\in\langle F^{h}\rangle$. Here, $h_{1}$ is homogeneous of degree $d_{0}$, and $y\nmid t$, whence $\mathrm{LM}(h_{1})=t$. Therefore, we have $\mathrm{LT}(g)\mid t$ for some $g\in G$. Since $\mathrm{deg}(t)=d_{0}$, we also obtain $\mathrm{deg}(g)\leq d_{0}$, so that $g\in G_{\leq d_{0}}$.
Next, assume that $y\mid\mathrm{LM}(f)$. In this case, it follows from the definition of $\prec^{h}$ that any other term in $f$ is also divisible by $y$, so that $f\in\langle y\rangle$. Hence, we can write $f=yf_{1}$ for some homogeneous $f_{1}\in R^{\prime}$. By $d-1\geq d_{0}$, the multiplication map $(R^{\prime}/\langle F^{h}\rangle)_{d-1}\longrightarrow(R^{\prime}/\langle F^{h}\rangle)_{d}$ by $y$ is injective, so that $f_{1}\in\langle F^{h}\rangle_{d-1}$. By the induction hypothesis, there exists $g\in G_{\leq d_{0}}$ such that $\mathrm{LM}(g)\mid\mathrm{LM}(f_{1})$. Since $\mathrm{LM}(f)=y\mathrm{LM}(f_{1})$, we obtain $\mathrm{LM}(g)\mid\mathrm{LM}(f)$. We have proved that $G_{\leq d_{0}}$ is a Gröbner basis of $\langle F^{h}\rangle$ with respect to $\prec^{h}$. ∎
Lazard proved that we can take $d_{1}+\cdots+d_{\ell}-\ell+1$ in Theorem 2.2.1 as $d_{0}$ in Lemma 2.2.2. Lazard’s bound given in (2.2.2) is also referred to as the Macaulay bound, and it provides an upper-bound for the solving degree of $F$ with respect to a DRL ordering.
As for the maximal Gröbner basis degree of $\langle F\rangle$, if $\langle F^{\rm top}\rangle$ is Aritinian, we have
$\mathrm{max.GB.deg}_{\prec^{\prime}}(F)\leq d_{\rm reg}(\langle F^{\rm top}\rangle)$ (2.2.3)
for any graded monomial ordering $\prec^{\prime}$ on $R$, see [8, Remark 15] or Lemma 4.2.4 below for a proof. Both $d_{\rm reg}(\langle F^{\rm top}\rangle)$ and $\mathrm{\rm sd}_{\prec}^{\rm mac}(F)$ are greater than or equal to ${\rm max.GB.deg}_{\prec}(F)$, whereas it is pointed out in [4], [8], and [9] by explicit examples that any of the degree of regularity and the first fall degree does not produce an estimate for the solving degrees in general, even when $F$ is an affine (cryptographic) semi-regular sequence. Caminata-Gorla proved in [9] that the solving degree $\mathrm{sd}_{\prec}^{\rm mut}(F)$ is nothing but the last fall degree if it is greater than the maximal Gröbner basis degree:
Theorem 2.2.3 ([9, Theorem 3.1]).
With notation as above, for any degree-compatible monomial ordering $\prec^{\prime}$ on $R$, we have the following equality:
$\mathrm{sd}_{\prec^{\prime}}^{\rm mut}(F)=\max\{d_{F},{\rm max.GB.deg}_{{\color[rgb]{0,0,0}\prec^{\prime}}}(F)\},$
where $d_{F}$ denotes the last fall degree of $F$ defined in [9, Definition 1.5] (originally in [27], [26]).
By this theorem, if $d_{\rm reg}({\langle F^{\rm top}\rangle})<d_{F}$, the degree of regularity is no longer an upper-bound on the solving degrees $\mathrm{sd}_{\prec}^{\rm mac}(F)$ and $\mathrm{sd}_{\prec}^{\rm mut}(F)$. Recently, Salizzoni [39] proved the following theorem:
Theorem 2.2.4 ([39, Theorem 1.1]).
With notation as above, we also set $D=d_{\rm reg}(\langle F^{\rm top}\rangle)$, and assume that $D\geq\max\{\deg(f):f\in F\}$. Then, for any graded monomial ordering $\prec^{\prime}$ on $R$, we have $\mathrm{sd}_{\prec^{\prime}}^{\rm mut}(F)\leq D+1$. Moreover, a Gröbner basis of $F$ can be find in $O((n+1)^{4(d+1)})$ operations in $K$.
On the other hand, Semaev and Tenti proved that the solving degree $\mathrm{sd}_{\prec}^{\mathcal{A}}(F)$ for some algorithm $\mathcal{A}$ is linear in the degree of regularity, if $K$ is a (small) finite field, and if the input system contains polynomials related to the field equations, say $x_{i}^{q}-x_{i}$ for $1\leq i\leq n$:
Theorem 2.2.5 ([40, Theorem 2.1], [41, Theorem 3.65 & Corollary 3.67]).
With notation as above, assume that $K=\mathbb{F}_{q}$, and that $F$ contains $x_{i}^{q}-x_{i}$ for all $1\leq i\leq n$. If $D\geq\max\{\deg(f):f\in F\}$ and $D\geq q$, then there exists a Buchberger-like algorithm $\mathcal{A}$ to compute the reduced Gröbner basis of $F$ with S-polynomials such that
$\mathrm{sd}^{\mathcal{A}}_{\prec}(F)\leq 2D-1.$ (2.2.4)
and
$\mathrm{sd}^{\mathcal{A}}_{\prec}(F)\leq 2D-2.$ (2.2.5)
in the strict sense (see the definition (I) of the solving degree for details). Furthermore, the complexity of the algorithm $\mathcal{A}$ is
$O(L_{q}(n,D)^{2}L_{q}(n,D-1)^{2}L_{q}(n,2D-2))$
operations in $K$, where $L_{q}(n,d)$ denotes the number of monomials in $\mathbb{F}_{q}[x_{1},\ldots,x_{n}]/\langle x_{1}^{q},\ldots,x_{n}^{q}\rangle$ of degree $\leq d$.
In Subsection 4.2 below, we will prove the same inequality as in (2.2.4), in the case where $F$ not necessarily contains a field equation but is cryptographic semi-regular.
2.3 Extension of the notion of degree of regularity
In this subsection, we shall extend the notion of degree of regularity of a homogeneous ideal $I$ of $R$. Let $\prec$ be a graded order on the monomials of $R$. Recall from Definition 2.1.7 that we set $d_{\rm reg}(I)=\infty$ if $R/I$ is not Artinian, but this is not feasible to analyzing the Gröbner basis computation of $F^{h}$ unless a system defined by $F^{h}$ has no non-trivial root over an algebraic closure. For the feasibility, we here give an alternative definition for degree of regularity, which is viewed as a generalization of the original definition (Definition 2.1.7):
Definition 2.3.1.
For a homogeneous ideal $I$ of $R$, we define its generalized degree of regularity $\widetilde{d}_{\rm reg}(I)$ as follows: If there exists an integer $d_{0}$ such that $\dim_{K}(R/I)_{d}=\dim_{K}(R/I)_{d_{0}}$ for any $d$ with $d\geq d_{0}$, we set $\widetilde{d}_{\rm reg}(I):=\mathrm{min}\{d_{0}:\dim_{K}(R/I)_{d}=\dim_{K}(R/I)_{d_{0}}\mbox{ for any $d$ with $d\geq d_{0}$}\}$, and otherwise we set $\widetilde{d}_{\rm reg}(I):=\infty$.
The existence of $d_{0}$ in Definition 2.3.1 is equivalent to that the number of projective zeros of $I$ is finite at most. Note also that, in Definition 2.3.1, we have $\widetilde{d}_{\rm reg}(I)={d}_{\rm reg}(I)<\infty$ if $R/I$ is Artinian, and otherwise $\widetilde{d}_{\rm reg}(I)<d_{\rm reg}(I)$ or $\widetilde{d}_{\rm reg}(I)={d}_{\rm reg}(I)=\infty$. We also extend the cryptographic semi-regularity (Definition 2.1.9) of a sequence of homogeneous polynomials, as follows:
Definition 2.3.2.
A sequence $(f_{1},\ldots,f_{m})\in(R\smallsetminus K)^{m}$ of homogeneous polynomials is said to be generalized cryptographic semi-regular if it is $\widetilde{d}_{\rm reg}(I)$-regular, where we set $I=\langle f_{1},\ldots,f_{m}\rangle_{R}$.
A sequence $\bm{F}=(f_{1},\ldots,f_{m})\in(R\smallsetminus K)^{m}$ of not necessarily homogeneous polynomials $f_{1},\ldots,f_{m}$ is said to be generalized cryptographic semi-regular if $\bm{F}^{h}=(f_{1}^{h},\ldots,f_{m}^{h})$ is generalized cryptographic semi-regular. In this case, we call $\bm{F}$ an affine generalized cryptographic semi-regular sequence.
Here, we relate the solving degree of $F^{h}$ (namely the maximal Gröbner basis degree of $F^{h}$) with our generalized degree of regularity, under some assumptions.
Proposition 2.3.3.
Let $\bm{F}=(f_{1},\ldots,f_{m})\in(R\smallsetminus K)^{m}$ be a sequence of not necessarily homogeneous polynomials. We set $F=\{f_{1},\ldots,f_{m}\}$. Assume that $R/\langle F^{\rm top}\rangle$ is Artinian, and that $\bm{F}^{\rm top}$ is cryptographic semi-regular. Then we have the following:
  1. 1.
    $\widetilde{d}_{\rm reg}(\langle F^{h}\rangle)\geq d_{\rm reg}(\langle F^{\rm top}\rangle)-1$.
  2. 2.
    $\mathrm{max.GB.deg}_{\prec^{h}}(F^{h})\leq\max\{{d}_{\rm reg}(\langle F^{\rm top}\rangle),\widetilde{d}_{\rm reg}(\langle F^{h}\rangle)\}$ for any graded monomial order $\prec$.
Moreover, the equality holds in the second inequality if $\langle\mathrm{LM}(\langle F^{h}\rangle)\rangle$ is weakly reverse lexicographic. Here, a weakly reverse lexicographic ideal is a monomial ideal $J$ such that if $x^{\alpha}$ is one of the minimal generators of $J$ then every monomial of the same degree which preceeds $x^{\alpha}$ must belong to $J$ as well (see [36, Section 4] for the original definition).
Proof.
Put $D:=\widetilde{d}_{\rm reg}(\langle F^{\rm top}\rangle)$ and $D^{\prime}:=\widetilde{d}_{\rm reg}(\langle F^{h}\rangle)$. By the proof of Theorem 3.1.1 below, we have the following exact sequence:
[…]
for each $d$, where $K_{\bullet}$ denotes the Koszul complex on $\bm{F}^{h}$. Here we have $H_{1}(K_{\bullet}\otimes_{R^{\prime}}R)_{d}=0$ and $H_{0}(K_{\bullet}\otimes_{R^{\prime}}R)_{d}\neq 0$ for $d\leq D-1$, whence $\dim_{K}(R^{\prime}/\langle F^{h}\rangle)_{d-1}<\dim_{K}(R^{\prime}/\langle F^{h}\rangle)_{d}$ for any such $d$. Therefore, the first argument to be proved holds. On the other hand, it follows from the definition of $d_{\rm reg}$ that $H_{0}(K_{\bullet}\otimes_{R^{\prime}}R)_{d}=0$ for any $d$ with $d\geq D$. Thus, for such any $d$, the multiplication-by-$y$ map $H_{0}(K_{\bullet})_{d-1}\longrightarrow H_{0}(K_{\bullet})_{d}$ is surjective, and it is bijective if and only if $\dim_{K}(R^{\prime}/\langle F^{h}\rangle)_{d-1}=\dim_{K}(R^{\prime}/\langle F^{h}\rangle)_{d}$. If $D^{\prime}\geq D$, for any $d$ with $d\geq D^{\prime}$, the multiplication-by-$y$ map $H_{0}(K_{\bullet})_{d}\longrightarrow H_{0}(K_{\bullet})_{d+1}$ is bijective, since $d+1\geq D^{\prime}+1\geq D$. By this together with the surjectivity of $H_{0}(K_{\bullet})_{D^{\prime}-1}\longrightarrow H_{0}(K_{\bullet})_{D^{\prime}}$, it follows from Lemma 2.2.2 that $\mathrm{max.GB.deg}_{\prec^{h}}(F^{h})$ is upper-bounded by $D^{\prime}=\max\{D,D^{\prime}\}$. Also if $D^{\prime}=D-1$, for any $d$ with $d\geq D^{\prime}$, the multiplication-by-$y$ map $H_{0}(K_{\bullet})_{d}\longrightarrow H_{0}(K_{\bullet})_{d+1}$ is bijective, since $d+1\geq D^{\prime}+1=D$. However, $H_{0}(K_{\bullet})_{D^{\prime}-1}\longrightarrow H_{0}(K_{\bullet})_{D^{\prime}}$ is injective but not surjective, we cannot apply Lemma 2.2.2 for $d_{0}=D^{\prime}$, but apply it for $d_{0}=D^{\prime}+1=D$. Hence, $\mathrm{max.GB.deg}_{\prec^{h}}(F^{h})$ is upper-bounded by $D^{\prime}+1=D=\max\{D,D^{\prime}\}$, as desired.
Assume that $\langle\mathrm{LM}(\langle F^{h}\rangle)\rangle$ is a weakly reverse lexicographic ideal. First we consider the case where $\widetilde{d}_{\rm reg}(\langle F^{h}\rangle)\geq d_{\rm reg}(\langle F^{\rm top}\rangle)$, so that $\max\{{d}_{\rm reg}(\langle F^{\rm top}\rangle),\widetilde{d}_{\rm reg}(\langle F^{h}\rangle)\}=\widetilde{d}_{\rm reg}(\langle F^{h}\rangle)$. We put $d:=\mathrm{max.GB.deg}_{\prec^{h}}(F^{h})$, and let $\{t_{1},\ldots,t_{r}\}$ be a standard monomial basis of $R^{\prime}_{d}/\langle F^{h}\rangle_{d}$, that is,
$\{t\in R^{\prime}_{d}:\mbox{$t$ is a monomial and }{t\notin\langle{\rm LM}(\langle F^{h}\rangle)\rangle}\}.$
For $s\geq 1$, we prove that $\{t_{1}y^{s},\ldots,t_{r}y^{s}\}$ is a basis of $R^{\prime}_{d+s}/\langle F^{h}\rangle_{d+s}$, from which we have $\dim_{K}R^{\prime}_{d}/\langle F^{h}\rangle_{d}=\dim_{K}R^{\prime}_{d+s}/\langle F^{h}\rangle_{d+s}$ for any positive integer $s$, so that $d\geq\widetilde{d}_{\rm reg}(\langle F^{h}\rangle)$ and therefore $d=\widetilde{d}_{\rm reg}(\langle F^{h}\rangle)$.
By Remark 3.1.2 and Lemma 2.3.5 below, the multiplication-by-$y^{s}$ map from $R_{d}^{\prime}/\langle F^{h}\rangle_{d}$ to $R^{\prime}_{d+s}/\langle F^{h}\rangle_{d+s}$ is surjective. Therefore $B_{s}:=\{t_{1}y^{s},\ldots,t_{r}y^{s}\}$ generates $R^{\prime}_{d+s}/\langle F^{h}\rangle_{d+s}$. Suppose to contrary that $B_{s}$ is not a basis. In this case, there exists an $i$ such that $t_{i}y^{s}$ is divisible by $\mathrm{LM}(g)$ for some $g\in G_{\rm hom}$, but $t_{i}$ is not divisible by $\mathrm{LM}(g)$, where $G_{\rm hom}$ is the reduced Gröbner basis of $\langle F^{h}\rangle$ with respect to $\prec^{h}$. Therefore, we can write $t_{i}y^{s}=s_{i}u_{i}y^{s}$ with $s_{i}u_{i}=t_{i}$ and $u_{i}y^{s^{\prime}}=\mathrm{LM}(g)$ for some $s^{\prime}\leq s$. (We note that $\deg s_{i}=d-\deg u_{i}\geq\deg{\rm LM}(g)-\deg u_{i}=s^{\prime}$ and $u_{i}={\rm GCD}(t_{i},{\rm LM}(g))$.) Note that $s^{\prime}\geq 1$ since otherwise $t_{i}$ is divisible by $\mathrm{LM}(g)$.
Take an arbitrary monomial $s^{\prime}_{i}$ such that $\deg s^{\prime}_{i}=s^{\prime}$ and $s_{i}^{\prime}$ divides $s_{i}$. Then, by the weakly reverse lexicographicness, as ${\rm LM}(g)=u_{i}y^{s^{\prime}}\preceq u_{i}s^{\prime}_{i}$, the monomial $s^{\prime}_{i}u_{i}$ should belong to $\langle{\rm LM}(G_{\rm hom})\rangle$. Moreover, since $s^{\prime}_{i}u_{i}$ divides $t_{i}=s_{i}u_{i}$, the monomial $t_{i}=s_{i}u_{i}$ also belongs to $\langle{\rm LM}(G_{\rm hom})\rangle$, which is a contradiction.
Finally, we consider the remaining case, namely we have $d_{\rm reg}(\langle F^{\rm top}\rangle)>\widetilde{d}_{\rm reg}(\langle F^{h}\rangle)$, so that $\max\{{d}_{\rm reg}(\langle F^{\rm top}\rangle),\widetilde{d}_{\rm reg}(\langle F^{h}\rangle)\}=d_{\rm reg}(\langle F^{\rm top}\rangle)$. In this case, it follows from Lemma 2.3.5 below that $d\geq d_{\rm reg}(\langle F^{\rm top}\rangle)$, whence $d=d_{\rm reg}(\langle F^{\rm top}\rangle)$. ∎
Remark 2.3.4.
In the proof of Proposition 2.3.3, the case where $D^{\prime}\geq D$ (i.e., $D^{\prime}>D-1$) means that
$\cdots<{\rm HF}_{A^{\prime}}(D-2)<{\rm HF}_{A^{\prime}}(D-1)\geq{\rm HF}_{A^{\prime}}(D)\geq\cdots\geq{\rm HF}_{A^{\prime}}(D^{\prime})={\rm HF}_{A^{\prime}}(D^{\prime}+1)=\cdots$
with $A^{\prime}:=R^{\prime}/\langle F^{h}\rangle$, and the case where $D^{\prime}=D-1$ (i.e., $D^{\prime}+1=D$) means that
$\cdots<{\rm HF}_{A^{\prime}}(D-2)<{\rm HF}_{A^{\prime}}(D-1)={\rm HF}_{A^{\prime}}(D)=\cdots.$
Lemma 2.3.5.
Let $\bm{F}=(f_{1},\ldots,f_{m})\in(R\smallsetminus K)^{m}$ be a sequence of not necessarily homogeneous polynomials. We set $F=\{f_{1},\ldots,f_{m}\}$, and assume that $R/\langle F^{\rm top}\rangle$ is Artinian and that $\bm{F}^{\rm top}$ is cryptographic semi-regular. We also suppose that $\langle\mathrm{LM}(\langle F^{h}\rangle)\rangle$ is weakly reverse lexicographic. Then we have $\mathrm{max.GB.deg}_{\prec^{h}}(F^{h})\geq d_{\rm reg}(\langle F^{\rm top}\rangle)$ and $\mathrm{max.GB.deg}_{\prec}(F^{\rm top})=d_{\rm reg}(\langle F^{\rm top}\rangle)$.
Proof.
Put $d:=\mathrm{max.GB.deg}_{\prec^{h}}(F^{h})$ and assume for a contradiction that $d<D:=d_{\rm reg}(\langle F^{\rm top}\rangle)$. By Lemma 4.1.4 below, we have ${\rm LM}(G_{\rm hom})={\rm LM}(G_{\rm top})_{\leq d}\subset R$. Then $\langle{\rm LM}(G_{\rm hom})\rangle_{R}$ contains ${\rm LM}(G_{\rm top})$. This can be shown as follows: For any $t\in{\rm LM}(G_{\rm top})$, there are polynomials $a_{1},\ldots,a_{m}$ in $R$ such that $t={\rm LM}(\sum_{i=1}^{m}a_{i}f_{i}^{\rm top})$. In this case, it can be easily checked that $t={\rm LM}(\sum_{i=1}^{m}a_{i}f_{i}^{h})$, and thus $t$ is divisible by some element of ${\rm LM}(G_{\rm hom})$.
Thus, as $R/\langle F^{\rm top}\rangle=R/\langle G_{\rm top}\rangle$ is Artinian, it follows from $\langle{\rm LM}(G_{\rm hom})\rangle_{R}\supset{\rm LM}(G_{\rm top})$ that there is an element $g$ in $G_{\rm hom}$ with ${\rm LM}(g)=x_{n}^{d^{\prime}}$ for some $d^{\prime}\leq d$. Then, for any monomial $t\in R_{d^{\prime}}$, as $t\succeq x_{n}^{d^{\prime}}$, it belongs to $\langle{\rm LM}(\langle F^{h}\rangle)\rangle_{R^{\prime}}(=\langle{\rm LM}(G_{\rm hom})\rangle_{R^{\prime}}=\langle{\rm LM}(G_{\rm top})_{\leq d}\rangle_{R^{\prime}})$ by its weakly reverse lexicographicness. This implies that $t$ also belongs to $\langle{\rm LM}(G_{\rm top})\rangle_{R}$, and hence $R_{d}/\langle F^{\rm top}\rangle_{d}=0$. Thus, we have $d\geq d^{\prime}\geq d_{\rm reg}(\langle F^{\rm top}\rangle)$, a contradiction.
By the same argument as above, we can show that $\mathrm{max.GB.deg}_{\prec}(F^{\rm top})=d_{\rm reg}(\langle F^{\rm top}\rangle)$ as follows. It is obvious that $\mathrm{max.GB.deg}_{\prec}(F^{\rm top})\leq d_{\rm reg}(\langle F^{\rm top}\rangle)$. Thus, we assume to the contrary that $d:=\mathrm{max.GB.deg}_{\prec}(F^{\rm top})<D:=d_{\rm reg}(\langle F^{\rm top}\rangle)$. Then, it follows from Lemma 4.1.4 below that ${\rm LM}(G_{\rm top})={\rm LM}(G_{\rm hom})_{\leq d}$ and the ideal $\langle{\rm LM}(G_{\rm top})\rangle_{R}$ has the weak reverse lexicographicness up to $d$. Since there is an element $g$ in $G_{\rm top}$ with ${\rm LM}(g)=x_{n}^{d^{\prime}}$ for some $d^{\prime}\leq d$, any monomial in $R_{d^{\prime}}$ belongs to $\langle{\rm LM}(G_{\rm top})\rangle$ and so $R_{d^{\prime}}/\langle F^{\rm top}\rangle_{d^{\prime}}=0$, which implies $d\geq d^{\prime}\geq D$, a contradiction. ∎
Remark 2.3.6.
In Lemma 2.3.5, when $R/\langle F^{\rm top}\rangle$ is Artinian, we can easily prove the equality $\mathrm{max.GB.deg}_{\prec}(F^{\rm top})=d_{\rm reg}(\langle F^{\rm top}\rangle)$ if $\langle\mathrm{LM}(\langle F^{\rm top}\rangle)\rangle$ is weakly reverse lexicographic, not assuming that $\bm{F}^{\rm top}$ is cryptographic semi-regular nor that $\langle\mathrm{LM}(\langle F^{h}\rangle)\rangle$ is weakly reverse lexicographic.
3 Proofs of Theorems 1 and 2
In this section, we shall prove Theorems 1 and 2 stated in Section 1. As in the previous section, let $K$ be a field, and $R=K[X]=K[x_{1},\ldots,x_{n}]$ denote the polynomial ring of $n$ variables over $K$. We denote by $R_{d}$ the homogeneous part of degree $d$, that is, the set of homogeneous polynomials of degree $d$ and 00. As in Theorems 1 and 2, let $F=\{f_{1},\ldots,f_{m}\}$ be a set of not necessarily homogeneous polynomials in $R$ of positive degrees $d_{1},\ldots,d_{m}$, and put $\bm{F}=(f_{1},\ldots,f_{m})$. Recall Definition 2.1.9 for the definition of cryptographic semi-regular sequences.
3.1 Bounded regularity of homogenized sequences
Here we revisit the main results in [30, Section 4]. For the readability, we remain the proofs. Also, as additional remarks, we explicitly give an important property of the Hilbert-Poincaré series of $R^{\prime}/\langle F^{h}\rangle$ with $R^{\prime}=R[y]$, and also give an alternative proof for [30, Theoem 7] (Theorem 3.1.1 below).
The Hilbert-Poincaré series associated to a (homogeneous) cryptographic semi-regular sequence is given by (2.1.2). On the other hand, the Hilbert-Poincaré series associated to the homogenizaton $F^{h}$ cannot be computed without knowing its Gröbner basis in general, but we shall prove that it can be computed up to the degree $d_{\rm reg}(\langle F^{\rm top}\rangle)-1$ if $\bm{F}$ is affine cryptographic semi-regular, namely $\bm{F}^{\rm top}=(f_{1}^{\rm top},\ldots,f_{m}^{\rm top})$ is cryptographic semi-regular.
Theorem 3.1.1 (Theorem 1 (1); [30, Theoem 7]).
Let $R^{\prime}=R[y]$, and let $\bm{F}=(f_{1},\ldots,f_{m})$ be a sequence of not necessarily homogeneous polynomials in $R$ of positive degrees. Assume that $\bm{F}$ is affine cryptographic semi-regular. Then, for each $d$ with $d<D:=d_{\mathrm{reg}}(\langle F^{\rm top}\rangle)$, we have
${\rm HF}_{R^{\prime}/\langle F^{h}\rangle}(d)={\rm HF}_{R/\langle F^{\rm top}\rangle}(d)+{\rm HF}_{{\color[rgb]{0,0,0}R^{\prime}}/\langle F^{h}\rangle}(d-1),$ (3.1.1)
and hence
${\rm HF}_{R^{\prime}/\langle F^{h}\rangle}(d)={\rm HF}_{R/\langle F^{\rm top}\rangle}(d)+\cdots+{\rm HF}_{R/\langle F^{\rm top}\rangle}(0),$ (3.1.2)
whence we can compute the value ${\rm HF}_{R^{\prime}/\langle F^{h}\rangle}(d)$ from the formula (2.1.2).
Proof.
Let $K_{\bullet}=K_{\bullet}(f_{1}^{h},\ldots,f_{m}^{h})$ be the Koszul complex on $(f_{1}^{h},\ldots,f_{m}^{h})$, which is given by (A.1.1). By tensoring $K_{\bullet}$ with $R^{\prime}/\langle y\rangle_{R^{\prime}}\cong K[x_{1},\ldots,x_{n}]=R$ over $R^{\prime}$, we obtain the following exact sequence of chain complexes:
[…]
where ×$\times y$ is a graded homomorphism of degree $1$ multiplying each entry of a vector with $y$, and where $\pi_{i}$ is a canonical homomorphism sending $v\in K_{i}$ to $v_{i}\otimes 1\in K_{i}\otimes_{R^{\prime}}R$. Note that there is an isomorphism
$K_{i}\otimes_{R^{\prime}}R\cong\bigoplus_{1\leq j_{1}<\cdots<j_{i}\leq m}R(-d_{j_{1}\cdots j_{i}})\mathbf{e}_{j_{1}\cdots{\color[rgb]{0,0,0}{j_{i}}}},$
via which we can interpret $\pi_{i}:K_{i}\to K_{i}\otimes_{R^{\prime}}R$ as a homomorphism that projects each entry of a vector in $K_{i}$ modulo $y$. In particular, we have
$K_{0}\otimes_{R^{\prime}}R$ $=$ $R^{\prime}/\langle f_{1}^{h},\ldots,f_{m}^{h}\rangle_{R^{\prime}}\otimes_{R^{\prime}}R^{\prime}/\langle y\rangle_{R^{\prime}}$
$\cong$ $R^{\prime}/\langle f_{1}^{h},\ldots,f_{m}^{h},y\rangle_{R^{\prime}}$
$\cong$ $R/\langle f_{1}^{\rm top},\ldots,f_{m}^{\rm top}\rangle_{R}$
for $i=0$. This means that the chain complex $K_{\bullet}\otimes_{R^{\prime}}R$ gives rise to the Kosuzul complex on $({f}_{1}^{\rm top},\ldots,f_{m}^{\rm top})$. We induce a long exact sequence of homology groups. In particular, for each degree $d$, we have the following long exact sequence:
[…]
where $\delta_{i+1}$ is a connecting homomorphism produced by the Snake lemma. For $i=0$, we have the following exact sequence:
[…]
From our assumption that $F^{\rm top}$ is cryptographic semi-regular, it follows from Proposition 2.1.10 that $H_{1}(K_{\bullet}\otimes_{R^{\prime}}R)_{\leq D-1}=0$ for $D:=d_{\rm reg}(\langle F^{\rm top}\rangle)$. Therefore, if $d\leq D-1$, we have an exact sequence
[…]
of $K$-linear spaces, so that
$\mathrm{dim}_{K}H_{0}(K_{\bullet})_{d}=\mathrm{dim}_{K}H_{0}(K_{\bullet}\otimes_{R^{\prime}}R)_{d}+\mathrm{dim}_{K}H_{0}(K_{\bullet})_{d-1}$ (3.1.3)
by the dimension theorem. Since $H_{0}(K_{\bullet})=R^{\prime}/\langle F^{h}\rangle$ and ${\color[rgb]{0,0,0}H_{0}(K_{\bullet}\otimes_{R^{\prime}}R)}=R/\langle F^{\rm top}\rangle$, we have the equality (3.1.1), as desired. ∎
Remark 3.1.2 (Theorem 1 (2), (3); [30, Remark 6]).
Note that, in the proof of Theorem 3.1.1, the multiplication map $H_{0}(K_{\bullet})_{d-1}\to H_{0}(K_{\bullet})_{d}$ by $y$ is injective for all $d<D$, whence ${\rm HF}_{R^{\prime}/\langle F^{h}\rangle}(d)$ is monotonically increasing for $d<D$. On the other hand, since $H_{0}(K_{\bullet}\otimes_{R^{\prime}}R)_{d}=(R/\langle F^{\rm top}\rangle)_{d}=0$ for all $d\geq D$ by the definition of the degree of regularity, the multiplication map $H_{0}(K_{\bullet})_{d-1}\to H_{0}(K_{\bullet})_{d}$ by $y$ is surjective for all $d\geq D$, whence ${\rm HF}_{R^{\prime}/\langle F^{h}\rangle}(d)$ is monotonically decreasing for $d\geq D-1$. By this together with [10, Theorem 3.3.4], the homogeneous ideal $\langle F^{h}\rangle$ is zero-dimensional or trivial, i.e., there are at most a finite number of projective zeros of $F^{h}$ (and thus there are at most a finite number of affine zeros of $F$).
Remark 3.1.3.
We note that, for each $d\geq D$, the condition $\dim_{K}H_{0}(K_{\bullet})_{d-1}\!=\!\dim_{K}H_{0}(K_{\bullet})_{d}$ is equivalent to that the multiplication map $H_{0}(K_{\bullet})_{d-1}\to H_{0}(K_{\bullet})_{d}$ by $y$ is injective (and thus bijective). By this together with Lemma 2.2.2, letting $d_{0}$ be the smallest number with $d_{0}\geq D$ such that $\dim_{K}H_{0}(K_{\bullet})_{d_{0}}=\dim_{K}H_{0}(K_{\bullet})_{d_{0}+1}$, the maximal Gröbner basis degree of $\langle F^{h}\rangle$ is upper-bounded by $d_{0}$.
Remark 3.1.4.
We have another proof of Theorem 1 (1), (2) by using the following exact sequence:
[…]
Then, as an easy consequence, for $d\in\mathbb{N}$, we have
${\rm HF}_{R^{\prime}/\langle F^{h}\rangle}(d)={\rm HF}_{R^{\prime}/(\langle F^{h}\rangle+\langle y\rangle)}(d)+{\rm HF}_{R^{\prime}/(\langle F^{h}\rangle:\langle y\rangle)}(d-1),$
see [25, Lemmas 5.2.1 and 5.2.2]. Note that ${\rm HF}_{R^{\prime}/(\langle F^{h}\rangle+\langle y\rangle)}(d)={\rm HF}_{R/\langle F^{\rm top}\rangle}(d)$ for any positive integer $d$. On the other hand, for $d<D$, any degree-fall does not occur, that is, if $yf\in\langle F^{h}\rangle_{d}$ with $f\in R^{\prime}$ then $f\in\langle F^{h}\rangle_{d-1}$. This can be shown by some semantic argument (see Remark 4.1.3) or also rigidly by the injectiveness of the multiplication map of $y$ in (3.1.3). Thus, we also have $\langle f\in R[y]:fy\in\langle F^{h}\rangle\rangle_{d-1}=\langle F^{h}\rangle_{d-1}$, so that
$\dim_{K}(R^{\prime}/(\langle F^{h}\rangle:\langle y\rangle))_{d-1}=\dim_{K}(R^{\prime}/\langle F^{h}\rangle)_{d-1},$
namely ${\rm HF}_{R^{\prime}/(\langle F^{h}\rangle:\langle y\rangle)}(d-1)={\rm HF}_{R^{\prime}/\langle F^{h}\rangle}(d-1)$, and hence we have (3.1.1) for $d<D$. For $d\geq D$, since $(R/\langle F^{\rm top}\rangle)_{d}=0$ by the definition of $D$, we have
$\dim_{K}(R^{\prime}/\langle F^{h}\rangle)_{d}={\rm HF}_{R^{\prime}/\langle F^{h}\rangle}(d)={\rm HF}_{R^{\prime}/(\langle F^{h}\rangle:\langle y\rangle)}(d-1)=\dim_{K}(R^{\prime}/(\langle F^{h}\rangle:\langle y\rangle))_{d-1}.$ (3.1.4)
Now we consider the following multiplication map by $y$:
×$\times y:(R^{\prime}/\langle F^{h}\rangle)_{d-1}\longrightarrow(R^{\prime}/\langle F^{h}\rangle)_{d}\ ;\ g\mapsto yg.$
Since ${\rm Ker}(\times y)=(\langle F^{h}\rangle:\langle y\rangle)_{d-1}/\langle F^{h}\rangle_{d-1}$, we have
$\dim_{K}R^{\prime}_{d}/\langle F^{h}\rangle_{d}$$\geq$ $\dim_{K}({\rm Im}(\times y))$ (3.1.5)
$=$ $\dim_{K}(R^{\prime}/\langle F^{h}\rangle)_{d-1}-\dim_{K}((\langle F^{h}\rangle:\langle y\rangle)/\langle F^{h}\rangle)_{d-1}$
$=$ $\dim_{K}R^{\prime}_{d-1}-\dim_{K}(\langle F^{h}\rangle:\langle y\rangle)_{d-1}$
$=$ $\dim_{K}(R^{\prime}/(\langle F^{h}\rangle:\langle y\rangle))_{d-1}.$
Since the both ends of (3.1.4) and (3.1.5) coincide, we have ${\rm Im}(\times y)=(R^{\prime}/\langle F^{h}\rangle)_{d}$, that is, the multiplication map by $y$ is surjective.
The Hilbert-Poincaré series of $R^{\prime}/\langle F^{h}\rangle$ satisfies the following equality (3.1.6):
Corollary 3.1.5 (Theorem 1 (3); [30, Corollary 1]).
Let $D=d_{\rm reg}(\langle F^{\rm top}\rangle)$. Then we have
${\rm HS}_{R^{\prime}/\langle F^{h}\rangle}(z)\equiv\frac{\prod_{i=1}^{m}(1-z^{d_{i}})}{(1-z)^{n+1}}\pmod{z^{D}}.$ (3.1.6)
Therefore, by Theorem 2.1.4 ([14, Theorem 1]), the sequence $\bm{F}^{h}$ is $D$-regular. Here, we note that $D=\deg({\rm HS}_{R/\langle F^{\rm top}\rangle}{)}+1={\deg\left(\left[\frac{\prod_{i=1}^{m}(1-z^{d_{i}})}{(1-z)^{n}}\right]\right)}+1$.
3.2 Solving degree for homogenized sequences
Here we assume that $\bm{F}^{\rm top}$ is semi-regular and that all degrees $d_{i}=\deg(f_{i})$ are smaller or equal to the degree of regularity $d_{\rm reg}(\langle F^{\rm top}\rangle)$. Then, any $n$-subsequence of $\bm{F}^{\rm top}$ is regular. Under this assumption, we can give a detailed discussion on the solving degree of $F^{h}$. From now on, we assume that $m\geq n$, and set $\bm{F}_{k}:=(f_{1},\ldots,f_{n+k})$ and $D_{k}:=d_{\rm reg}(\langle F_{k}^{\rm top}\rangle)$ for each $k\geq 0$. As $\bm{F}_{0}^{\rm top}$ is regular and $\bm{F}_{1}^{\rm top}$ is semi-regular, we have $D_{0}=d_{1}+\cdots+d_{n}-n+1$ and $D_{1}=\lfloor\frac{d_{1}+\cdots+d_{n+1}-n-1}{2}\rfloor+1$, see [4, Theorem 4.1]. Thus, by setting $d_{1}\leq d_{2}\leq\cdots\leq d_{m}$, we can minimize the values $D_{0}$ and $D_{1}$.
Remark 3.2.1.
Our estimations on the solving degree below require that $\bm{F}_{1}^{\rm top}$ is semi-regular. Thus, even when $\bm{F}^{\rm top}$ is not semi-regular, if there is an $(n+1)$-subset which is semi-regular, we may assume that $\bm{F}_{1}^{\rm top}$ is semi-regular and apply our arguments below.
We denote by $K_{\bullet}^{(j,{\rm top})}$ the Koszul complex on $(f_{1}^{\rm top},\ldots,f_{j}^{\rm top})$, and let
$K_{\bullet}^{(j-1,{\rm top})}(-d_{j})\xrightarrow{\times f_{j}^{\rm top}}K_{\bullet}^{(j-1,{\rm top})}$
be a graded homomorphism of degree $d_{j}$ multiplying each entry of a vector with $f_{j}^{\rm top}$. (This kind of complex is also used in [14].) Regarding $K_{\bullet}^{(j,{\rm top})}$ as the mapping cone of the above ×$\times f_{j}^{\rm top}$, we obtain the following short exact sequence of complexes
[…]
where $K_{\bullet}^{(j-1,{\rm top})}[-1]$ is a shifted complex defined by $K_{\bullet}^{(j-1,{\rm top})}[-1]_{i}=K_{i-1}^{(j-1,{\rm top})}$, and where $K_{i}^{(j,{\rm top})}\cong K_{i}^{(j-1,{\rm top})}\oplus K_{i-1}^{(j-1,{\rm top})}(-d_{j})$, for example
$K_{1}^{(j,{\rm top})}=\bigoplus_{s=1}^{j}R(-d_{s})\cong\left(\bigoplus_{s=1}^{j-1}R(-d_{s})\right)\oplus R(-d_{j})=K_{1}^{(j-1,{\rm top})}\oplus K_{0}^{(j-1,{\rm top})}(-d_{j}).$
Note also that $K_{\bullet}^{(j-1,{\rm top})}\longrightarrow K_{\bullet}^{(j,{\rm top})}$ and $K_{\bullet}^{(j,{\rm top})}\longrightarrow K_{\bullet}^{(j-1,{\rm top})}[-1](-d_{j})$ are the canonical inclusion and projection respectively. Then we deduce the following exact sequence from the Snake lemma:
[…]
where $\delta_{i}$ denotes a connecting homomorphism. Note that $\delta_{i}$ coincides with the multiplication map by $f_{j}^{\rm top}$ on
$H_{i}(K_{\bullet}^{(j-1,{\rm top})}(-d_{j}))\longrightarrow H_{i}(K_{\bullet}^{(j-1,{\rm top})})$
induced from that on $K_{\bullet}^{(j-1,{\rm top})}(-d_{j})\longrightarrow K_{\bullet}^{(j-1,{\rm top})}$ (this is also derived from general facts in homological algebra). Since $H_{-1}(K_{\bullet}^{(j-1,{\rm top})})=0$, we can rewrite the above long exact sequence as
[…]
In particular, for $i=0$ and for each degree $d$, we have the following exact sequence:
[…]
Now consider $H_{1}(K_{\bullet}^{(m,{\rm top})})$ for $m\geq n+1$. Here we remark that $H_{i}(K_{\bullet}^{(n,{\rm top})})=0$ for all $i$ with $i\geq 1$, since the sequence $\bm{F}_{0}^{\rm top}=(f_{1}^{\rm top},\ldots,f_{n}^{\rm top})$ is regular by our assumption.
Proposition 3.2.2.
Suppose that $d_{1}\leq d_{2}\leq\cdots\leq d_{m}$ and $m>n$. If $\bm{F}^{\rm top}$ is semi-regular, then $H_{1}(K_{\bullet}^{(m,{\rm top})})_{d}=0$ for any $d$ with $d\geq D_{0}+d_{m}$. Moreover, if $d_{m}\leq D_{1}$, then $H_{1}(K_{\bullet}^{(m,{\rm top})})_{d}=0$ for any $d$ with $d\geq D_{0}+d_{n+1}$.
Proof.
First consider the case where $m=n+1$. For $d\geq D_{0}+d_{n+1}$, as $d-d_{n+1}\geq D_{0}$, we have $H_{0}(K_{\bullet}^{(n,{\rm top})})_{d-d_{n+1}}=0$. Therefore, for any $d$ with $d\geq D_{0}+d_{n+1}$, we obtain an exact sequence
[…]
so that $H_{1}(K_{\bullet}^{(n+1,{\rm top})})_{d}=0$.
Next we consider the case where $m\geq n+1$ and we show that $H_{1}(K_{\bullet}^{(m,{\rm top}})_{d}=0$ for $d\geq D_{0}+d_{m}$ by the induction on $m$. So we assume that $H_{1}(K_{\bullet}^{(m,{\rm top}})_{d}=0$ for $d\geq D_{0}+d_{m}$. Then, for $d\geq D_{0}+d_{m+1}\geq D_{0}+d_{m}$, we have an exact sequence
[…]
It follows from $H_{0}(K_{\bullet}^{(n,{\rm top})})_{d^{\prime}}=0$ for $d^{\prime}\geq D_{0}$ that $H_{0}(K_{\bullet}^{(m,{\rm top})})_{d^{\prime}}=0$ by $F_{m-n}^{{\rm top}}\supset F_{0}^{{\rm top}}$. Therefore, we also have $H_{0}(K_{\bullet}^{(m,{\rm top})})_{d-d_{m+1}}=0$ by $d-d_{m+1}\geq D_{0}$, whence $H_{1}(K_{\bullet}^{(m+1,{\rm top})})_{d}=0$.
Finally we consider the case where $d_{m}\leq D_{1}$ and show $H_{1}(K_{\bullet}^{(m,{\rm top})})_{d}=0$ for $d\geq D_{0}+d_{n+1}$ by the induction on $m$ in a similar manner as above. So we assume that $H_{1}(K_{\bullet}^{(m,{\rm top})})_{d}=0$ for $d\geq D_{0}+d_{n+1}$. Then, we consider the sequence (3.2.1) for $d\geq D_{0}+d_{n+1}$ again. Thus it suffices to show that $H_{0}(K_{\bullet}^{(m,{\rm top})})_{d-d_{m+1}}=0$.
Using $D_{1}=\left\lfloor\frac{d_{1}+\cdots+d_{n+1}-n-1}{2}\right\rfloor+1\geq d_{m+1}$, we have
$d-d_{m+1}$$\geq$ $D_{0}+d_{n+1}-d_{m+1}$
$\geq$ $(d_{1}+\cdots+d_{n+1}-n-1)+2-\left(\left\lfloor\frac{d_{1}+\cdots+d_{n+1}-n-1}{2}\right\rfloor+1\right)$
$\geq$ $\left\lfloor\frac{d_{1}+\cdots+d_{n+1}-n-1}{2}\right\rfloor+1=D_{1}.$
Thus, it follows that $H_{0}(K_{\bullet}^{(n+1,{\rm top})})_{d-d_{m+1}}=0$. Since one has $\langle F_{m-n}^{\rm top}\rangle\supset\langle F_{1}^{\rm top}\rangle$, the condition $H_{0}(K_{\bullet}^{(n+1,{\rm top})})_{d-d_{m+1}}=0$ implies $H_{0}(K_{\bullet}^{(m,{\rm top})})_{d-d_{m+1}}=0$, as desired. ∎
Theorem 3.2.3 (Theorem 2).
Suppose that that $d_{1}\leq d_{2}\leq\cdots\leq d_{m}$ and $m>n$. If $\bm{F}^{\rm top}$ is semi-regular, then the generalized degree of regularity of $\langle F^{h}\rangle$ is upper-bounded by $d_{1}+d_{2}+\cdots+d_{n}+d_{m}-n$ and so the solving degree of $F^{h}$. Moreover, if $d_{m}\leq D_{1}$, the generalized degree of regularity of $\langle F^{h}\rangle$ is upper-bounded by $d_{1}+\cdots+d_{n}+d_{n+1}-n$ and so the solving degree of $F^{h}$.
Proof.
We recall the long exact sequence of homology groups derived from the following exact sequence considered in the proof of Theorem 3.1.1:
[…]
For $i=0$ and $d\in\mathbb{N}$, we have the following exact sequence:
[…]
Then, for $d\geq D_{0}+d_{m}$ (or $d\geq D_{0}+d_{n+1}$ if $d_{m}\leq D_{1}$), it follows from Proposition 3.2.2 that $H_{1}(K_{\bullet}(F^{\rm top}))_{d}=0$. Moreover, $H_{0}(K_{\bullet}(F^{\rm top}))_{d}=0$ also holds, since $d>D_{0}\geq D$. Therefore, we have an exact sequence
[…]
and, by letting $A=R^{\prime}/\langle F^{h}\rangle$, we have
$A_{d-1}=H_{0}(K_{\bullet}(F^{h}))_{d-1}\cong H_{0}(K_{\bullet}(F^{h}))_{d}=A_{d}$
for any $d\geq D_{0}+d_{m}$ (or $d\geq D_{0}+d_{n+1}$ if $d_{m}\leq D_{1}$). Moreover, the multiplication map by $y$ from $A_{d-1}$ to $A_{d}$ is a bijection. Thus, the generalized degree of regularity of $\langle F^{h}\rangle$ is bounded by $D_{0}+d_{m}-1$ (or $D_{0}+d_{n+1}-1$ if $d_{m}\leq D_{1}$). Then, by Proposition 2.3.3, it bounds the solving degree ($={\rm max.GB.deg}_{\prec^{h}}(F^{h})$). ∎
Remark 3.2.4.
The bound in Theorem 3.2.3 looks the same as Lazard’s bound (Theorem 2.2.1). However, in our bound, except $d_{m}$, the degrees $d_{1},\ldots,d_{n}$ are set in ascending order, while in Lazard’s bound they are set in descending order. We note that, when $d_{1}=\cdots=d_{m}$, these two bounds coincide with one another.
Finally in this subsection, under the assumption that $\bm{F}^{\rm top}$ is cryptographic semi-regular, we show that the generalized degree of regularity of $\langle F^{h}\rangle$ (and thus solving degree of $F^{h}$) can be bounded by $D$ plus the saturation exponent, say $S_{0}$ here, that is, the minimal integer $k$ such that $\langle F\rangle^{h}=(\langle F^{h}\rangle:y^{\infty})=(\langle F^{h}\rangle:y^{k})$. See [25, p. 81] for the definition of saturation exponent.
Proposition 3.2.5.
The generalized degree of regularity $\langle F^{h}\rangle$ and the solving degree of $F^{h}$ are both bounded by $D+S_{0}$.
Proof.
Consider the following exact sequence:
[…]
where $R^{\prime}/(\langle F^{h}\rangle:\langle y^{S_{0}}\rangle)=R^{\prime}/\langle F\rangle^{h}$. Then, we have
${\rm HS}_{R^{\prime}/\langle F^{h}\rangle}(z)={\rm HS}_{R^{\prime}/(\langle F^{h}\rangle+\langle y^{S_{0}}\rangle)}(z)+z^{S_{0}}{\rm HS}_{R^{\prime}/\langle F\rangle^{h}}(z).$
First, we show ${\rm HF}_{R^{\prime}/(\langle F^{h}\rangle+\langle y^{S_{0}}\rangle)}(d)=0$ for $d\geq D+S_{0}$, by which we have ${\rm HF}_{R^{\prime}/\langle F^{h}\rangle}(d)={\rm HF}_{R^{\prime}/\langle F\rangle^{h}}(d-S_{0})$. Suppose for a contradiction that $(R^{\prime}/(\langle F^{h}\rangle+\langle y^{S_{0}}\rangle))_{d}\neq 0$. Then, it follows from Macaulay’s basis theorem (cf. [28, Theorem 1.5.7]) that
$LB_{d}:=\{t\in R^{\prime}_{d}:\mbox{$t$ is a monomial and }{t\notin\langle{\rm LM}(\langle F^{h}\rangle+\langle y^{S_{0}}\rangle)\rangle}\}$
is a non-empty basis for the $K$-vector space $(R^{\prime}/(\langle F^{h}\rangle+\langle y^{S_{0}}\rangle))_{d}$. For any element $T$ in $LB_{d}$, if $T$ is divisible by $y^{S_{0}}$, then $T$ belongs to $(\langle F^{h}\rangle+\langle y^{S_{0}}\rangle)_{d}$, which is a contradiction. Otherwise, the degree of the $X$-part of $T$ is not smaller than $D$. Since ${\rm LM}(\langle F^{h}\rangle)$ contains any monomial in $X$ of degree $D$ by Lemma 4.1.4, it also contains $T$. Therefore $T\in{\rm LM}(\langle F^{h}\rangle+\langle y^{S_{0}}\rangle))$, which is a contradiction.
Next we show that ${\rm HF}_{R^{\prime}/\langle F\rangle^{h}}(d)$ becomes constant for $d\geq D$, which implies that ${\rm HF}_{R^{\prime}/\langle F^{h}\rangle}(d)$ becomes constant for $d\geq D+S_{0}$. Then, the generalized degree of regularity of $\langle F^{h}\rangle$ is bounded by $D+S_{0}$, and by Proposition 2.3.3, it follows that the solving degree of $F^{h}$ is bounded by $D+S_{0}$.
Let $G$ be the reduced Gröbner basis of $\langle F\rangle$ with respect to $\prec$. Then $G^{h}$ is a Gröbner basis of $\langle F\rangle^{h}$. By Lemma 4.2.4 below, we have ${\rm max.GB.deg}(F)\leq D$ and thus, any element of $G^{h}$ is of degree not greater than $D$. Then, let $\{t_{1},\ldots,t_{r}\}$ be the standard monomial basis of $R/\langle F\rangle$ as a $K$-vector space, that is, $\{t_{1},\ldots,t_{r}\}=\{t:{\rm LM}(g)\nmid t\mbox{ for any }g\in G\}$ with $r:=\dim_{K}R/\langle F\rangle$.
Again by Macaulay’s basis theorem, as a basis of the $K$-linear space $(R^{\prime}/\langle F\rangle^{h})_{d}$, we can take $LB_{d}^{\prime}=\{t\in R^{\prime}_{d}:\mbox{$t$ is a monomial and }{\rm LM}(g)\nmid t\mbox{ for any }g\in G^{h}\}$, which is equal to $\{t_{1}y^{k_{1}},\ldots,t_{r}y^{k_{r}}\}$ for $d\geq D$, where $\deg(t_{i}y^{k_{i}})=d$ for $1\leq i\leq r$. Thus, for $d\geq D$, it follows that $\dim_{K}(R^{\prime}/\langle F\rangle^{h})_{d}$ is equal to the constant $r$. ∎
4 Behaviors of Gröbner bases computation
Here we show certain correspondences in the Gröbner basis computations among inputs $F^{h}$, $F^{\rm top}$, and $F$. First we revisit the correspondence among the computation of the Gröbner basis of $F^{h}$ and that of $F^{\rm top}$ given in [30, Section 5.1]. Then, we explicitly give an important correspondence between the computation of the Gröbner basis of $F^{h}$ and that of $F$, which brings an upper-bound (Lemma 4.2.4 below) on the solving degree of $F$ related to Samaev-Tenti’s bound [40].
Here we use the same notation as in the previous section, and unless otherwise noted, assume that $\bm{F}$ is cryptographic semi-regular. Let $G$, $G_{\rm hom}$, and $G_{\rm top}$ be the reduced Gröbner bases of $\langle F\rangle$, $\langle F^{h}\rangle$, and $\langle F^{\rm top}\rangle$, respectively, where their monomial orderings are DRL $\prec$ or its extension $\prec^{h}$. Also we let $D=d_{\rm reg}(\langle F^{\rm top}\rangle)$, and assume $D<\infty$. Moreover, we extend the notion of top part to a homogeneous polynomial $h$ in $R^{\prime}=R[y]$ as follows. We call $h|_{y=0}$ the top part of $h$ and denote it by $h^{\rm top}$. Thus, if $h^{\rm top}$ is not zero, it coincides with the top part $(h|_{y=1})^{\rm top}$ of the dehomogenization $h|_{y=1}$ of $h$. We remark that $g^{\rm top}=(g^{h})^{\rm top}$ for a polynomial $g$ in $R$.
4.1 Correspondence between $G_{\rm hom}$ and $G_{\rm top}$
Here we revisit the results in [30, Section 5.1].
Corollary 4.1.1 ([30, Corollary 2]).
With notation as above, assume that $\bm{F}=(f_{1},\ldots,f_{m})\in R^{m}$ is affine cryptographic semi-regular. Put $\overline{I}:=\langle F^{\rm top}\rangle_{R}$ and $\tilde{I}:=\langle F^{h}\rangle_{R^{\prime}}$. Then, we have $(\langle\mathrm{LM}(\tilde{I})\rangle_{R^{\prime}})_{d}=(\langle\mathrm{LM}(\overline{I})\rangle_{R^{\prime}})_{d}$ for each $d$ with $d<D:=d_{\rm reg}(\overline{I})$.
Since $\bm{F}^{\rm top}$ is cryptographic semi-regular and since $\bm{F}^{h}$ is $D$-regular by Corollary 3.1.5, we obtain $H_{1}(K_{\bullet}(F^{\rm top}))_{{\color[rgb]{0,0,0}<D}}=H_{1}(K_{\bullet}(F^{h}))_{{\color[rgb]{0,0,0}<D}}=0$. Moreover, as $H_{1}(K_{\bullet}(F^{h}))={\rm syz}(F^{h})/{\rm tsyz}(F^{h})$ and $H_{1}(K_{\bullet}(F^{\rm top}))={\rm syz}(F^{\rm top})/{\rm tsyz}(F^{h})$ (see (A.1.2)), we have the following corollary, where ${\rm tsyz}$ denotes the module of trivial syzygies (see Definition A.1.1).
Corollary 4.1.2 ([14, Theorem 1]).
With notation as above, we have ${\rm syz}(F^{\rm top})_{<D}={\rm tsyz}(F^{\rm top})_{<D}$ and ${\rm syz}(F^{h})_{<D}={\rm tsyz}(F^{h})_{<D}$.
Remark 4.1.3.
[…]
In addition to the above facts, as mentioned (somehow implicitly) in [1, Section 3.5] and [3], when we compute a Gröbner basis of $\langle F^{h}\rangle$ for the degree less than $D$ by the $F_{5}$ algorithm with respect to $\prec^{h}$, for each computed non-zero polynomial $g$ from an S-polynomial, say $S(g_{1},g_{2})$, of degree less than $D$, its signature does not come from any trivial syzygy and so the reductions of $S(g_{1},g_{2})$ are done only at its top part. This implies that any degree-fall does not occur at each step degree less than $D$. This can be rigidly shown by using the injectiveness of the multiplication map by $y$ shown in Remark 3.1.2.
Now we recall that the Gröbner basis computation process of $\langle F^{h}\rangle$ corresponds exactly to that of $\langle F^{\rm top}\rangle$ at each step degree less than $D$. (We also discuss similar correspondences among the Gröbner basis computation of $\langle F^{h}\rangle$ and that of $\langle F\rangle$ in the next subsection.) Especially, the following lemma holds.
Lemma 4.1.4 ([30, Lemma 2]).
With notation as above, assume that $\bm{F}=(f_{1},\ldots,f_{m})\in R^{m}$ is affine cryptographic semi-regular. For each degree $d<D$, we have
${\rm LM}(G_{\rm hom})_{d}={\rm LM}(G_{\rm top})_{d}.$ (4.1.1)
We also note that the argument and the proof of Lemma 4.1.4 can be considered as a corrected version of [38, Theorem 4].
Next we consider $(G_{\rm hom})_{D}$. The following lemma holds, not assuming that $\bm{F}$ is affine cryptographic semi-regular:
Lemma 4.1.5 ([30, Lemma 3]).
Assume that $D=d_{\rm reg}(\langle F^{\rm top}\rangle)<\infty$ (the assumption that $\bm{F}$ is affine cryptographic semi-regular is not necessary). Then, for each monomial $M$ in $X$ of degree $D$, there is an element $g$ in $(G_{\rm hom})_{\leq D}$ such that ${\rm LM}(g)$ divides $M$. Therefore,
$\langle{\rm LM}((G_{\rm hom})_{\leq D})\rangle_{R^{\prime}}\cap R_{D}=R_{D}.$ (4.1.2)
Moreover, for each element $g$ in $(G_{\rm hom})_{D}$ with $g^{\rm top}\not=0$, the top-part $g^{\rm top}$ consists of one term, that is, $g^{\rm top}={\rm LT}(g)$, where ${\rm LT}$ denotes the leading term of $g$. (We recall ${\rm LT}(g)={\rm LC}(g){\rm LM}(g)$.)
Remark 4.1.6.
If we apply a signature-based algorithm such as the $F_{5}$ algorithm or its variant to compute the Gröbner basis of $\langle F^{h}\rangle$, its $\Sigma$-Gröbner basis is a Gröbner basis, but is not always reduced in the sense of ordinary Gröbner basis, in general. In this case, we have to compute so called inter-reduction among elements of the $\Sigma$-Gröbner basis to obtain the reduced Gröbner basis.
4.2 Correspondence between the computations of $G_{\rm hom}$ and $G$
In this subsection, we show that, at early stages, there is a strong correspondence between the computation of $G_{\rm hom}$ and that of $G$, from which we shall extend the upper bound on solving degree given in [40, Theorem 2.1] to our case.
Remark 4.2.1.
In [40], polynomial ideals over $R={\mathbb{F}}_{q}[x_{1},\ldots,x_{n}]$ are considered. Under the condition where the generating set $F$ contains the field equations $x_{i}^{q}-x_{i}$ for $1\leq i\leq n$, recall from Theorem 2.2.5 ([41, Theorem 6.5 & Corollary 3.67]) that the solving degree $\mathrm{sd}_{\prec}^{\mathcal{A}}(F)$ in the strict sense (see the definition (I) of Subsection 2.2 for the definition) with respect to a Buchberger-like algorithm $\mathcal{A}$ for $\langle F\rangle$ is upper-bounded by $22D-$ 2, where $D=d_{\rm deg}(\langle F^{\rm top}\rangle)$. In the proofs of [41, Theorem 6.5 & Corollary 3.67], the property $\langle F^{\rm top}\rangle_{D}=R_{D}$ was essentially used for obtaining the upper-bound. As the property also holds in our case, we may apply their arguments. Also in [4, Section 3.2], the case where $F^{h}$ is cryptographic semi-regular is considered. The results on the solving degree and the maximal degree of the Gröbner basis are heavily related to our results in this subsection.
Here we examine how two computations look like each other in early stages when we use the normal selection strategy on the choice of S-polynomials with respect to the monomial ordering $\prec^{h}$. Here we denote by $\mathcal{G}_{\rm hom}$ the set of intermediate polynomials during the computation of $G_{\rm hom}$, and denote by $\mathcal{G}$ that of $G$, namely, $\mathcal{G}$ and $\mathcal{G}_{\rm hom}$ may not be reduced and $G$ and $G_{\rm hom}$ are obtained by applying so-called ”inter-reduction” to $\mathcal{G}$ and $\mathcal{G}_{\rm hom}$, respectively.
Phase 1: Before degree fall in the computation of $G$:
The computation of $\mathcal{G}$ can simulate faithfully that of $\mathcal{G}_{\rm hom}$ until the degree of computed polynomials becomes $D-1$. Here, we call this stage an early stage and denote by $\mathcal{G}^{(e)}$ and $\mathcal{G}_{\rm hom}^{(e)}$ the set of all elements in $\mathcal{G}$ and that in $\mathcal{G}_{\rm hom}$ computed in an early stage, respectively.
In this process, we can make the following correspondence among $\mathcal{G}^{(e)}$ and that of $\mathcal{G}_{\rm hom}^{(e)}$ by carefully choosing S-polynomials and their reducers:
$\mathcal{G}_{\rm hom}^{(e)}\ni g\longleftrightarrow g^{\rm deh}\in\mathcal{G}^{(e)}.$
We can show it by induction on the degree. Consider a step where two polynomial $g_{1}$ and $g_{2}$ in $\mathcal{G}_{\rm hom}^{(e)}$ are chosen such that its S-polynomial $S(g_{1},g_{2})=t_{1}g_{1}-t_{2}g_{2}$ is of degree $d<D$, where $t_{1}$ and $t_{2}$ are terms (monomials with non-zero coefficients), $\deg(t_{1}g_{1})=\deg(t_{2}g_{2})=d$ and ${\rm LCM}({\rm LM}(g_{1}),{\rm LM}(g_{2}))={\rm LM}(t_{1}g_{1})={\rm LM}(t_{2}g_{2})$. From $S(g_{1},g_{2})$, we obtain a new element $g_{3}\neq 0$ by using some $h_{1},\ldots,h_{t}$ in $\mathcal{G}^{(e)}_{\rm hom}$ as reducers, where $h_{1},\ldots,h_{t}$ are already produced before the computation of $S(g_{1},g_{2})$. That is, $g_{3}$ can be written as
$g_{3}=t_{1}g_{1}-t_{2}g_{2}-\sum_{i=1}^{t}b_{i}h_{i}$
for some $b_{1},\ldots,b_{t}$ in $R$ such that ${\rm LM}(b_{i}h_{i})\preceq{\rm LM}(S(g_{1},g_{2}))$ for every $i$. Simultaneously, for the counter part in $\mathcal{G}^{(e)}$, two polynomial $g_{1}^{\rm deh}$ and $g_{2}^{\rm deh}$ are chosen by induction. Then we can make the obtained new element from the S-polynomial $S(g_{1}^{\rm deh},g_{2}^{\rm deh})$ equal to $g_{3}^{\rm deh}$. Indeed, as there is no degree-fall for $<D$ by Lemma 4.1.4 (since $F^{\rm top}$ is cryptographic semi-regular), we have ${\rm LM}(S(g_{1},g_{2}))={\rm LM}(S(g_{1}^{\rm deh},g_{2}^{\rm deh}))$, whence the condition ${\rm LM}(b_{i}h_{i})\preceq^{h}{\rm LM}(S(g_{1},g_{2}))$ is equivalent to ${\rm LM}(b_{i}^{\rm deh}h_{i}^{\rm deh})\preceq{\rm LM}(S(g_{1}^{\rm deh},g_{2}^{\rm deh}))$. Since $h_{1}^{\rm deh},\ldots,h_{t}^{\rm deh}$ are already computed before the computation of $S(g_{1}^{\rm deh},g_{2}^{\rm deh})$ by induction, the following expression
$g_{3}^{\rm deh}=t_{1}g_{1}^{\rm deh}-t_{2}g_{2}^{\rm deh}-\sum_{i=1}^{t}b_{i}^{\rm deh}h_{i}^{\rm deh}$
matches to the reduction process of $S(g_{1}^{\rm deh},g_{2}^{\rm deh})$. (It can be easily checked by our induction hypothesis that $g_{3}^{\rm deh}$ cannot be reduced by any element in $\mathcal{G}^{(e)}$ already computed before the computation of $S(g_{1}^{\rm deh},g_{2}^{\rm deh})$.) Here we note that, since we use the normal selection strategy, each pair $(g_{1},g_{2})$ is chosen simply by checking ${\rm LCM}({\rm LM}(g_{1}),{\rm LM}(g_{2}))$. Moreover, also by synchronizing the choice of reducers, the computation of reduction of $S(g_{1}^{\rm deh},g_{2}^{\rm deh})$ can be synchronized faithfully with that of $g_{3}$ in $G_{\rm hom}^{(e)}$ at this early stage.
Conversely, we can make the computation of $\mathcal{G}_{\rm hom}^{(e)}$ to match with that of $\mathcal{G}^{(e)}$ at an early stage in the same manner. Thus, we have ${\rm LM}(\mathcal{G}^{(e)})={\rm LM}(\mathcal{G}_{\rm hom}^{(e)})$ in this case. Of course, the reduction computation for each S-polynomial depends on the choice of reducers, and some elements might be not synchronized faithfully in actual computation. However, the set ${\rm LM}(\mathcal{G}_{\rm hom}^{(e)})$ is automatically minimal, that is, it has no element $g$ in $\mathcal{G}_{\rm hom}^{(e)}$ such that ${\rm LM}(g)$ is divisible by ${\rm LM}(g^{\prime})$ for some its another element $g^{\prime}$ in $\mathcal{G}_{\rm hom}^{(e)}$. Thus, ${\rm LM}(\mathcal{G}_{\rm hom}^{(e)})$ coincides with ${\rm LM}((G_{\rm hom})_{<D})$, that is, it does not depend on the process for the computation of $G_{\rm hom}$. Hence, we have the following:
Lemma 4.2.2.
$LM(\mathcal{G}^{(e)})$ coincides with $LM(\mathcal{G}_{\rm hom}^{(e)})={\rm LM}((G_{\rm hom})_{<D})$.
Phase 2: At the step degree $D$:
Next we investigate the computation of $G_{\rm hom}$ at the step degree $D$. In this phase, there might occur some degree fall, from which the computation process would become very complicated. Thus, to simply our investigation, we also assume to use the sugar strategy for the computation of $G$, by which the computational behaviour becomes very close to that for $G_{\rm hom}$. See [13] for details on the sugar strategy.
After the computation at the step degree $D-1$, we enter the computation at step degree $D$. In this phase, pairs of degree $D$ in $\mathcal{G}_{\rm hom}^{(e)}$ are chosen. Simultaneously, corresponding pairs in $\mathcal{G}_{\rm hom}$ of degree $D$ are chosen. (Here we continue to synchronize the computation of $\mathcal{G}^{(e)}$ and that of $\mathcal{G}^{(e)}_{\rm hom}$ as in Phase 1.) Thus, we extend the notations $\mathcal{G}^{(e)}_{\rm hom}$ and $\mathcal{G}^{(e)}$ to the step degree $D$. Let $\mathcal{G}^{(e),D}_{\rm hom}$ be the set of all elements obtained at the step degree $D$, each of which is computed from an S-polynomial $(g_{1},g_{2})$ such that $g_{1}$ and $g_{2}$ belong to $\mathcal{G}^{(e)}_{\rm hom}$ and $S(g_{1},g_{2})$ is of degree $D$. Similarly we let $\mathcal{G}^{(e),D}$ be the set of all elements in $\mathcal{G}$ obtained at the step degree $D$. We note that no element in $\mathcal{G}^{(e),D}_{\rm hom}$ is used for constructing an S-polynomial at this phase, and so for $\mathcal{G}^{(e),D}$.
Let $(g_{1},g_{2})$ be a pair in $\mathcal{G}_{\rm hom}^{(e)}$ such that its S-polynomial $S(g_{1},g_{2})$ is reduced to $g_{3}$ and ${\rm LM}(g_{3})$ is not divisible by $y$. Consider the step where $(g_{1},g_{2})$ is chosen, and simultaneously, its corresponding pair $(g_{1}^{\rm deh},g_{2}^{\rm deh})$ is also chosen. Let $g^{\prime}$ be an element computed from the corresponding S-polynomial $S(g_{1}^{\rm deh},g_{2}^{\rm deh})$. Then $g_{3}$ is obtained from $S(g_{1},g_{2})=t_{1}g_{1}-t_{2}g_{2}$ as
$g_{3}=t_{1}g_{1}-t_{2}g_{2}-\sum_{i=1}^{t}b_{i}h_{i}$
by reducers $h_{1},\ldots,h_{t}$ in $\mathcal{G}^{(e)}_{\rm hom}$. Simultaneously, $S(g_{1}^{\rm deh},g_{2}^{\rm,deh})$ can be also reduced to $g_{3}^{\rm deh}$ by reducers $h_{1}^{\rm deh},\ldots,h_{t}^{\rm deh}$;
$g_{3}^{\rm deh}=t_{1}g_{1}^{\rm deh}-t_{2}g_{2}^{\rm deh}-\sum_{i=1}^{t}b_{i}^{\rm deh}h_{i}^{\rm deh}.$
If $g_{3}^{\rm deh}$ is not reducible by any element in $\mathcal{G}^{(e)}\cup\mathcal{G}^{(e),D}$ already computed before the computation of $S(g_{1}^{\rm deh},g_{2}^{\rm deh})$, then ${\rm LM}(g_{3}^{\rm{deh}})={\rm LM}(g^{\prime})$. So, there is still a correspondence, and $\langle{\rm LM}(\mathcal{G}^{(e)}\cup\mathcal{G}^{(e),D})\rangle$ contains ${\rm LM}(g_{3}^{\rm deh})$. Otherwise, ${\rm LM}(g_{3}^{\rm{deh}})$ is divisible by ${\rm LM}(g^{\prime\prime})$ for some $g^{\prime\prime}$ already computed elements in $\mathcal{G}^{(e)}\cup\mathcal{G}^{(e),D}$ at the step degree $D$. This implies that $\langle{\rm LM}(\mathcal{G}^{(e)}\cup\mathcal{G}^{(e),D})\rangle$ contains ${\rm LM}(g_{3}^{\rm deh})$, which holds for any pair $(g_{1},g_{2})$ generated at the step degree $D$. Hence, $\langle{\rm LM}(\mathcal{G}^{(e)}\cup\mathcal{G}^{(e),D})\rangle$ includes ${\rm LM}((G_{\rm hom})_{\leq D})\cap R_{D}$. Therefore, $\langle{\rm LM}(\mathcal{G}^{(e)}\cup\mathcal{G}^{(e),D})\rangle$ contains all monomials of degree $D$ in $X$, since $\langle{\rm LM}((G_{\rm hom})_{\leq D})\rangle_{R^{\prime}}\cap R_{D}=R_{D}$ by Lemma 4.1.5. Thus, we have the following lemma.
Lemma 4.2.3.
$\langle{\rm LM}(\mathcal{G}^{(e)}\cup\mathcal{G}^{(e),D})\rangle$ contains all monomials in $X$ of degree $\geq D$.
Solving degree of $F$ as the highest step degree:
Here we show an upper-bound on the highest step degree appeared in the computation of $G$ with respect to the DRL ordering by a Buchberger-like algorithm $\mathcal{A}$ based on S-polynomials with the normal strategy and the sugar strategy. We note that, in [30, Lemma 4.2.4], we restart the computation of the Gröbner basis of $F$ from $H=\{g|_{y=1}:g\in(G_{\rm hom})_{\leq D}\}$. However, here we do not need $(G_{\rm hom})_{\leq D}$. We refer to [8, Remark 15] for another proof of ${\rm max.GB.}\deg_{\prec}(F)\leq D$.
Lemma 4.2.4 (cf. [30, Lemma 4]).
Assume that $D\geq{\rm max}\{\deg(f):f\in F\}$, and that $\prec$ is a DRL ordering on the set of monomials in $R$. Then, it follows that ${\rm max.GB.}\deg_{\prec}(F)\leq D$. Moreover, there exists a Buchberger-like algorithm $\mathcal{A}$ with normal strategy such that
${\rm sd}_{\prec}^{\mathcal{A}}(F)\leq{\color[rgb]{0,0,0}2D-1},$
and
${\rm sd}_{\prec}^{\mathcal{A}}(F)\leq{\color[rgb]{0,0,0}2D-2}.$
in the strict sense (see (I) in Subsection 2.2 for details on the definition of these solving degrees). Namely, the maximal degree of S-polynomials generated during the execution of $\mathcal{A}$ is bounded by $22D-$ 2.
Remark 4.2.5.
We refer to [8, Remark 15] for another proof of ${\rm max.GB.}\deg_{\prec}({F})\leq D$. We also note that, if $D=d_{\rm reg}({F}^{\rm top})<\infty$, Lemma 4.2.3 and Lemma 4.2.4 hold without the assumption that ${F^{\rm top}}$ is cryptographic semi-regular.
Remark 4.2.6 (cf. [30, Section 5.2]).
As to the computation of $G_{\rm hom}$, we have a result similar to Lemma 4.2.4. Since $\langle{\rm LM}(G_{\rm hom})_{\leq D}\rangle$ contains all monomials in $X$ of degree $D$, for any polynomial $g$ generated in the middle of the computation of $G_{\rm hom}$ the degree of the $X$-part of ${\rm LM}(g)$ is less than $D$. Because $g$ is reduced by $(G_{\rm hom})_{\leq D}$. Thus, letting ${\mathcal{U}}$ be the set of all polynomials generated during the computation of $G_{\rm hom}$, we have
$\{\mbox{The $X$-part of }{\rm LM}(g):g\in{\mathcal{U}}\}\subset\{x_{1}^{e_{1}}\cdots x_{n}^{e_{n}}:e_{1}+\cdots+e_{n}\leq D\}.$
As different $g,g^{\prime}\in{\mathcal{U}}$ can not have the same $X$-part in their leading terms, the size #$\#{\mathcal{U}}$ is upper-bounded by the number of monomials in $X$ of degree not greater than $D$, that is $\binom{n+D}{n}$. By using the $F_{5}$ algorithm or its efficient variant, under an assumption that every unnecessary S-polynomial can be avoided, the number of computed S-polynomials during the computation of $G_{\rm hom}$ coincides with the number #$\#{\mathcal{U}}$ and is upper-bounded by $\binom{n+D}{n}$.
We review a simple example shown in [30, Example 1] and examine the correspondences discussed in this and the previous subsections.
Example 4.2.7.
We give a simple example. Let $p=73$, $K={\mathbb{F}}_{p}$, and
$f_{1}$ $=$ $x_{1}^{2}+3x_{1}x_{2}+x_{2}^{2}-2x_{1}x_{3}-2x_{2}x_{3}+x_{3}^{2}-x_{1}-2x_{2}+x_{3},$
$f_{2}$ $=$ $4x_{1}^{2}+3x_{1}x_{2}+4x_{1}x_{3}+x_{3}^{2}-2x_{1}-x_{2}+2x_{3},$
$f_{3}$ $=$ $3x_{1}^{2}+9x_{2}^{2}-6x_{2}x_{3}+x_{3}^{2}-x_{1}+x_{2}-x_{3},$
$f_{4}$ $=$ $x_{1}^{2}-6x_{1}x_{2}+9x_{2}^{2}+2x_{1}x_{3}-6x_{2}x_{3}+2x_{3}^{2}-2x_{1}+x_{2}.$
Then, $d_{1}=d_{2}=d_{3}=d_{4}=2$. As their top parts (maximal total degree parts) are
$f_{1}^{\rm top}$ $=$ $x_{1}^{2}+3x_{1}x_{2}+x_{2}^{2}-2x_{1}x_{3}-2x_{2}x_{3}+x_{3}^{2},$
$f_{2}^{\rm top}$ $=$ $4x_{1}^{2}+3x_{1}x_{2}+4x_{1}x_{3}+x_{3}^{2},$
$f_{3}^{\rm top}$ $=$ $3x_{1}^{2}+9x_{2}^{2}-6x_{2}x_{3}+x_{3}^{2},$
$f_{4}^{\rm top}$ $=$ $x_{1}^{2}-6x_{1}x_{2}+9x_{2}^{2}+2x_{1}x_{3}-6x_{3}x_{2}+2x_{3}^{2},$
one can verify that $\bm{F}^{\rm top}$ is cryptographic semi-regular (and furthermore, $\bm{F}^{\rm top}$ is semi-regular). Then its degree of regularity is equal to $3$. Indeed, the reduced Gröbner basis $G_{\rm top}$ of the ideal $\langle F^{\rm top}\rangle$ with respect to the DRL ordering $x_{1}\succ x_{2}\succ x_{3}$ is
$\{\underline{x_{2}x_{3}^{2}},\ \underline{x_{3}^{3}},\ \underline{x_{1}^{2}}+68x_{2}x_{3}+55x_{3}^{2},\ \underline{x_{1}x_{2}}+27x_{2}x_{3}+29x_{3}^{2},\ \underline{x_{2}^{2}}+x_{2}x_{3}+71x_{3}^{2},\ \underline{x_{1}x_{3}}+3x_{2}x_{3}+33x_{3}^{2}\}.$
Then its leading monomials are $x_{2}x_{3}^{2},x_{3}^{3},x_{1}^{2},x_{1}x_{2},x_{2}^{2},x_{1}x_{3}$ and its Hilbert-Poincaré series satisfies
${\rm HS}_{R/\langle F^{\rm top}\rangle}(z)=2z^{2}+3z+1=\left(\frac{(1-z^{2})^{4}}{(1-z)^{3}}\bmod{z^{3}}\right),$
whence the degree of regularity of $\langle F^{\rm top}\rangle$ is 3.
On the other hand, the reduced Gröbner basis $G_{\rm hom}$ of the ideal $\langle F^{h}\rangle$ with respect to the DRL ordering $x_{1}\succ x_{2}\succ x_{3}\succ y$ is
$\{\underline{x_{1}y^{3}},\ \underline{x_{2}y^{3}},\ \underline{x_{3}y^{3}},\ \underline{x_{2}x_{3}^{2}}+60x_{1}y^{2}+22x_{2}y^{2}+39x_{3}y^{2},$
$\underline{x_{3}^{3}}+72x_{1}y^{2}+14x_{2}y^{2}+56x_{3}y^{2},\ \underline{x_{2}x_{3}y}+16x_{1}y^{2}+55x_{2}y^{2}+38x_{3}y^{2},$
$\underline{x_{3}^{2}y}+72x_{1}y^{2}+66x_{2}y^{2}+70x_{3}y^{2},\ \underline{x_{1}^{2}}+68x_{2}x_{3}+55x_{3}^{2}+72x_{1}y+40x_{2}y+14x_{3}y,$
$\underline{x_{1}x_{2}}+27x_{2}x_{3}+29x_{3}^{2}+20x_{1}y+37x_{2}y+12x_{3}y,$
$\underline{x_{2}^{2}}+x_{2}x_{3}+71x_{3}^{2}+57x_{1}y+3x_{2}y+52x_{3}y,$
$\underline{x_{1}x_{3}}+3x_{2}x_{3}+33x_{3}^{2}+22x_{1}y+5x_{2}y+14x_{3}y\}$
and its leading monomials are $x_{1}y^{3},x_{2}y^{3},x_{3}y^{3},x_{2}x_{3}^{2},x_{3}^{3},x_{2}x_{3}y,x_{3}^{2}y,x_{1}^{2},x_{1}x_{2},x_{2}^{2},x_{1}x_{3}$. Then the Hilbert-Poincaré series of $R^{\prime}/\langle F^{h}\rangle$ satisfies
$\left({\rm HS}_{R^{\prime}/\langle F^{h}\rangle}(z)\bmod{z^{3}}\right)=\left(6z^{2}+4z+1\bmod{z^{3}}\right)=\left(\frac{(1-z^{2})^{4}}{(1-z)^{4}}\bmod{z^{3}}\right).$
We note that ${\rm HF}_{R^{\prime}/\langle F^{h}\rangle}(3)=4$ and ${\rm HF}_{R^{\prime}/\langle F^{h}\rangle}(4)=1$. We can also examine the correspondence ${\rm LM}(G_{\rm hom})_{<D}={\rm LM}(G_{\rm top})_{<D}$ and, for $g\in G_{\rm hom}$, if ${\rm LM}(g)$ is divided by $y$, then $\deg(g)\geq D=3$. Thus, any degree-fall cannot occur at degree less than $3=D$.
Finally, we examine the correspondence between $\mathcal{G}^{(e)}\cup\mathcal{G}^{(e),D}$ and $(G_{\rm hom})_{\leq D}$. The reduced Gröbner basis of $\langle F\rangle$ with respect to $\prec$ is $\{x,y,z\}$ and we can examine that ${\rm LM}(\mathcal{G}^{(e)})$ coincides with ${\rm LM}(G_{\rm hom})_{<3}$. Because we have the following $\mathcal{G}$ without inter-reduction (see the paragraph just after Remark 4.2.1 for the definition of $\mathcal{G}$);
$\{\underline{x_{1}^{2}}+3x_{1}x_{2}+x_{2}^{2}+71x_{1}x_{3}+71x_{2}x_{3}+x_{3}^{2}+72x_{1}+71x_{2}+x_{3},$
$\underline{x_{1}x_{2}}+41x_{2}^{2}+23x_{1}x_{3}+64x_{2}x_{3}+49x_{3}^{2}+16x_{1}+56x_{2}+57x_{3},$
$\underline{x_{2}^{2}}+{14x_{1}x_{3}}+43x_{2}x_{3}+22x_{3}^{2}+29x_{3},\ \underline{x_{1}x_{3}}+3x_{2}x_{3}+33x_{3}^{2}+22x_{1}+5x_{2}+14x_{3},$
$\underline{x_{2}x_{3}^{2}}+41x_{3}^{3}+5x_{2}x_{3}+35x_{3}^{2}+64x_{1}+42x_{2}+11x_{3},\ \underline{x_{3}^{3}}+35x_{3}^{2}+37x_{1}+61x_{2}+24x_{3},$
$\underline{x_{3}x_{2}}+13x_{3}^{2}+3x_{1}+37x_{2}+72x_{3},\ \underline{x_{3}^{2}}+72x_{1}+66x_{2}+70x_{3},$
$\underline{x_{1}}+61x_{2}+51x_{3},\ \underline{x_{2}}+70x_{3},\ \underline{x_{3}}\},$
and ${\rm LM}(\mathcal{G}^{(e)})=\{x_{1}^{2},x_{1}x_{2},x_{2}^{2},x_{1}x_{3}\}$. Moreover, ${\rm LM}({G_{\rm hom}})_{D}$ coincides with ${\rm LM}(\mathcal{G}^{(e),D})$, as it is $\{x_{2}x_{3}^{2},x_{3}^{2},x_{2}x_{3},x_{3}^{2}\}$. We note that we have removed $f_{2},f_{3},f_{4}$ from $\mathcal{G}$ as they have the same LM as $f_{1}$. Interestingly, in this case, we can see that the whole ${\rm LM}(\mathcal{G})$ corresponds to ${\rm LM}(G_{\rm hom})$.
4.3 Experimental observation and a variant of Fröberg’s conjecture
In this subsection, we observe actual behavior of Gröbner basis computation from a part of our experimental results. For experiments, we used Magma V2.25-3 [5]. In particular, the built-in functions Dimension, HilbertSeries, and GroebnerBasis were applied. In our experiments, given $n$, $m$, $(d_{1},\ldots,d_{m})$, and an odd prime $q$ with $n<m$ and $d_{i}\geq 2$ (for the case where $n=m$, see Remark 4.3.3 below), we generate inhomogeneous polynomials $f_{1},\ldots,f_{m}$ in $R=\mathbb{F}_{q}[x_{1},\ldots,x_{n}]$ of degrees $d_{1},\ldots,d_{m}$ whose constant terms are all zero (as in Example 4.2.7), where coefficients are chosen uniformly at random. Note that, for any sequence $\bm{F}=(f_{1},\ldots,f_{m})$ generated in this way, the homogenization $F^{h}$ with $F=\{f_{1},\ldots,f_{m}\}$ always has a non-trivial root $(0,\ldots,0,1)$: Any sequence of $m$ polynomials (in $R$) that has at least one root is transformed by a linear coordinate change into a polynomial sequence $(f_{1},\ldots,f_{m})\in R^{m}$ such that each $f_{i}$ has no constant term. Moreover, $\bm{F}^{\rm top}=(f_{1}^{\rm top},\ldots,f_{m}^{\rm top})$ is expected to be cryptographic semi-regular (in fact, semi-regular) with high probability. Therefore, our setting causes no loss of generality.
As an experimental result, we heuristically find an upper-bound
${\deg\left(\left[\frac{\prod_{i=1}^{m}(1-z^{d_{i}})}{(1-z)^{n+1}}\right]\right)}+1$ (4.3.1)
on the maximal Gröbner basis degree $\mathrm{max.GB.deg}_{\prec^{h}}(F^{h})$, which is sharper than our upper-bound provided in Theorem 2 (1). This sharper bound comes from a property that $\bm{F}^{h}$ is generalized cryptographic semi-regular (i.e., $D^{\prime}$-regular), where we set $D^{\prime}:=\widetilde{d}_{\rm reg}(F^{h})$: We confirmed in our experiments that $\bm{F}^{h}$ satisfies this property in most cases. Here we use the following lemma (the proof is straightfoward):
Lemma 4.3.1.
Let $\bm{F}=(f_{1},\ldots,f_{m})\in(R\smallsetminus K)^{m}$ be a sequence of not necessarily homogeneous polynomial. Assume that $n<m$. Then we have the following:
  1. 1.
    If ${F}^{h}$ has at least one solution and if $\bm{F}^{h}$ is $D^{\prime}$-regular (and thus $D^{\prime}<\infty$), then $D^{\prime}$ is upper-bounded by (4.3.1).
  2. 2.
    If $\bm{F}^{\rm top}$ is cryptographic semi-regular, then $D:=d_{\rm reg}(F^{\rm top})$ is also upper-bounded by (4.3.1) (this also holds even if $n=m$).
Supposing both the assumptions in this lemma, we have
$\mathrm{max.GB.deg}_{\prec^{h}}(F^{h})\leq\max\{D,D^{\prime}\}\leq{\deg\left(\left[\frac{\prod_{i=1}^{m}(1-z^{d_{i}})}{(1-z)^{n+1}}\right]\right)}+1$
by Proposition 2.3.3. We compare exact values for several bounds, for $n=9$ and $10$ with some conditions on $m$ and $(d_{1},\ldots,d_{m})$ in Tables 14, where we set $D:=d_{\rm reg}(F^{\rm top})$ for the case where $\bm{F}^{\rm top}$ is cryptographic semi-regular.
Table 1: Exact values for several upper-bounds on $\mathrm{max.GB.deg}_{\prec^{h}}(F^{h})$ in the case where $n=9$, $n+1\leq m\leq 2n$, and $d_{1}=\cdots=\cdots=d_{m}=2$.
$m$ $10$ $11$ $12$ $13$ $14$ $15$ $16$ $17$ $18$
Lazard’s bound (Theorem 2.2.1) $11$ $11$ $11$ $11$ $11$ $11$ $11$ $11$ $11$
Our bound in Theorem 2 (1) $11$ $11$ $11$ $11$ $11$ $11$ $11$ $11$ $11$
Heuristic bound (4.3.1) (Conjecture 4.3.4) $11$ 66{6}6 66{6}6 55{5}5 55{5}5 44{4}4 44{4}4 44{4}4 44{4}4
$D$ for semi-regular ${F}^{\rm top}$ $6$ $5$ $5$ $4$ $4$ $4$ $4$ $4$ $4$
$12D-$ 1 (Theorem 3 (4)) $1${11}11 $9$ $9$ $7$ $7$ $7$ $7$ $7$ $7$
Table 2: Exact values for several upper-bounds on $\mathrm{max.GB.deg}_{\prec^{h}}(F^{h})$ in the case where $n=10$, $n+1\leq m\leq 2n$, and $d_{1}=\cdots=d_{m}=2$.
$m$ $11$ $12$ $13$ $14$ $15$ $16$ $17$ $18$ $19$ $20$
Lazard’s bound (Theorem 2.2.1) $12$ $12$ $12$ $12$ $12$ $12$ $12$ $12$ $12$ $12$
Our bound in Theorem 2 (1) $12$ $12$ $12$ $12$ $12$ $12$ $12$ $12$ $12$ $12$
Heuristic bound (4.3.1) (Conjecture 4.3.4) $12$ 77{7}7 66{6}6 55{5}5 55{5}5 55{5}5 55{5}5 44{4}4 44{4}4 44{4}4
$D$ for semi-regular $\bm{F}^{\rm top}$ $6$ $6$ $5$ $5$ $4$ $4$ $4$ $4$ $4$ $4$
$12D-$ 1 (Theorem 3 (4)) $1${11}11 $11$ $9$ $9$ $7$ $7$ $7$ $7$ $7$ $7$
Table 3: Exact values for several upper-bounds on $\mathrm{max.GB.deg}_{\prec^{h}}(F^{h})$ in the case where $n=9$, $n+1\leq m\leq 2n$, $d_{1}=\cdots=d_{n}=3$, and $d_{n+1}=\cdots=d_{m}=2$.
$m$ $10$ $11$ $12$ $13$ $14$ $15$ $16$ $17$ $18$
Lazard’s bound (Theorem 2.2.1) $20$ $20$ $20$ $20$ $20$ $20$ $20$ $20$ $20$
Our bound in Theorem 2 (1) $20$ $19$ $18$ $17$ $16$ $15$ $14$ $13$ $12$
Heuristic bound (4.3.1) (Conjecture 4.3.4) $20$ $1${11}11 99{9}9 88{8}8 77{7}7 77{7}7 66{6}6 66{6}6 55{5}5
$D$ for semi-regular ${F}^{\rm top}$ $10$ $9$ $8$ $7$ $6$ $6$ $6$ $5$ $5$
$12D-$ 1 (Theorem 3 (4)) 1919{19}19 $17$ $15$ $13$ $11$ $11$ $9$ $9$ $9$
Table 4: Exact values for several upper-bounds on $\mathrm{max.GB.deg}_{\prec^{h}}(F^{h})$ in the case where $n=10$, $n+1\leq m\leq 2n$, $d_{1}=\cdots=d_{n}=3$, and $d_{n+1}=\cdots=d_{m}=2$.
$m$ $11$ $12$ $13$ $14$ $15$ $16$ $17$ $18$ $19$ $20$
Lazard’s bound (Theorem 2.2.1) $22$ $22$ $22$ $22$ $22$ $22$ $22$ $22$ $22$ $22$
Our bound in Theorem 2 (1) $22$ $21$ $20$ $19$ $18$ $17$ $16$ $15$ $14$ $13$
Heuristic bound (4.3.1) (Conjecture 4.3.4) $22$ 1212{12}12 1010{10}10 99{9}9 88{8}8 77{7}7 77{7}7 66{6}6 66{6}6 66{6}6
$D$ for semi-regular $\bm{F}^{\rm top}$ $11$ $10$ $9$ $8$ $7$ $6$ $6$ $6$ $5$ $5$
$12D-$ 1 (Theorem 3 (4)) 2121{21}21 $19$ $17$ $15$ $13$ $11$ $11$ $11$ $9$ $9$
Remark 4.3.2.
Note that, if $m=n+1$, then it follows from $\frac{\prod_{i=1}^{m}(1-z^{d_{i}})}{(1-z)^{n+1}}=\prod_{i=1}^{n+1}(1+z+\cdots+z^{d_{i}-1})$ that the bound (4.3.1) is equal to $\sum_{j=1}^{n+1}(d_{j}-1)+1$, which is equal to Lazard’s bound. As a more particular case, if $d_{i}=2$ for all $i$, then it is equal to $n+2$. On the other hand, recall from [4, Theorem 4.1] that $D=\lfloor(n+1)/2\rfloor+1$, and thus $12D-$ 1 is equal to $n+2$ if $n$ is odd and to $n+1$ if $n$ is even.
Remark 4.3.3.
If $m=n$, then it follows from $\frac{\prod_{i=1}^{m}(1-z^{d_{i}})}{(1-z)^{n}}=\prod_{i=1}^{n}(1+z+\cdots+z^{d_{i}-1})$ that $D=\sum_{j=1}^{n}(d_{j}-1)+1$, and we have
$\frac{\prod_{i=1}^{m}(1-z^{d_{i}})}{(1-z)^{n+1}}=\frac{\prod_{i=1}^{n}(1-z^{d_{i}})}{(1-z)^{n}}\cdot(1+z+z^{2}+\cdots),$
whose degree-$d$ coefficients are equal to a constant for $d\geq D-1$. In this case, the $D^{\prime}$-regularity of $F^{h}$ implies that $D^{\prime}\leq D$, whence $D^{\prime}\in\{D-1,D\}$ and $\mathrm{max.GB.deg}_{\prec^{h}}(F^{h})\leq D$ by Proposition 2.3.3. Note that the equality $\mathrm{max.GB.deg}_{\prec^{h}}(F^{h})=D$ holds if $\langle\mathrm{LM}(\bm{F}^{h})\rangle$ is a weakly reverse lexicographic ideal.
Based on our experiments, we here raise the following conjecture:
Conjecture 4.3.4.
Let $K$ be an infinite field, and let $R=K[x_{1},\ldots,x_{n}]$. Let $d_{1},\ldots,d_{m}$ are integers larger than $1$, and let $f_{1},\ldots,f_{m}$ be polynomials in $R$ of degrees $d_{1},\ldots,d_{m}$ such that each $f_{i}$ has no constant term. Then, for given $K$, $n$, $m$, and $(d_{1},\ldots,d_{m})$, the property that $\bm{F}^{h}=(f_{1}^{h},\ldots,f_{m}^{h})$ is generalized cryptographic semi-regular (i.e., $D^{\prime}$-regular) is generic, where $D^{\prime}=\widetilde{d}_{\rm reg}(F^{h})$ is the generalized degree of regularity of $\bm{F}^{h}$ defined in Definition 2.3.1.
Conjecture 4.3.4 can be viewed as Fröberg’s conjecture [21]: A generic sequence of homogeneous polynomials is $D$-regular, where $D$ is the degree of regularity of the sequence.
As a consequence of Conjecture 4.3.4, if $F$ is generated in the way described in the beginning of this subsection, we may expect the following properties:
  1. 1.
    As we described as above, the solving degree $\mathrm{sd}^{\rm mac}_{\prec^{h}}(F^{h})$ (which is equal to the solving degree $\mathrm{sd}^{\rm mac}_{\prec}(F)$ if $\prec$ is a DRL order, see Subsection 2.2) is also upper-bounded by $D^{\prime}$, which can be quite smaller than $12D-$ 1.
  2. 2.
    It follows from Theorem 2.1.4 that ${\rm HS}_{A^{\prime}}(z)\equiv\frac{\prod_{j=1}^{m}(1-z^{d_{j}})}{(1-z)^{n}}\pmod{z^{D^{\prime}}}$ for $A^{\prime}=R^{\prime}/\langle F^{h}\rangle$ with $R^{\prime}=K[x_{1},\ldots,x_{n},y]$. Hence, by the definition of $D^{\prime}$, the Hilbert series of $A^{\prime}$ is computed as follows:
    ${\rm HS}_{A^{\prime}}(z)=\left(\frac{\prod_{j=1}^{m}(1-z^{d_{j}})}{(1-z)^{n}}\bmod{z^{D^{\prime}}}\right)+\sum_{d=D^{\prime}}^{\infty}N_{F^{h}}z^{d},$ (4.3.2)
    where $N_{F^{h}}$ is the number of projective roots of $F^{h}$ counted with multiplicity. This implies that the Hilbert driven algorithm can be effectively applied to the Gröbner basis computation of $F^{h}$, from which a Gröbner basis of $F$ is easily obtained.
  3. 3.
    As to the shape of the Hilbert function ${\rm HF}_{A^{\prime}}(z)$ of $A^{\prime}$, its unimodality and symmetry (up to degree $D^{\prime}$) can be easily examined by the formula (4.3.2).
Here, we also raise interesting questions:
Question 4.3.5.
  1. 1.
    Does Fröberg’s conjecture imply Conjecture 4.3.4? (Or, does the converse hold?)
  2. 2.
    If $R/\langle F^{\rm top}\rangle$ is Artinian and if $\bm{F}^{\rm top}$ is cryptographic semi-regular, does one of the following conditions hold generically?
    • (A)
      $\bm{F}^{h}$ is generalized cryptographic semi-regular (i.e., $D^{\prime}$-regular).
    • (B)
      $\langle\mathrm{LM}(\bm{F}^{h})\rangle$ is a weakly reverse lexicographic ideal. (Cf. Moreno-Socías conjecture [35].)
  3. 3.
    Are the conditions (A) and (B) are equivalent to each other?
We refer to [36] for several conjectures equivalent to Fröberg’s conjecture. It is one of our future works to give answers to these questions.
Final remark for security analysis in cryptography
The assumption that $\bm{F}^{h}$ is generalized cryptographic semi-regular (i.e., $D^{\prime}$-regular) could be useful to estimate the security of multivariate cryptosystems and signature schemes (or algebraic attacks based on Gröbner basis computation), see e.g., [22, Subsection 2.3], where the authors of [22] assume that the bound (4.3.1) gives a degree bound of the XL algorithm [11]. In fact, the bound (4.3.1) has been sometimes used in the cryptographic community without assuming the $D^{\prime}$-regularity of $\bm{F}^{h}$, see [22, Subsection 2.3] (in particular [22, Remark 1]) for details. Supposing the $D^{\prime}$-regularity of the homogenization of a target polynomial system, one could give a mathematically rigid explanation about the complexity of computing the Gröbner basis of the system.
Acknowledgement
The contents of this paper will be presented at Effective Methods in Algebraic Geometry (MEGA 2024). The authors thank the anonymous referees assigned by the organization of this conference for their comments and suggestions. The authors are grateful to Yuta Kambe and Shuhei Nakamura for helpful comments. They are also thankful to Masayuki Noro for providing computational tools for our computational experiments. This work was supported by JSPS Grant-in-Aid for Young Scientists 20K14301 and 23K12949, JSPS Grant-in-Aid for Scientific Research (C) 21K03377, and JST CREST Grant Number JPMJCR2113.
Appendix A Koszul complex and homogenization
A.1 Koszul complex
Let $f_{1},\ldots,f_{m}\in R$ be homogeneous polynomials of positive degrees $d_{1},\ldots,d_{m}$ respectively, and put $d_{j_{1}\cdots j_{i}}:=\sum_{k=1}^{i}d_{j_{k}}$. For each $0\leq i\leq m$, we define a free $R$-module of rank $\binom{m}{i}$
$K_{i}(f_{1},\ldots,f_{m}):=\left\{\begin{array}[]{cl}\displaystyle\bigoplus_{1\leq j_{1}<\cdots<j_{i}\leq m}R({-d_{j_{1}\cdots j_{i}}})\mathbf{e}_{j_{1}\cdots j_{i}}&(i\geq 1)\\ R&(i=0),\end{array}\right.$
where $\mathbf{e}_{j_{1}\cdots j_{i}}$ is a standard basis. We also define a graded homomorphism
$\varphi_{i}:K_{i}(f_{1},\ldots,f_{m})\longrightarrow K_{i-1}(f_{1},\ldots,f_{m})$
of degree 00 by
$\varphi_{i}(\mathbf{e}_{j_{1}\cdots j_{i}}):=\sum_{k=1}^{i}(-1)^{k-1}f_{j_{k}}\mathbf{e}_{j_{1}\cdots\hat{j_{k}}\cdots j_{i}}.$
Here, $\hat{j_{k}}$ means to omit $j_{k}$. For example, we have $\mathbf{e}_{1\hat{2}3}=\mathbf{e}_{13}$. To simplify the notation, we set $K_{i}:=K_{i}(f_{1},\ldots,f_{m})$. Then,
$K_{\bullet}:0\to K_{m}\xrightarrow{\varphi_{m}}\cdots\xrightarrow{\varphi_{3}}K_{2}\xrightarrow{\varphi_{2}}K_{1}\xrightarrow{\varphi_{1}}K_{0}\to 0$ (A.1.1)
is a complex of graded free $R$-modules, and we call it the Koszul complex on $(f_{1},\ldots,f_{m})$. The $i$-th homology group of $K_{\bullet}$ is given by
$H_{i}(K_{\bullet})=\mathrm{Ker}(\varphi_{i})/\mathrm{Im}(\varphi_{i+1}).$
In particular, we have
$H_{0}(K_{\bullet})=R/\langle f_{1},\ldots,f_{m}\rangle_{R}.$
We also note that $H_{m}(K_{\bullet})=0$, since $\varphi_{m}$ is clearly injective by definition. The kernel and the image of a graded homomorphism are both graded submodules in general, so that $\mathrm{Ker}(\varphi_{i})$ and $\mathrm{Im}(\varphi_{i+1})$ are graded $R$-modules, and so is the quotient module $H_{i}(K_{\bullet})$ (and each homogeneous part is finite-dimensional $K$-vector space). In the following, we denote by $H_{i}(K_{\bullet})_{d}$ the degree-$d$ homogeneous part of $H_{i}(K_{\bullet})$.
Note that $\mathrm{Ker}(\varphi_{1})=\mathrm{syz}(f_{1},\ldots,f_{m})$ (the right hand side is the module of syzygies), and that $\mathrm{Im}(\varphi_{2})\subset K_{1}=\bigoplus_{j=1}^{m}R(-d_{j})\mathbf{e}_{j}$ is generated by
$\{\mathbf{t}_{i,j}:=f_{i}\mathbf{e}_{j}-f_{j}\mathbf{e}_{i}:1\leq i<j\leq m\}.$
Hence, putting
$\mathrm{tsyz}(f_{1},\ldots,f_{m}):=\langle\mathbf{t}_{i,j}:1\leq i<j\leq m\rangle_{R},$
we have
$H_{1}(K_{\bullet})=\mathrm{syz}(f_{1},\ldots,f_{m})/\mathrm{tsyz}(f_{1},\ldots,f_{m}).$ (A.1.2)
Definition A.1.1 (Trivial syzygies).
With notation as above, we call each generator $\mathbf{t}_{i,j}$ (or each element of $\mathrm{tsyz}(f_{1},\ldots,f_{m})$) a trivial syzygy for $(f_{1},\ldots,f_{m})$. We also call $\mathrm{tsyz}(f_{1},\ldots,f_{m})$ the module of trivial syzygies.
A.2 Homogenization of polynomials and monomial orders
We here recall the notion of homogenization; see [29, Chapter 4] for details. Let $R=K[x_{1},\ldots,x_{n}]$ be the polynomial ring of $n$ variables over a field $K$, and $\mathcal{T}$ the set of all monomials in $x_{1},\ldots,x_{n}$. Put $R^{\prime}=R[y]$ for an extra variable $y$.
  1. (1)
    For a non-homogeneous and non-zero polynomial $f=\sum_{{t}\in\mathcal{T}}c_{t}{t}$ in $R$ with $c_{t}\in K$, its homogenization $f^{h}$ is defined, by introducing an extra variable $y$, as
    $f^{h}=\sum_{{t}\in\mathcal{T}}c_{t}{t}{y^{\deg(f)-{\color[rgb]{0,0,0}\deg(t)}}}\in R^{\prime}=R[y].$
    Thus $f^{h}$ is a homogeneous polynomial in $R^{\prime}$ with total degree $d=\deg(f)$. Also for a set $F$ (or a sequence ${F}=(f_{1},\ldots,f_{m})\in R^{m}$) of non-zero polynomials, its homogenization $F^{h}$ (or ${F}^{h}$) is defined as $F^{h}=\{f^{h}\;|\;f\in F\}$ (or ${F}^{h}=(f_{1}^{h},\ldots,f_{m}^{h})\in(R^{\prime})^{m}$).
  2. (2)
    Conversely, for a homogeneous polynomial $h$ in $R^{\prime}$, its dehomogenization $h^{\rm deh}$ is defined by substituting $y$ with $1$, that is, $h^{\rm deh}=h(x_{1},\ldots,x_{n},1)$ (it is also denoted by $h|_{y=1}$). For a set $H$ of homogeneous polynomials in $R^{\prime}$, its dehomogenization $H^{\rm deh}$ (or $H|_{y=1}$) is defined as $H^{\rm deh}=\{h^{\rm deh}:h\in H\}$. We also apply the dehomogenization to sequences of polynomials.
  3. (3)
    For an ideal $I$ of $R$, its homogenization $I^{h}$, as an ideal, is defined as $\langle I^{h}\rangle_{R^{\prime}}$. We remark that, for a set $F$ of polynomials in $R$, we have $\langle F^{h}\rangle_{R^{\prime}}\subset I^{h}$ with $I=\langle F\rangle_{R}$, and the equality does not hold in general.
  4. (4)
    For a homogeneous ideal $J$ in $R^{\prime}$, its dehomogenization $J^{\rm deh}$, as a set, is an ideal of $R$. We note that if a homogeneous ideal $J$ is generated by $H$, then $J^{\rm deh}=\langle H^{\rm deh}\rangle_{R}$ and for an ideal $I$ of $R$, we have $(I^{h})^{\rm deh}=I$.
  5. (5)
    For a monomial ordering $\prec$ on the set of monomials $\mathcal{T}$ in $X$, its homogenization $\prec_{h}$ on the set of monomials $\mathcal{T}^{h}$ in $x_{1},\ldots,x_{n},y$ is defined as follows: For two monomials $X^{\alpha}y^{a}$ and $X^{\beta}y^{b}$ in $\mathcal{T}^{h}$, we say $X^{\alpha}y^{a}\prec_{h}X^{\beta}y^{b}$ if and only if one of the following holds:
    • (i)
      $a+|\alpha|<b+|\beta|$, or
    • (ii)
      $a+|\alpha|=b+|\beta|$ and $X^{\alpha}\prec X^{\beta}$,
    where $\alpha=(\alpha_{1},\ldots,\alpha_{n})\in\mathbb{Z}_{\geq 0}^{n}$ and $|\alpha|=\alpha_{1}+\cdots+\alpha_{n}$, and where $X^{\alpha}$ denotes $x_{1}^{\alpha_{1}}\cdots x_{n}^{\alpha_{n}}$. Here, for a monomial $X^{\alpha}{\color[rgb]{0,0,0}y^{a}}$, we call $X^{\alpha}$ and ${\color[rgb]{0,0,0}y^{a}}$ the $X$-part and the $y$-part, respectively. If a monomial ordering $\prec$ is graded, that is, it first compares the total degrees, the restriction $\prec_{h}|_{\mathcal{T}}$ of $\prec_{h}$ on $\mathcal{T}$ coincides with $\prec$.
It is well-known that, for a Gröbner basis $H$ of $\langle F^{h}\rangle$ with respect to $\prec^{h}$, its dehomogenization $H^{\rm deh}=\{h^{\rm deh}:h\in H\}$ is also a Gröbner basis of $\langle F\rangle$ with respect to $\prec$ if $\prec$ is graded. Moreover, we have $\langle F\rangle^{h}=(\langle F^{h}\rangle:\langle y\rangle^{\infty})=(\langle F^{h}\rangle:\langle y^{k}\rangle)$ for some integer $k$, where $(\langle F^{h}\rangle:\langle y^{k}\rangle)$ is the ideal quotient of $\langle F^{h}\rangle$ by $\langle y^{k}\rangle$, namely $\{f\in R^{\prime}:f\langle y^{k}\rangle\subset\langle F^{h}\rangle\}$ see [29, Corollary 4.3.8].

### v3

@@
 Determining the complexity of computing Gröbner bases is an important problem both in theory and in practice, and for that the solving degree plays a key role.
-In this paper, we study the solving degrees of affine semi-regular sequences and their homogenized sequences.
+In this paper, we study the solving degrees for affine semi-regular sequences and their homogenized sequences.
 Some of our results are considered to give mathematically rigorous proofs of the correctness of methods for computing Gröbner bases of the ideal generated by an affine semi-regular sequence.
@@
 We also define the generalized degree of regularity for a sequence of homogeneous polynomials.
-For the homogenization of an affine semi-regular sequence, we relate its generalized degree of regularity with its maximal Gröbner basis degree (i.e., the solving degree of the homogenized sequence).
-The definition of a generalized (cryptographic) semi-regular sequence is also given, and it derives a new cryptographic assumption to estimate the security of cryptosystems and signature schemes.
+For the ideal generated by the homogenization of an affine semi-regular sequence, we relate its generalized degree of regularity with its maximal Gröbner basis degree (i.e., the solving degree for the homogenized sequence).
+The definition of a generalized (cryptographic) semi-regular sequence is also given, and it derives a new cryptographic assumption to estimate the security of cryptosystems.
 From our experimental observation, we raise a conjecture and some questions related to this generalized semi-regularity.
-These new definitions and our results provide a theoretical formulation of (somehow heuristic) discussions done so far in the cryptographic community.
+These definitions and our results provide a theoretical formulation of (somehow heuristic) discussions done so far in the cryptographic community.
 1 Introduction
@@
 Let $R=K[x_{1},\ldots,x_{n}]$ be the polynomial ring in $n$ variables over $K$.
-For a given monomial ordering $\prec$ on the set of monomials in $R$, let $\mathrm{LM}(f)$ denote the leading monomial of $f\in R\smallsetminus\{0\}$ with respect to it.
+For a given monomial ordering $\prec$ on (the set of monomials in) $R$, let $\mathrm{LM}(f)$ denote the leading monomial of $f\in R\smallsetminus\{0\}$ with respect to it.
 For a non-empty subset $F\subset R\smallsetminus\{0\}$, put $\mathrm{LM}(F):=\{\mathrm{LM}(f):f\in F\}$.
@@
 A typical application of computing Gröbner bases is solving the multivariate polynomial (MP) problem: Given a sequence $\bm{F}=(f_{1},\ldots,f_{m})$ of $m$ polynomials $f_{1},\ldots,f_{m}$ in $R\smallsetminus\{0\}$, find $(a_{1},\ldots,a_{n})\in K^{n}$ such that $f_{i}(a_{1},\ldots,a_{n})=0$ for all $i$ with $1\leq i\leq m$.
-A particular case where $f_{1},\ldots,f_{m}$ are all quadratic is called the MQ problem, and its hardness is applied to constructing public-key cryptosystems and digital signature schemes that are expected to be quantum resistant.
+A particular case where $f_{1},\ldots,f_{m}$ are all quadratic is called the MQ problem (cf.
+[44]), and its hardness is applied to constructing public-key cryptosystems that are expected to be quantum resistant.
 Therefore, analyzing the complexity of computing Gröbner bases is one of the most important problems both in theory and in practice.
 An algorithm for computing Gröbner bases was proposed first by Buchberger [7], and so far a number of its improvements such as the $F_{4}$ [19] and $F_{5}$ [20] algorithms have been proposed.
-In determining the complexity of computing Gröbner bases, as we will see in the first paragraph of Subsection 2.2 below, one of the most important cases is the case where the input system is zero-dimensional and where the monomial ordering is graded (i.e., degree-compatible), and we focus on that case in the rest of this paper.
+In determining the complexity of computing Gröbner bases, as we will see in the first paragraph of Subsection 2.2 below, one of the most important cases is the case where the input system is zero-dimensional (see Terminology below for the meaning of zero-dimensional) and the monomial ordering is graded (i.e., degree-compatible).
+Therefore, we focus on that case in the rest of this paper.
 Namely, we suppose that the input sequence $\bm{F}=(f_{1},\ldots,f_{m})$ admits a finite number of zeros in $\mathbb{A}^{n}_{\overline{K}}$ (resp.
@@
 Note that, when $F$ is homogeneous, these three solving degrees coincide with one another (for $\mathcal{A}$ with suitable setting) and we call them the solving degree simply; they are equal to the maximal Gröbner basis degree $\mathrm{max.GB.deg}_{\prec}(F)$ of $F$ with respect to $\prec$.
-In this case, we can apply a well-known bound [31, Theorem 2] by Lazard.
+In this case, we can apply a well-known bound [32, Theorem 2] by Lazard.
 In the following, we mainly treat with the case where $F$ is affine.
@@
 In particular, Caminata-Gorla [9] proved the following upper-bound on $\mathrm{sd}_{\prec}^{\rm mac}(F)$ by using Lazard’s bound:
-  • ([8, Theorem 11]) When $K=\mathbb{F}_{q}$, the solving degree $\mathrm{sd}_{\prec}^{\rm mac}(F)$ for a DRL ordering $\prec$ can be upper-bounded by the Macaulay bound $d_{1}+\cdots+d_{\ell}-\ell+1$ with $d_{1}\geq d_{2}\geq\cdots\geq d_{m}$ and $\ell=\min\{n+1,\ell\}$, if $F$ contains the field equations $x_{i}^{q}-x_{i}$ for all $1\leq i\leq n$.
+  • ([9, Theorem 11]) When $K=\mathbb{F}_{q}$, the solving degree $\mathrm{sd}_{\prec}^{\rm mac}(F)$ for a DRL ordering $\prec$ can be upper-bounded by the Macaulay bound $d_{1}+\cdots+d_{\ell}-\ell+1$ with $d_{1}\geq d_{2}\geq\cdots\geq d_{m}$ and $\ell=\min\{n+1,\ell\}$, if $F$ contains the field equations $x_{i}^{q}-x_{i}$ for all $1\leq i\leq n$.
 As for upper-bounds on the solving degrees $\mathrm{sd}_{\prec}^{\mathcal{A}}(F)$ and $\mathrm{sd}_{\prec}^{\rm mut}(F)$, we know the following:
-  • Semaev-Tenti [40] (see also Tenti’s PhD thesis [41]) constructed a Buchberger-like algorithm $\mathcal{A}$ for the case $K=\mathbb{F}_{q}$ such that $\mathrm{sd}_{\prec}^{\mathcal{A}}(F)\leq 2D-2$ with $D:=d_{\rm reg}(\langle F^{\rm top}\rangle)$ for a DRL ordering $\prec$, assuming that $\{x_{i}^{q}-x_{i}:1\leq i\leq n\}\subset F$ and $\max\{q,\deg(f_{1}),\ldots,\deg(f_{m})\}\leq D$.
-Here $d_{\rm reg}(\langle F^{\rm top}\rangle)$ is the degree of regularity for $F^{\rm top}$, i.e., the smallest integer $d$ with $R_{d}=\langle F^{\rm top}\rangle_{d}$, where $R_{d}$ denotes the homogeneous part (component) of degree $d$ and where we set $I_{d}=I\cap R_{d}$ for a homogeneous ideal $I$ of $R$.
-  • Caminata-Gorla proved in [9, Theorem 3.1] that $\mathrm{sd}_{\prec}^{\rm mut}(F)=\max\{d_{F},{\rm max.GB.deg}_{{\prec}}(F)\}$ for any graded monomial ordering $\prec$, where $d_{F}$ denotes the last fall degree of $F$ defined in [9, Definition 1.5] (originally in [27], [26]).
+  • Semaev-Tenti [41] (see also Tenti’s PhD thesis [42]) constructed a Buchberger-like algorithm $\mathcal{A}$ for the case $K=\mathbb{F}_{q}$ such that $\mathrm{sd}_{\prec}^{\mathcal{A}}(F)\leq 2D-2$ with $D:=d_{\rm reg}(\langle F^{\rm top}\rangle)$ for a DRL ordering $\prec$, assuming that $F=\{f_{1},\ldots,f_{m}\}\cup\{x_{i}^{q}-x\mid 1\leq i\leq n\}$ and $\max\{q,\deg(f_{1}),\ldots,\deg(f_{m})\}\leq D$.
+Here $d_{\rm reg}(\langle F^{\rm top}\rangle)$ is the degree of regularity of $\langle F^{\rm top}\rangle$, i.e., the smallest non-negative integer $d$ with $R_{d}=\langle F^{\rm top}\rangle_{d}$, where $R_{d}$ denotes the homogeneous part (component) of degree $d$ and where we set $I_{d}=I\cap R_{d}$ for a homogeneous ideal $I$ of $R$.
+  • Caminata-Gorla proved in [10, Theorem 3.1] that $\mathrm{sd}_{\prec}^{\rm mut}(F)=\max\{d_{F},{\rm max.GB.deg}_{{\prec}}(F)\}$ with $F=\{f_{1},\ldots,f_{m}\}$ for any graded monomial ordering $\prec$, where $d_{F}$ denotes the last fall degree of $F$ defined in [10, Definition 1.5] (originally in [28], [27]).
 Recently, Salizzoni [40] also proved $\mathrm{sd}_{\prec}^{\rm mut}(F)\leq D+1$, in the case where $\max\{\deg(f_{1}),\ldots,\deg(f_{m})\}\leq D<\infty$.
@@
 affine cryptographic semi-regular) if $\bm{F}^{\rm top}=(f_{1}^{\rm top},\ldots,f_{m}^{\rm top})$ is semi-regular (resp.
-cryptographic semi-regular), see Definitions 2.1.3, 2.1.9, and 2.1.12 for details.
-Note that homogeneous semi-regular sequences are conjectured by Pardue [36, Conjecture B] to be generic sequences of polynomials (see e.g., [36] for the definition of genericness), and affine (cryptographic) semi-regular sequences are often appearing in the construction of multivariate public key cryptosystems and digital signature schemes.
+cryptographic semi-regular), see Definitions 2.1.3, 2.1.10, and 2.1.13 for details.
+Note that homogeneous semi-regular sequences are conjectured by Pardue [37, Conjecture B] to be generic sequences of polynomials (see e.g., [37] for the definition of genericness), and affine (cryptographic) semi-regular sequences are often appearing in the construction of multivariate public key cryptosystems.
 As a sequel of the authors’ previous work [31], we investigate further results on the solving degrees and on behaviors of the computation of Gröbner bases.
@@
 This characterization is useful to analyze the Gröbner basis computation for both $F$ and $F^{h}$.
-Theorem 1 (Theorem 3.1.1, Remark 3.1.2, Remark 3.1.3 and Corollary 3.1.5).
+Theorem 1 (Theorem 3.1.1, Remark 3.1.2, and Corollary 3.1.4).
 With notation as above, assume that $\bm{F}$ is affine cryptographic semi-regular, and put $D:=d_{\rm reg}(\langle F^{\rm top}\rangle)$. Then, we have the following:
@@
     There exists $d_{0}$ such that ${\rm HF}_{R^{\prime}/\langle F^{h}\rangle}(d_{0})={\rm HF}_{R^{\prime}/\langle F^{h}\rangle}(d)$ for all $d$ with $d\geq d_{0}$, namely the number of projective zeros for $F^{h}$ is finite at most.
-Here, $d_{0}$ gives an upper-bound on the solving degree of $F^{h}$ (or equivalently the maximal degree of the Gröbner basis of $\langle F^{h}\rangle$).
   4. (4)
-    ${\rm HS}_{R^{\prime}/\langle F^{h}\rangle}(z)\equiv\prod_{i=1}^{m}(1-z^{d_{i}})/(1-z)^{n+1}\pmod{z^{D}}$, so that $F^{h}$ is $D$-regular, equivalently $\mathrm{syz}(F^{h})_{<D}=\mathrm{tsyz}(F^{h})_{<D}$.
+    ${\rm HS}_{R^{\prime}/\langle F^{h}\rangle}(z)\equiv\prod_{i=1}^{m}(1-z^{d_{i}})/(1-z)^{n+1}\pmod{z^{D}}$, so that $\bm{F}^{h}$ is $D$-regular, equivalently $\mathrm{syz}(F^{h})_{<D}=\mathrm{tsyz}(F^{h})_{<D}$.
 Here we denote by $\mathrm{syz}(F^{h})$ and $\mathrm{tsyz}(F^{h})$ the module of syzygies of $F^{h}$ and that of trivial syzygies of $F^{h}$, respectively (see Appendix A.1 for the definition of $\mathrm{syz}(F^{h})$ and $\mathrm{tsyz}(F^{h})$).
-As for (3) of Theorem 1, it follows from the proof of Lazard’s bound [31, Theorem 2] that $\mathrm{max.GB.deg}_{\prec^{h}}(F^{h})\leq d_{0}$ for a DRL ordering $\prec$ (we give an explicit proof for this in Lemma 2.2.2 below), where $\prec^{h}$ is the homogenization of $\prec$.
-As in [8, Theorem 11] recalled above, we can apply Lazard’s bound to obtaining $d_{0}\leq d_{1}+\cdots+d_{\ell}-\ell+1$ with $d_{i}=\deg(f_{i})$ and $\ell=\min\{n+1,\ell\}$, assuming $d_{1}\geq\cdots\geq d_{m}$ in descending order.
+As for (3) of Theorem 1, similarly to the proof of Lazard’s bound [32, Theorem 2], it can be proved (see Proposition 2.3.6 below) that $\mathrm{max.GB.deg}_{\prec^{h}}(F^{h})\leq\mathrm{max}\{D,D^{\prime}\}$, where $\prec^{h}$ is the homogenization of $\prec$ and where $D^{\prime}:=\min\{d_{0}\mid{\rm HF}_{R^{\prime}/\langle F^{h}\rangle}(d_{0})={\rm HF}_{R^{\prime}/\langle F^{h}\rangle}(d)\mbox{ for all $d\geq d_{0}$}\}$.
+As in [9, Theorem 11] recalled above, we can apply Lazard’s bound to obtaining $\mathrm{max.GB.deg}_{\prec^{h}}(F^{h})\leq d_{1}+\cdots+d_{\ell}-\ell+1$ with $d_{i}=\deg(f_{i})$ and $\ell=\min\{n+1,\ell\}$, assuming $d_{1}\geq\cdots\geq d_{m}$ in descending order.
 As an additional result in this paper, we also obtain the following upper-bound on the solving degree of $F^{h}$:
@@
   2. (2)
-    Let $S_{0}$ be the saturation exponent of $(\langle F^{h}\rangle:\langle y^{\infty}\rangle)$, that is, the minimum integer $s$ such that $(\langle F^{h}\rangle:\langle y^{s}\rangle)=(\langle F^{h}\rangle:\langle y^{\infty}\rangle)$.
+    Let $S_{0}$ be the saturation exponent of $(\langle F^{h}\rangle:\langle y^{\infty}\rangle)$, that is, the minimum non-negative integer $s$ such that $(\langle F^{h}\rangle:\langle y^{s}\rangle)=(\langle F^{h}\rangle:\langle y^{\infty}\rangle)$.
 Then the solving degree of $F^{h}$ is upper-bounded by $D+{S_{0}}$.
 Based on Theorem 1, we can explore the computations of reduced Gröbner bases of $\langle F\rangle$, $\langle F^{h}\rangle$, and $\langle F^{\rm top}\rangle$ in Section 4 below, dividing the cases into the degree less than $D$ or not.
-More precisely, denoting by $G$, $G_{\rm hom}$, and $G_{\rm top}$ the reduced Gröbner bases of $\langle F\rangle$, $\langle F^{h}\rangle$, and $\langle F^{\rm top}\rangle$ respectively, where their monomial orderings are DRL $\prec$ or its extension $\prec^{h}$, we revisit [30, Section 5] and obtain more precise results:
-Theorem 3 (Section 4; cf. [30, Section 5]).
+More precisely, denoting by $G$, $G_{\rm hom}$, and $G_{\rm top}$ the reduced Gröbner bases of $\langle F\rangle$, $\langle F^{h}\rangle$, and $\langle F^{\rm top}\rangle$ respectively, where their monomial orderings are DRL $\prec$ or its homogenization $\prec^{h}$, we revisit [31, Section 5] and obtain more precise results:
+Theorem 3 (Section 4; cf. [31, Section 5]).
 With notation as above, assume that $\bm{F}$ is affine cryptographic semi-regular, and that $D:=d_{\rm reg}(\langle F^{\rm top}\rangle)<\infty$.
@@
 Note that (2) and the first half of (4) hold not necessarily assuming the affine cryptographic semi-regularity of $\bm{F}$.
+Remark 1.
+In (4) of Theorem 3, the complexity of the algorithm $\mathcal{A}$ (with respect to the number of arithmetic operations on $K$) is
+$O\left(m\binom{n+D}{D}^{\omega}+\binom{n+D}{D}^{2}\binom{n+D-1}{D-1}^{2}\binom{n+2D-2}{2D-2}\right),$
+which is reduced to
+$O\left(m\binom{n+D}{D}^{\omega}+\binom{n+D-1}{D-1}^{2}\binom{n+2D-2}{2D-2}\right)$
+if we can avoid every 0-reduction completely, where $32\leq\omega<$ 3 is the exponent of matrix multiplication.
+This can be proved in a way similar to the proof of [42, Theorem 3.65] together with Theorem A.4.1 below: A proof will be given in Appendix A.5 below for the readers’ convenience.
 In particular, we rigorously prove some existing results, which are often used for analyzing the complexity of computing Gröbner bases, and moreover extend them to our case.
-Furthermore, based on Lemma 2.2.2 below, for homogeneous ideals of projective dimension 0, we naturally extend the notion of degree of regularity.
-The new notion of generalized degree of regularity plays a very important role on the computation of Gröbner bases for such ideals, see Subsections 2.3 and 4.3.
-The following proposition summarizes theoretical results on this generalized degree of regularity:
-Proposition 1 (Proposition 2.3.3).
+Furthermore, based on Lemma 2.2.2 below, for zero-dimensional homogeneous ideals (see Terminology below), we naturally extend the notion of degree of regularity: We shall define the generalized degree of regularity $\tilde{d}_{\rm reg}(I)$ of such a homogeneous ideal $I$, as the index of regularity (or called the Hilbert regularity) $i_{\rm reg}(I)$ of $I$.
+The generalized degree of regularity of $\langle F^{h}\rangle$ plays a very important role in analyzing the computation of Gröbner bases for such ideals, see Subsections 2.3 and 4.3.
+The following proposition summarizes several theoretical results, which are proved rigidly in this paper, on this generalized degree of regularity:
+Proposition 1 (Lemma 2.3.5 and Proposition 2.3.6; see also Subsection A.3).
 With notation as above, assume that $R/\langle F^{\rm top}\rangle$ is Artinian, and that $\bm{F}^{\rm top}$ is cryptographic semi-regular. Then we have the following:
@@
   2. 2.
-    $\mathrm{max.GB.deg}_{\prec^{h}}(F^{h})\leq\max\{{d}_{\rm reg}(\langle F^{\rm top}\rangle),\widetilde{d}_{\rm reg}(\langle F^{h}\rangle)\}$ for any graded monomial order $\prec$.
+    $\mathrm{max.GB.deg}_{\prec^{h}}(F^{h})\leq\max\{{d}_{\rm reg}(\langle F^{\rm top}\rangle),\widetilde{d}_{\rm reg}(\langle F^{h}\rangle)\}$ for any graded monomial ordering $\prec$.
 Moreover, the equality holds in the second inequality if $\langle\mathrm{LM}(\langle F^{h}\rangle)\rangle$ is weakly reverse lexicographic.
-Here, a weakly reverse lexicographic ideal is a monomial ideal $J$ such that if $x^{\alpha}$ is one of the minimal generators of $J$ then every monomial of the same degree which preceeds $x^{\alpha}$ must belong to $J$ as well (see [36, Section 4] for the original definition).
-Finally, in Subsection 4.3, we give some observation on the behavior of Gröbner basis computation based on our experiments, from which we arrive at a conjecture on the Hilbert–Poincaré series for affine polynomial sequences without constant terms.
-For this conjecture, we generalize the notion of cryptographic semi-regular in Subsection 2.3.
+Here, a weakly reverse lexicographic ideal is a monomial ideal $J$ such that if $x^{\alpha}$ is one of the minimal generators of $J$ then every monomial of the same degree which preceeds $x^{\alpha}$ must belong to $J$ as well (see [37, Section 4] for the original definition).
+Corollary 1 (Corollary A.4.2 and Remark A.4.4; see also Subsection 4.3).
+Under the same setting as in Proposition 1, when $d_{1},\ldots,d_{m}$ are fixed, a Gröbner basis of $\langle F\rangle_{R}$ can be computed in
+$O\left(m\binom{n+\widetilde{d}_{\rm reg}(\langle F^{h}\rangle)}{\widetilde{d}_{\rm reg}(\langle F^{h}\rangle)}^{\omega}\right)$
+if the complexity of substituting $y=1$ to $G_{\rm hom}$ is negligible.
+Furthermore, if $m=n$, then $\widetilde{d}_{\rm reg}(\langle F^{h}\rangle)=d_{\rm reg}(\langle F^{\rm top}\rangle)-1$ with $d_{\rm reg}(\langle F^{\rm top}\rangle)=\sum_{j=1}^{n}d_{j}-n+1$.
+If $m>n$ and if $\bm{F}^{h}$ is generalized cryptographic semi-regular, i.e., $\widetilde{d}_{\rm reg}(\langle F^{h}\rangle)$-regular, then we have
+$\widetilde{d}_{\rm reg}(\langle F^{h}\rangle)\leq{\deg\left(\left[\frac{\prod_{i=1}^{m}(1-z^{d_{i}})}{(1-z)^{n+1}}\right]\right)}+1,$
+where []delimited-[][$\cdot$][ ⋅ ] means truncating a formal power series over $\mathbb{Z}$ after the last consecutive positive coefficient.
+Finally, in Subsection 4.3, we give some observation on the behavior of Gröbner basis computation based on our experiments, from which we arrive at a conjecture (Conjecture 4.3.4 below) on the Hilbert–Poincaré series for affine polynomial sequences without constant terms.
+For this conjecture, we also generalize the notion of cryptographic semi-regular in Subsection 2.3.
 Notation
@@
   • $f^{h}$: The homogenization of $f\in R\smallsetminus\{0\}$ by an extra variable $y$, say $f^{h}:=y^{\deg(f)}f(x_{1}/y,\ldots,x_{n}/y)$.
-  • ${\rm HF}_{M}$: The Hilbert function of a finitely generated graded $R$-module $M=\bigoplus_{d\in\mathbb{Z}}M_{d}$, say ${\rm HF}_{M}(d)=\mathrm{dim}_{K}M_{d}$ for each $d\in{\color[rgb]{0,0,0}\mathbb{Z}_{\geq 0}}$.
-  • ${\rm HS}_{M}$: The Hilbert–Poincaré series of a finitely generated graded $R$-module $M=\bigoplus_{d\in\mathbb{Z}}M_{d}$, say ${\rm HS}_{M}(z)=\sum_{d=0}^{\infty}{\rm HF}_{M}(d)z^{d}\in\mathbb{Z}\llbracket z\rrbracket$.
+  • ${\rm HF}_{M}$: The Hilbert function of a finitely generated graded $R$-module $M=\bigoplus_{d\in\mathbb{Z}}M_{d}$, say ${\rm HF}_{M}(d)=\mathrm{dim}_{K}M_{d}$ for each $d\in\mathbb{Z}$.
+  • ${\rm HS}_{M}$: The Hilbert–Poincaré series of a finitely generated graded $R$-module $M=\bigoplus_{d\in\mathbb{Z}_{\geq 0}}M_{d}$, say ${\rm HS}_{M}(z)=\sum_{d=0}^{\infty}{\rm HF}_{M}(d)z^{d}\in\mathbb{Z}\llbracket z\rrbracket$.
   • $K_{\bullet}(f_{1},\ldots,f_{m})$: The Koszul complex on a sequence $(f_{1},\ldots,f_{m})$ of homogeneous polynomials in $R$.
@@
 As for the definition of Koszul complex and homogenization, see Appendix A for details.
+Terminology
+Following [14, Chapter IX], we use the terminology ‘zero-dimensional’ for a polynomial ideal by the following rule:
+  • If the ideal is not necessarily homogeneous, we say that it is zero-dimensional if the number of its affine zeros over $\overline{K}$ is finite at most.
+This is equivalent to that the affine Hilbert polynomial of the ideal is a constant.
+  • If the ideal is homogeneous, unless otherwise noted, we say that it is zero-dimensional if the number of its projective zeros over $\overline{K}$ is finite at most.
+This is equivalent to that the Hilbert polynomial of the ideal is a constant.
+As for the definition of affine Hilbert polynomials and Hilbert polynomials, we refer to [14, Chapter IX].
 2 Preliminaries
-In this section, we recall definitions of semi-regular sequences and solving degrees, and collect some known facts related to them.
+In this section, we recall definitions of semi-regular sequences and solving degrees, and collect some known facts related to them. Subsequently, we also extend the notion of degree of regularity.
 2.1 Semi-regular sequences
 We first review the notion of semi-regular sequence defined by Pardue [37].
-Definition 2.1.1 (Semi-regular sequences, [36, Definition 1]).
+Definition 2.1.1 (Semi-regular sequences, [37, Definition 1]).
 Let $I$ be a homogeneous ideal of $R$.
@@
 The semi-regularity is characterized by equivalent conditions in Proposition 2.1.2 below. In particular, the fourth condition enables us to compute the Hilbert–Poincaré series of each $A^{(i)}$.
-Proposition 2.1.2 (cf. [36, Proposition 1]).
+Proposition 2.1.2 (cf. [37, Proposition 1]).
 With notation as above, the following are equivalent:
@@
     For each $i$ with $1\leq i\leq m$, we have ${\rm HS}_{A^{(i)}}(z)=\left[\frac{\prod_{j=1}^{i}(1-z^{d_{j}})}{(1-z)^{n}}\right]$.
-When $K$ is an infinite field, Pardue also conjectured in [36, Conjecture B] that generic polynomial sequences are semi-regular.
+When $K$ is an infinite field, Pardue also conjectured in [37, Conjecture B] that generic polynomial sequences are semi-regular.
 We next review the notion of cryptographic semi-regular sequence, which is defined by a condition weaker than one for semi-regular sequence.
@@
 Diem [15] also formulated cryptographic semi-regular sequences, in terms of commutative and homological algebra.
-The terminology “cryptographic” was named by Bigdeli et al.
+The terminology ‘cryptographic’ was named by Bigdeli et al.
 in their recent work [5], in order to distinguish such a sequence from a semi-regular one defined by Pardue (see Definition 2.1.1).
-Definition 2.1.3 ([2, Definition 3]; see also [14, Definition 1]).
+Definition 2.1.3 ([2, Definition 3]; see also [15, Definition 1]).
 Let $f_{1},\ldots,f_{m}\in R$ be homogeneous polynomials of positive degrees $d_{1},\ldots,d_{m}$ respectively, and put $I=\langle f_{1},\ldots,f_{m}\rangle_{R}$.
@@
 Diem [15] determined the (truncated) Hilbert series of $d$-regular sequences as in the following proposition:
-Theorem 2.1.4 (cf. [14, Theorem 1]).
+Theorem 2.1.4 (cf. [15, Theorem 1]).
 With the same notation as in Definition 2.1.3, the following are equivalent for each $d$ with $d\geq\mathrm{max}\{d_{i}:1\leq i\leq m\}$:
@@
     $H_{1}(K_{\bullet}(f_{1},\ldots,f_{m}))_{\leq d-1}=0$.
-Proposition 2.1.5 ([14, Proposition 2 (a)]).
+Proposition 2.1.5 ([15, Proposition 2 (a)]).
 With the same notation as in Definition 2.1.3, let $D$ and $i$ be natural numbers.
@@
 A finitely generated graded $R$-module $M$ is said to be Artinian if there exists a sufficiently large $D\in\mathbb{Z}$ such that $M_{d}=0$ for all $d\geq D$.
-Definition 2.1.7 ([2, Definition 4], [3, Definition 4]).
+Definition 2.1.7 ([2, Definition 4], [4, Definition 4]).
 For a homogeneous ideal $I$ of $R$, we define its degree of regularity $d_{\rm reg}(I)$ as follows: If the finitely generated graded $R$-module $R/I$ is Artinian, we set $d_{\rm reg}(I):=\mathrm{min}\{d:R_{d}=I_{d}\}$, and otherwise we set $d_{\rm reg}(I):=\infty$.
-As for an upper-bound on the degree of regularity, we refer to [23, Theorem 21].
+As for an upper-bound on the degree of regularity, we refer to [24, Theorem 21].
+An elementary but important fact that relates $d_{\rm reg}(I)$ and ${\rm max.GB.deg}_{\prec}(I)$ is the following (the proof is straightforward, but write it here for the readers’ convenience):
+Lemma 2.1.8.
+For any homogeneous ideal $I$ of $R$ and any graded ordering $\prec$ on the set of monomials in $R$, we have ${\rm max.GB.deg}_{\prec}(I)\leq d_{\rm reg}(I)$.
+Proof.
+The case $d_{\rm reg}(I)=\infty$ is trivial, so we consider the case where $d_{\rm reg}(I)<\infty$, namely $R/I$ is Artinian.
+Let $G$ be the reduced Gröbner basis of $I$ with respect to $\prec$, and put $D:=d_{\rm reg}(I)$.
+Assume for a contradiction that there were an element $g\in G$ such that $\deg(g)>D$.
+Since $\prec$ is graded, we have $\deg(g)=\deg\mathrm{LM}(g)$, so we choose any monomial $M\in R_{D}$ dividing $\mathrm{LM}(g)$.
+Then, it follows from $R_{D}=I_{D}$ that $M$ is divisible by $\mathrm{LM}(g^{\prime})$ for some $g^{\prime}\in G$ with $\deg(g^{\prime})=\deg\mathrm{LM}(g^{\prime})\leq D$, so that $\mathrm{LM}(g)$ is divisible by $\mathrm{LM}(g^{\prime})$ with $g\neq g^{\prime}$.
+This contradicts to that $G$ is reduced.
+∎
 Remark 2.1.9.
 In Definition 2.1.7, since $R/I$ is Noetherian, it is Artinian if and only if it is of finite length.
-In this case, the degree of regularity $d_{\rm reg}(I)$ is equal to the Castelnuovo-Mumford regularity $\mathrm{reg}(I)$ of $I$ (see e.g., [16, §20.5] for the definition), whence $d_{\rm reg}(I)=\mathrm{reg}(I)=\mathrm{reg}(R/I)+1$.
-Definition 2.1.9 ([2, Definition 5], [3, Definition 5]; see also [14, Section 2]).
+In this case, the degree of regularity $d_{\rm reg}(I)$ is equal to the Castelnuovo-Mumford regularity $\mathrm{reg}(I)$ of $I$ (see e.g., [17, §20.5] for the definition), whence $d_{\rm reg}(I)=\mathrm{reg}(I)=\mathrm{reg}(R/I)+1$.
+Definition 2.1.10 ([2, Definition 5], [4, Definition 5]; see also [15, Section 2]).
 A sequence $(f_{1},\ldots,f_{m})\in(R\smallsetminus K)^{m}$ of homogeneous polynomials is said to be cryptographic semi-regular if it is $d_{\rm reg}(I)$-regular, where we set $I=\langle f_{1},\ldots,f_{m}\rangle_{R}$.
 The cryptographic semi-regularity is characterized by equivalent conditions in Proposition 2.1.11 below.
-Proposition 2.1.10 ([14, Proposition 1 (d)]; see also [3, Proposition 6]).
+Proposition 2.1.11 ([15, Proposition 1 (d)]; see also [4, Proposition 6]).
 With the same notation as in Definition 2.1.3, we put $D=d_{\rm reg}(I)$. Then, the following are equivalent:
@@
 It can be proved (cf.
-[36]) that Fröberg’s conjecture is equivalent to Pardue’s one [36, Conjecture B].
-We also note that Moreno-Socías conjecture [35] is stronger than the above two conjectures, see [36, Theorem 2] for a proof.
+[37]) that Fröberg’s conjecture is equivalent to Pardue’s one [37, Conjecture B].
+We also note that Moreno-Socías conjecture [36] is stronger than the above two conjectures, see [37, Theorem 2] for a proof.
 It follows from the fourth condition of Proposition 2.1.2 together with the second condition of Proposition 2.1.11 that the semi-regularity implies the cryptographic semi-regularity.
@@
 Even when $F$ is homogeneous, one can efficiently convert $G$ to $G^{\prime}$ by Gröbner walk [13] (or Hilbert driven [43] if both $\prec$ and $\prec^{\prime}$ are graded).
-From this, we focus on the case where the monomial ordering is graded, and if necessary we also assume that the ideal generated by the input polynomials is zero-dimensional.
+From this, we focus on the case where the monomial ordering is graded, and if necessary we also assume that the ideal generated by the input polynomials is zero-dimensional (see Terminology in Section 1 for the meaning of zero-dimensional).
 Definitions of solving degrees
@@
   • (I)
-    As the first definition, we define the solving degree of an algorithm to compute a Gröbner basis as the highest degree of the polynomials involved during the execution of the algorithm, see [15, p.
+    As the first definition, we define the solving degree of an algorithm to compute a Gröbner basis as the highest degree of the polynomials involved during the execution of the algorithm, see [16, p.
 36].
-For example, applying Buchberger’s algorithm or its variants such as $F_{4}$ with the normal strategy, we collect critical S-pairs with the lowest degree and then reduce the corresponding S-polynomials in each iteration of the main loop of reductions.
+For example, applying Buchberger’s algorithm or its variants such as $F_{4}$ with the normal strategy (or called normal selection strategy, see [14, \$II.10]), we collect critical S-pairs with the lowest degree and then reduce the corresponding S-polynomials in each iteration of the main loop of reductions.
 The lowest degree of each iteration is called the step degree.
 Then the solving degree is defined as the highest step degree.
-Instead, we may adopt the highest degree of $S$-polynomials appearing in the whole computation as in [41] and [40] by Semaeve-Tenti, and in this case we use the terminology “the solving degree in the strict sense”.
+Instead, we may adopt the highest degree of $S$-polynomials appearing in the whole computation as in [42] and [41] by Semaeve-Tenti, and in this case we use the terminology ‘the solving degree in the strict sense’.
   • (I)’
     Their is a variant of the above first definition, where the solving degree is defined as a value depending not only on an algorithm but also on its implementation.
-More precisely, in [15, Section 2.1], the authors use the term solving degree for the step degree at which it takes the most amount of time among all iterations.
+More precisely, in [16, Section 2.1], the authors use the term solving degree for the step degree at which it takes the most amount of time among all iterations.
 In the cryptographic literature, the term solving degree often means this solving degree.
@@
   • (II)
-    We define the solving degree of $F$ with respect to a fixed (graded) monomial ordering as the lowest degree $d$ at which the reduced row echelon form (RREF) of $M_{\leq d}(F)$ produces a Gröbner basis of $F$.
+    We define the solving degree of $F$ with respect to a fixed (graded) monomial ordering $\prec$ as the lowest degree $d$ at which the reduced row echelon form (RREF) of $M_{\leq d}(F)$ produces a Gröbner basis of $F$ with respect to $\prec$.
 Note that the computation of the RREF of $M_{\leq d}(F)$ corresponds to the standard XL algorithm [12], which is based on an idea of Lazard [32].
@@
 If the resulting matrix produces a Gröbner basis of $F$, then we stop, and otherwise we proceed to the next degree, $d+1$.
-This strategy computes a basis of $V_{F,d}$ for each $d$, and therefore the smallest $d$ for which the mutant strategy terminates is equal to the solving degree of $F$ in terms of (III), see [24, Theorem 1].
+This strategy computes a basis of $V_{F,d}$ for each $d$, and therefore the smallest $d$ for which the mutant strategy terminates is equal to the solving degree of $F$ in terms of (III), see [25, Theorem 1].
 As in [24], we refer to the algorithms such as Mutant-XL [8] and MXL2 [35] that employ this strategy as mutant algorithms.
@@
 [9], [5], [25], [10], [24]), Gorla et al.
-provided a mathematical formulation for the relations between the solving degrees $\mathrm{sd}_{\prec}^{\rm mac}(F)$ and $\mathrm{sd}_{\prec}^{\rm mut}(F)$ and algebraic invariants coming from $F$, such as the maximal Gröbner basis degree, the degree of regularity, the Castelnuovo–Mumford regularity, the first and last degrees, and so on.
+provided a mathematical formulation for the relations between the solving degrees $\mathrm{sd}_{\prec}^{\rm mac}(F)$ and $\mathrm{sd}_{\prec}^{\rm mut}(F)$ and algebraic invariants coming from $F$, such as the maximal Gröbner basis degree, the degree of regularity, the Castelnuovo–Mumford regularity, the first and last fall degrees, and so on.
 Here, the maximal Gröbner basis degree of the ideal $\langle F\rangle_{R}$ is the maximal degree of elements in the reduced Gröbner basis of $\langle F\rangle_{R}$ with respect to a fixed monomial ordering $\prec$, and is denoted by $\mathrm{max.GB.deg}_{\prec}(F)$.
@@
 Since we are now considering the zero-dimensional case, we can apply Lazard’s upper-bound below.
-In the non-homogeneous case, i.e., $F$ contains at least one non-homogeneous element, the euqalities in (2.2.1) do not hold in general, and it is not so easy to estimate any of the solving degrees.
-A straightforward way of bounding the solving degrees in the non-homogeneous case is to apply the homogenization as follows.
+In the inhomogeneous case, i.e., $F$ contains at least one inhomogeneous element, the equalities in (2.2.1) do not hold in general, and it is not so easy to estimate any of the solving degrees.
+A straightforward way of bounding the solving degrees in the inhomogeneous case is to apply the homogenization as follows.
 We set $\prec$ as the DRL ordering on $R$ with $x_{n}\prec\cdots\prec x_{1}$, and fix it throughout the rest of this subsection.
@@
 see [9] for a proof. Here, we also recall Lazard’s bound for the maximal Gröbner basis degree of $\langle F^{h}\rangle_{R^{\prime}}$ with $R^{\prime}=R[y]$:
-Theorem 2.2.1 (Lazard; [31, Theorem 2], [32, Théorèm 3.3]).
+Theorem 2.2.1 (Lazard; [32, Theorem 2], [33, Théorèm 3.3]).
 With notation as above, we assume that the number of projective zeros of $F^{h}$ is finite (and therefore $m\geq n$), and that $f_{1}^{h}=\cdots=f_{m}^{h}=0$ has no non-trivial solution over the algebraic closure $\overline{K}$ with $y=0$, i.e., $F^{\rm top}$ has no solution in $\overline{K}^{n}$ other than $(0,\ldots,0)$.
@@
 Lemma 2.2.2.
-With notation as above, let $d_{0}$ be a positive integer satisfying the following two properties:
+Let $I$ be a homogeneous ideal of $R^{\prime}$, and let $d_{0}$ be a positive integer satisfying the following two properties:
   1. 1.
-    The multiplication-by-$y$ map $(R^{\prime}/\langle F^{h}\rangle)_{d_{0}-1}\longrightarrow(R^{\prime}/\langle F^{h}\rangle)_{d_{0}}$ is surjective.
+    The multiplication-by-$y$ map $(R^{\prime}/I)_{d_{0}-1}\longrightarrow(R^{\prime}/I)_{d_{0}}$ is surjective.
   2. 2.
-    For any $d\in\mathbb{Z}$ with $d\geq d_{0}$, the multiplication-by-$y$ map $(R^{\prime}/\langle F^{h}\rangle)_{d}\longrightarrow(R^{\prime}/\langle F^{h}\rangle)_{d+1}$ is injective.
-Then we have $\mathrm{max.GB.deg}_{\prec^{h}}(F^{h})\leq d_{0}$.
+    For any $d\in\mathbb{Z}$ with $d\geq d_{0}$, the multiplication-by-$y$ map $(R^{\prime}/I)_{d}\longrightarrow(R^{\prime}/I)_{d+1}$ is injective.
+Then we have $\mathrm{max.GB.deg}_{\prec^{\prime}}(I)\leq d_{0}$, where $\prec^{\prime}$ is a homogenization (with respect to $y$) of any graded monomial ordering on $R$, see Subsection A.2 below for details.
 Proof.
-Let $G$ be a Gröbner basis of $\langle F^{h}\rangle$ with respect to $\prec^{h}$.
+Let $G$ be a Gröbner basis of $I$ with respect to $\prec^{\prime}$.
 Clearly, we may suppose that each element of $G$ is homogeneous.
-It suffices to prove that $G_{\leq d_{0}}:=\{g\in G:\mathrm{deg}(g)\leq d_{0}\}$ is a Gröbner basis of $\langle F^{h}\rangle$ with respect to $\prec^{h}$.
-Indeed, the maximal degree of the reduced Gröbner basis of $\langle F^{h}\rangle$ with respect to $\prec^{h}$ is not greater than that of any Gröbner basis of $\langle F^{h}\rangle$ with respect to $\prec^{h}$.
-Let $f\in\langle F^{h}\rangle$, and $d:=\mathrm{deg}(f)$.
+It suffices to prove that $G_{\leq d_{0}}:=\{g\in G:\mathrm{deg}(g)\leq d_{0}\}$ is a Gröbner basis of $I$ with respect to $\prec^{\prime}$.
+Indeed, the maximal degree of the reduced Gröbner basis of $I$ with respect to $\prec^{\prime}$ is not greater than that of any Gröbner basis of $I$ with respect to $\prec^{\prime}$.
+Let $f\in I$, and $d:=\mathrm{deg}(f)$.
 We show that there exists $g\in G_{\leq d_{0}}$ with $\mathrm{LM}(g)\mid\mathrm{LM}(f)$, by the induction on $d$.
-It suffices to consider the case where $f$ is homogeneous, since $\langle F^{h}\rangle$ is homogeneous.
+It suffices to consider the case where $f$ is homogeneous, since $I$ is homogeneous.
 The case where $d\leq d_{0}$ is clear, and so we assume $d>d_{0}$.
-First, if $\mathrm{LM}(f)\in R=K[x_{1},\ldots,x_{n}]$ (namely $y\nmid\mathrm{LM}(f)$), we choose an arbitrary monomial $t\in R$ of degree $d_{0}$ with $t\mid\mathrm{LM}(f)$.
-Since the multiplication map $(R^{\prime}/\langle F^{h}\rangle)_{d_{0}-1}\longrightarrow(R^{\prime}/\langle F^{h}\rangle)_{d_{0}}$ by $y$ is surjective, there exists a homogeneous polynomial $h\in(R^{\prime})_{d_{0}-1}$ such that $h_{1}:=t-yh\in\langle F^{h}\rangle$.
+First, we consider the case where $y\nmid\mathrm{LM}(f)$ (namely $\mathrm{LM}(f)\in R=K[x_{1},\ldots,x_{n}]$).
+We choose an arbitrary monomial $t\in R$ of degree $d_{0}$ with $t\mid\mathrm{LM}(f)$.
+Since the multiplication map $(R^{\prime}/I)_{d_{0}-1}\longrightarrow(R^{\prime}/I)_{d_{0}}$ by $y$ is surjective, there exists a homogeneous polynomial $h\in(R^{\prime})_{d_{0}-1}$ such that $h_{1}:=t-yh\in I$.
 Here, $h_{1}$ is homogeneous of degree $d_{0}$, and $y\nmid t$, whence $\mathrm{LM}(h_{1})=t$.
@@
 Next, assume that $y\mid\mathrm{LM}(f)$.
-In this case, it follows from the definition of $\prec^{h}$ that any other term in $f$ is also divisible by $y$, so that $f\in\langle y\rangle$.
+In this case, it follows from the definition of $\prec^{\prime}$ that any other term in $f$ is also divisible by $y$, so that $f\in\langle y\rangle$.
 Hence, we can write $f=yf_{1}$ for some homogeneous $f_{1}\in R^{\prime}$.
-By $d-1\geq d_{0}$, the multiplication map $(R^{\prime}/\langle F^{h}\rangle)_{d-1}\longrightarrow(R^{\prime}/\langle F^{h}\rangle)_{d}$ by $y$ is injective, so that $f_{1}\in\langle F^{h}\rangle_{d-1}$.
+By $d-1\geq d_{0}$, the multiplication map $(R^{\prime}/I)_{d-1}\longrightarrow(R^{\prime}/I)_{d}$ by $y$ is injective, so that $f_{1}\in I_{d-1}$.
 By the induction hypothesis, there exists $g\in G_{\leq d_{0}}$ such that $\mathrm{LM}(g)\mid\mathrm{LM}(f_{1})$.
 Since $\mathrm{LM}(f)=y\mathrm{LM}(f_{1})$, we obtain $\mathrm{LM}(g)\mid\mathrm{LM}(f)$.
-We have proved that $G_{\leq d_{0}}$ is a Gröbner basis of $\langle F^{h}\rangle$ with respect to $\prec^{h}$.
+We have proved that $G_{\leq d_{0}}$ is a Gröbner basis of $I$ with respect to $\prec^{\prime}$.
 ∎
-Lazard proved that we can take $d_{1}+\cdots+d_{\ell}-\ell+1$ in Theorem 2.2.1 as $d_{0}$ in Lemma 2.2.2.
+Lazard proved that we can take $d_{1}+\cdots+d_{\ell}-\ell+1$ in Theorem 2.2.1 as $d_{0}$ in Lemma 2.2.2, where we also take $I$ and $\prec^{\prime}$ to be $\langle F^{h}\rangle$ and $\prec^{h}$ respectively.
 Lazard’s bound given in (2.2.2) is also referred to as the Macaulay bound, and it provides an upper-bound for the solving degree of $F$ with respect to a DRL ordering.
@@
 $\mathrm{max.GB.deg}_{\prec^{\prime}}(F)\leq d_{\rm reg}(\langle F^{\rm top}\rangle)$ (2.2.3)
-for any graded monomial ordering $\prec^{\prime}$ on $R$, see [8, Remark 15] or Lemma 4.2.4 below for a proof.
-Both $d_{\rm reg}(\langle F^{\rm top}\rangle)$ and $\mathrm{\rm sd}_{\prec}^{\rm mac}(F)$ are greater than or equal to ${\rm max.GB.deg}_{\prec}(F)$, whereas it is pointed out in [4], [8], and [9] by explicit examples that any of the degree of regularity and the first fall degree does not produce an estimate for the solving degrees in general, even when $F$ is an affine (cryptographic) semi-regular sequence.
-Caminata-Gorla proved in [9] that the solving degree $\mathrm{sd}_{\prec}^{\rm mut}(F)$ is nothing but the last fall degree if it is greater than the maximal Gröbner basis degree:
-Theorem 2.2.3 ([9, Theorem 3.1]).
-With notation as above, for any degree-compatible monomial ordering $\prec^{\prime}$ on $R$, we have the following equality:
-$\mathrm{sd}_{\prec^{\prime}}^{\rm mut}(F)=\max\{d_{F},{\rm max.GB.deg}_{{\color[rgb]{0,0,0}\prec^{\prime}}}(F)\},$
-where $d_{F}$ denotes the last fall degree of $F$ defined in [9, Definition 1.5] (originally in [27], [26]).
+for any graded monomial ordering $\prec^{\prime}$ on $R$, see [9, Remark 15] or Lemma 4.2.4 below for a proof.
+Both $d_{\rm reg}(\langle F^{\rm top}\rangle)$ and $\mathrm{\rm sd}_{\prec}^{\rm mac}(F)$ are greater than or equal to ${\rm max.GB.deg}_{\prec}(F)$, whereas it is pointed out in [5], [9], and [10] by explicit examples that any of the degree of regularity and the first fall degree does not produce an estimate for the solving degrees in general, even when $\bm{F}$ is an affine (cryptographic) semi-regular sequence.
+Caminata-Gorla proved in [10] that the solving degree $\mathrm{sd}_{\prec}^{\rm mut}(F)$ is nothing but the last fall degree $d_{F}$ if $d_{F}$ is greater than the maximal Gröbner basis degree:
+Theorem 2.2.3 ([10, Theorem 3.1]).
+With notation as above, for any graded monomial ordering $\prec^{\prime}$ on $R$, we have the following equality:
+$\mathrm{sd}_{\prec^{\prime}}^{\rm mut}(F)=\max\{d_{F},{\rm max.GB.deg}_{\prec^{\prime}}(F)\},$
+where $d_{F}$ denotes the last fall degree of $F$ defined in [10, Definition 1.5] (originally in [28], [27]).
 By this theorem, if $d_{\rm reg}({\langle F^{\rm top}\rangle})<d_{F}$, the degree of regularity is no longer an upper-bound on the solving degrees $\mathrm{sd}_{\prec}^{\rm mac}(F)$ and $\mathrm{sd}_{\prec}^{\rm mut}(F)$.
 Recently, Salizzoni [40] proved the following theorem:
-Theorem 2.2.4 ([39, Theorem 1.1]).
+Theorem 2.2.4 ([40, Theorem 1.1]).
 With notation as above, we also set $D=d_{\rm reg}(\langle F^{\rm top}\rangle)$, and assume that $D\geq\max\{\deg(f):f\in F\}$.
@@
 On the other hand, Semaev and Tenti proved that the solving degree $\mathrm{sd}_{\prec}^{\mathcal{A}}(F)$ for some algorithm $\mathcal{A}$ is linear in the degree of regularity, if $K$ is a (small) finite field, and if the input system contains polynomials related to the field equations, say $x_{i}^{q}-x_{i}$ for $1\leq i\leq n$:
-Theorem 2.2.5 ([40, Theorem 2.1], [41, Theorem 3.65 & Corollary 3.67]).
+Theorem 2.2.5 ([41, Theorem 2.1], [42, Theorem 3.65 & Corollary 3.67]).
 With notation as above, assume that $K=\mathbb{F}_{q}$, and that $F$ contains $x_{i}^{q}-x_{i}$ for all $1\leq i\leq n$.
@@
 In this subsection, we shall extend the notion of degree of regularity of a homogeneous ideal $I$ of $R$.
-Let $\prec$ be a graded order on the monomials of $R$.
-Recall from Definition 2.1.7 that we set $d_{\rm reg}(I)=\infty$ if $R/I$ is not Artinian, but this is not feasible to analyzing the Gröbner basis computation of $F^{h}$ unless a system defined by $F^{h}$ has no non-trivial root over an algebraic closure.
+Let $\prec$ be a graded ordering on the monomials of $R$.
+Recall from Definition 2.1.7 that we set $d_{\rm reg}(I)=\infty$ if $R/I$ is not Artinian, but this is not feasible to analyzing the Gröbner basis computation of $F^{h}$ unless a system defined by $F^{h}$ has no non-trivial root over the algebraic closure $\overline{K}$ of $K$.
 For the feasibility, we here give an alternative definition for degree of regularity, which is viewed as a generalization of the original definition (Definition 2.1.7):
@@
 For a homogeneous ideal $I$ of $R$, we define its generalized degree of regularity $\widetilde{d}_{\rm reg}(I)$ as follows: If there exists an integer $d_{0}$ such that $\dim_{K}(R/I)_{d}=\dim_{K}(R/I)_{d_{0}}$ for any $d$ with $d\geq d_{0}$, we set $\widetilde{d}_{\rm reg}(I):=\mathrm{min}\{d_{0}:\dim_{K}(R/I)_{d}=\dim_{K}(R/I)_{d_{0}}\mbox{ for any $d$ with $d\geq d_{0}$}\}$, and otherwise we set $\widetilde{d}_{\rm reg}(I):=\infty$.
-The existence of $d_{0}$ in Definition 2.3.1 is equivalent to that the number of projective zeros of $I$ is finite at most.
-Note also that, in Definition 2.3.1, we have $\widetilde{d}_{\rm reg}(I)={d}_{\rm reg}(I)<\infty$ if $R/I$ is Artinian, and otherwise $\widetilde{d}_{\rm reg}(I)<d_{\rm reg}(I)$ or $\widetilde{d}_{\rm reg}(I)={d}_{\rm reg}(I)=\infty$.
+Remark 2.3.2.
+For a homogeneous ideal $I$ in $R$, its generalized degree of regularity $\widetilde{d}_{\rm reg}(I)$ is nothing but the index of regularity $i_{\rm reg}(I)$, if $I$ has finitely many projective zeros over the algebraic closure $\overline{K}$.
+Here, the index of regularity of $I$ is defined as follows: Denoting by ${\rm HP}_{R/I}$ the Hilbert polynomial of $R/I$, we define the index of regularity of $I$ as the smallest non-negative integer $i_{\rm reg}(I)$ such that ${\rm HF}_{R/I}(d)={\rm HP}_{R/I}(d)$ for all $d$ with $d\geq i_{\rm reg}(I)$.
+Note that the index of regularity $i_{\rm reg}(I)$ is also called the Hilbert regularity of $I$, which is often denoted by ${\rm hilb}(I)$.
+The following three conditions are all equivalent: (1) $I$ has finite number of projective zeros over the algebraic closure $\overline{K}$.
+(2) ${\rm HP}_{R/I}$ is a constant.
+(3) $\widetilde{d}_{\rm reg}(I)<\infty$.
+Therefore, we have $\tilde{d}_{\rm reg}(I)=i_{\rm reg}(I)$.
+We also remark that ${\rm HS}_{R/I}$ is written as ${\rm HS}_{R/I}(z)=h(z)/(1-z)$ for a unique polynomial $h(z)$, for which $\widetilde{d}_{\rm reg}(I)=\deg(h)$ and $h(1)={\rm HF}_{R/I}(d)={\rm HP}_{R/I}(d)$ for all $d$ with $d\geq\widetilde{d}_{\rm reg}(I)=i_{\rm reg}(I)$.
+Note that the degree of regularity $d_{\rm reg}(I)$ is also equal to $i_{\rm reg}(I)$ if $R/I$ is Artinian, but these are distinguished in the literature.
+Following this, we distinguish $\tilde{d}_{\rm reg}(I)$ and $i_{\rm reg}(I)$.
+Note also that, in Definition 2.3.1, we have $\widetilde{d}_{\rm reg}(I)={d}_{\rm reg}(I)<\infty$ if $R/I$ is Artinian, and otherwise $\widetilde{d}_{\rm reg}(I)<d_{\rm reg}(I)=\infty$ or $\widetilde{d}_{\rm reg}(I)={d}_{\rm reg}(I)=\infty$.
 We also extend the cryptographic semi-regularity (Definition 2.1.10) of a sequence of homogeneous polynomials, as follows:
@@
 Here, we relate the solving degree of $F^{h}$ (namely the maximal Gröbner basis degree of $F^{h}$) with our generalized degree of regularity, under some assumptions.
+For this, we extend the notion of top part to a homogeneous polynomial in $R^{\prime}=R[y]$ as follows:
+Definition 2.3.4.
+For a homogeneous polynomial $h$ in $R^{\prime}=R[y]$, we call $h|_{y=0}$ the top part of $h$, and denote it by $h^{\rm top}$.
+For a set $H$ of homogeneous polynomials in $R^{\prime}$, its top part is defined by $H^{\rm top}:=\{h^{\rm top}:h\in F\}\subset R$, and similarly $\bm{H}^{\rm top}$ is defined for a sequence $\bm{H}$ of homogeneous polynomials in $R^{\prime}$.
+In Definition 2.3.4, if $h^{\rm top}$ is not zero, it coincides with the top part $(h|_{y=1})^{\rm top}$ of the dehomogenization $h|_{y=1}$ of $h$.
+We remark that $g^{\rm top}=(g^{h})^{\rm top}$ for a polynomial $g$ in $R$.
+First, we prove the following lemma (a generalization of Lemma 2.3.5, see Appendix A.3 below):
+Lemma 2.3.5.
+Let $H=\{h_{1},\ldots,h_{m}\}$ be a set of homogeneous polynomials in $R^{\prime}=R[y]$.
+Assume that $R^{\prime}/\langle H,y\rangle$ (which is isomorphic to $R/\langle H^{\rm top}\rangle$ with $H^{\rm top}=H|_{y=0}$) is Artinian, namely $D:=d_{\rm reg}(\langle H^{\rm top}\rangle_{R})<\infty$.
+We also assume that $D^{\prime}:=\widetilde{d}_{\rm reg}(\langle H\rangle_{R^{\prime}})<\infty$.
+Then, if $D^{\prime}\geq D$, we have $\mathrm{max.GB.deg}_{\prec^{\prime}}(H)\leq D^{\prime}$ for any graded monomial ordering $\prec^{\prime}$ on $R^{\prime}$ given in Lemma 2.2.2.
+Proof.
+As in the proof of Theorem 3.1.1 below (or considering a mapping cone of Koszul complexes), we have the following exact sequence:
+[…]
+for each $d$, where $K_{\bullet}$ (resp.
+$K_{\bullet}^{\prime}$) denotes the Koszul complex on the sequence $(h_{1},\ldots,h_{m})$ (resp.
+the sequence $(h_{1},\ldots,h_{m},y)$).
+It follows from the definition of $d_{\rm reg}$ that $H_{0}(K_{\bullet}^{\prime})_{d}=0$ for any $d$ with $d\geq D$.
+Thus, for any $d$ with $d\geq D$, the multiplication-by-$y$ map $H_{0}(K_{\bullet})_{d-1}\longrightarrow H_{0}(K_{\bullet})_{d}$ is surjective, and it is bijective if and only if $\dim_{K}(R^{\prime}/\langle H\rangle)_{d-1}=\dim_{K}(R^{\prime}/\langle H\rangle)_{d}$.
+Here, for any $d$ with $d\geq D^{\prime}$, the multiplication-by-$y$ map $H_{0}(K_{\bullet})_{d}\longrightarrow H_{0}(K_{\bullet})_{d+1}$ is bijective, since $d+1\geq D^{\prime}+1\geq D$.
+By this together with the surjectivity of $H_{0}(K_{\bullet})_{D^{\prime}-1}\longrightarrow H_{0}(K_{\bullet})_{D^{\prime}}$ (this surjectivity comes from $D^{\prime}\geq D$), it follows from Lemma 2.2.2 that $\mathrm{max.GB.deg}_{\prec^{\prime}}(H)$ is upper-bounded by $D^{\prime}$, as desired.
+∎
 Proposition 2.3.6.
-Let $\bm{F}=(f_{1},\ldots,f_{m})\in(R\smallsetminus K)^{m}$ be a sequence of not necessarily homogeneous polynomials.
-We set $F=\{f_{1},\ldots,f_{m}\}$.
-Assume that $R/\langle F^{\rm top}\rangle$ is Artinian, and that $\bm{F}^{\rm top}$ is cryptographic semi-regular.
+Let $H=\{h_{1},\ldots,h_{m}\}$ be a set of homogeneous polynomials in $R^{\prime}\smallsetminus K$ with $R^{\prime}=R[y]$, and put $\bm{H}:=(h_{1},\ldots,h_{m})\in(R^{\prime}\smallsetminus K)^{m}$.
+Assume that $R/\langle H^{\rm top}\rangle$ is Artinian, and that $\bm{H}^{\rm top}$ is cryptographic semi-regular (and hence $\widetilde{d}_{\rm reg}(\langle H\rangle_{R^{\prime}})<\infty$ by Theorem 3.1.1, see also Remark 3.1.2).
 Then we have the following:
   1. 1.
-    $\widetilde{d}_{\rm reg}(\langle F^{h}\rangle)\geq d_{\rm reg}(\langle F^{\rm top}\rangle)-1$.
+    $\widetilde{d}_{\rm reg}(\langle H\rangle_{R^{\prime}})\geq d_{\rm reg}(\langle H^{\rm top}\rangle_{R})-1$.
   2. 2.
-    $\mathrm{max.GB.deg}_{\prec^{h}}(F^{h})\leq\max\{{d}_{\rm reg}(\langle F^{\rm top}\rangle),\widetilde{d}_{\rm reg}(\langle F^{h}\rangle)\}$ for any graded monomial order $\prec$.
-Moreover, the equality holds in the second inequality if $\langle\mathrm{LM}(\langle F^{h}\rangle)\rangle$ is weakly reverse lexicographic.
-Here, a weakly reverse lexicographic ideal is a monomial ideal $J$ such that if $x^{\alpha}$ is one of the minimal generators of $J$ then every monomial of the same degree which preceeds $x^{\alpha}$ must belong to $J$ as well (see [36, Section 4] for the original definition).
+    $\mathrm{max.GB.deg}_{\prec^{\prime}}(H)\leq\max\{{d}_{\rm reg}(\langle H^{\rm top}\rangle_{R}),\widetilde{d}_{\rm reg}(\langle H\rangle_{R^{\prime}})\}$ for any graded monomial ordering $\prec^{\prime}$ on $R^{\prime}$ given in Lemma 2.2.2.
+Moreover, when we assume that $(H|_{y=1})^{h}=H$, the equality holds in the second inequality if $\prec^{\prime}$ is a DRL ordering and if $\langle\mathrm{LM}(\langle H\rangle)\rangle$ is weakly reverse lexicographic.
+Here, a weakly reverse lexicographic ideal is a monomial ideal $J$ such that if $x^{\alpha}$ is one of the minimal generators of $J$ then every monomial of the same degree which preceeds $x^{\alpha}$ must belong to $J$ as well (see [37, Section 4] for the original definition).
 Proof.
-Put $D:=\widetilde{d}_{\rm reg}(\langle F^{\rm top}\rangle)$ and $D^{\prime}:=\widetilde{d}_{\rm reg}(\langle F^{h}\rangle)$. By the proof of Theorem 3.1.1 below, we have the following exact sequence:
-[…]
-for each $d$, where $K_{\bullet}$ denotes the Koszul complex on $\bm{F}^{h}$.
-Here we have $H_{1}(K_{\bullet}\otimes_{R^{\prime}}R)_{d}=0$ and $H_{0}(K_{\bullet}\otimes_{R^{\prime}}R)_{d}\neq 0$ for $d\leq D-1$, whence $\dim_{K}(R^{\prime}/\langle F^{h}\rangle)_{d-1}<\dim_{K}(R^{\prime}/\langle F^{h}\rangle)_{d}$ for any such $d$.
-Therefore, the first argument to be proved holds.
-On the other hand, it follows from the definition of $d_{\rm reg}$ that $H_{0}(K_{\bullet}\otimes_{R^{\prime}}R)_{d}=0$ for any $d$ with $d\geq D$.
-Thus, for such any $d$, the multiplication-by-$y$ map $H_{0}(K_{\bullet})_{d-1}\longrightarrow H_{0}(K_{\bullet})_{d}$ is surjective, and it is bijective if and only if $\dim_{K}(R^{\prime}/\langle F^{h}\rangle)_{d-1}=\dim_{K}(R^{\prime}/\langle F^{h}\rangle)_{d}$.
-If $D^{\prime}\geq D$, for any $d$ with $d\geq D^{\prime}$, the multiplication-by-$y$ map $H_{0}(K_{\bullet})_{d}\longrightarrow H_{0}(K_{\bullet})_{d+1}$ is bijective, since $d+1\geq D^{\prime}+1\geq D$.
-By this together with the surjectivity of $H_{0}(K_{\bullet})_{D^{\prime}-1}\longrightarrow H_{0}(K_{\bullet})_{D^{\prime}}$, it follows from Lemma 2.2.2 that $\mathrm{max.GB.deg}_{\prec^{h}}(F^{h})$ is upper-bounded by $D^{\prime}=\max\{D,D^{\prime}\}$.
-Also if $D^{\prime}=D-1$, for any $d$ with $d\geq D^{\prime}$, the multiplication-by-$y$ map $H_{0}(K_{\bullet})_{d}\longrightarrow H_{0}(K_{\bullet})_{d+1}$ is bijective, since $d+1\geq D^{\prime}+1=D$.
+Put $D:=\widetilde{d}_{\rm reg}(\langle H^{\rm top}\rangle_{R})$ and $D^{\prime}:=\widetilde{d}_{\rm reg}(\langle H\rangle_{R^{\prime}})$.
+Recall from the proof of Lemma 2.3.5 that we have the exact sequence (2.3.1).
+For any $d$ with $d\leq D-1$, we have $H_{0}(K_{\bullet}^{\prime})_{d}\neq 0$, and it follows from the cryptographic semi-regularity of $\bm{H}^{\rm top}$ that $H_{1}(K_{\bullet}^{\prime})_{d}=0$ by Proposition 2.1.11, whence $\dim_{K}(R^{\prime}/\langle H\rangle)_{d-1}<\dim_{K}(R^{\prime}/\langle H\rangle)_{d}$ for any such $d$.
+Therefore, the first assertion to be proved holds.
+As in the proof of Lemma 2.3.5, it follows from the definition of $d_{\rm reg}$ that $H_{0}(K_{\bullet}^{\prime})_{d}=0$ for any $d$ with $d\geq D$.
+Thus, for such any $d$, the multiplication-by-$y$ map $H_{0}(K_{\bullet})_{d-1}\longrightarrow H_{0}(K_{\bullet})_{d}$ is surjective, and it is bijective if and only if $\dim_{K}(R^{\prime}/\langle H\rangle)_{d-1}=\dim_{K}(R^{\prime}/\langle H\rangle)_{d}$.
+Here, we prove the second assertion $\mathrm{max.GB.deg}_{\prec^{\prime}}(H)\leq\max\{D,D^{\prime}\}$.
+If $D^{\prime}\geq D$, the assertion is a consequence of Lemma 2.3.5.
+Thus, it suffices to consider the case where $D^{\prime}=D-1$.
+In this case, for any $d$ with $d\geq D^{\prime}$, the multiplication-by-$y$ map $H_{0}(K_{\bullet})_{d}\longrightarrow H_{0}(K_{\bullet})_{d+1}$ is bijective, since $d+1\geq D^{\prime}+1=D$.
 However, $H_{0}(K_{\bullet})_{D^{\prime}-1}\longrightarrow H_{0}(K_{\bullet})_{D^{\prime}}$ is injective but not surjective, we cannot apply Lemma 2.2.2 for $d_{0}=D^{\prime}$, but apply it for $d_{0}=D^{\prime}+1=D$.
-Hence, $\mathrm{max.GB.deg}_{\prec^{h}}(F^{h})$ is upper-bounded by $D^{\prime}+1=D=\max\{D,D^{\prime}\}$, as desired.
-Assume that $\langle\mathrm{LM}(\langle F^{h}\rangle)\rangle$ is a weakly reverse lexicographic ideal.
-First we consider the case where $\widetilde{d}_{\rm reg}(\langle F^{h}\rangle)\geq d_{\rm reg}(\langle F^{\rm top}\rangle)$, so that $\max\{{d}_{\rm reg}(\langle F^{\rm top}\rangle),\widetilde{d}_{\rm reg}(\langle F^{h}\rangle)\}=\widetilde{d}_{\rm reg}(\langle F^{h}\rangle)$.
-We put $d:=\mathrm{max.GB.deg}_{\prec^{h}}(F^{h})$, and let $\{t_{1},\ldots,t_{r}\}$ be a standard monomial basis of $R^{\prime}_{d}/\langle F^{h}\rangle_{d}$, that is,
-$\{t\in R^{\prime}_{d}:\mbox{$t$ is a monomial and }{t\notin\langle{\rm LM}(\langle F^{h}\rangle)\rangle}\}.$
-For $s\geq 1$, we prove that $\{t_{1}y^{s},\ldots,t_{r}y^{s}\}$ is a basis of $R^{\prime}_{d+s}/\langle F^{h}\rangle_{d+s}$, from which we have $\dim_{K}R^{\prime}_{d}/\langle F^{h}\rangle_{d}=\dim_{K}R^{\prime}_{d+s}/\langle F^{h}\rangle_{d+s}$ for any positive integer $s$, so that $d\geq\widetilde{d}_{\rm reg}(\langle F^{h}\rangle)$ and therefore $d=\widetilde{d}_{\rm reg}(\langle F^{h}\rangle)$.
-By Remark 3.1.2 and Lemma 2.3.5 below, the multiplication-by-$y^{s}$ map from $R_{d}^{\prime}/\langle F^{h}\rangle_{d}$ to $R^{\prime}_{d+s}/\langle F^{h}\rangle_{d+s}$ is surjective.
-Therefore $B_{s}:=\{t_{1}y^{s},\ldots,t_{r}y^{s}\}$ generates $R^{\prime}_{d+s}/\langle F^{h}\rangle_{d+s}$.
+Hence, $\mathrm{max.GB.deg}_{\prec^{\prime}}(H)$ is upper-bounded by $D^{\prime}+1=D=\max\{D,D^{\prime}\}$, as desired.
+Now, assuming that $(H|_{y=1})^{h}=H$, we show that the equality holds in the second inequality if $\prec^{\prime}$ is a DRL ordering and if $\langle\mathrm{LM}(\langle H\rangle)\rangle$ is weakly reverse lexicographic.
+Let $G_{H}$ be the reduced Gröbner basis of $\langle H\rangle_{R^{\prime}}$ with respect to $\prec^{\prime}$.
+Note that the elements of $\mathrm{LM}(G_{H})$ are the minimal generators of the monomial ideal $\langle\mathrm{LM}(\langle H\rangle)\rangle=\langle\mathrm{LM}(G_{H})\rangle$.
+First we consider the case where $D^{\prime}\geq D$, so that $\max\{D,D^{\prime}\}=D^{\prime}$.
+Put $d:=\mathrm{max.GB.deg}_{\prec^{\prime}}(H)$.
+Then, it follows from Macaulay’s basis theorem (cf.
+[29, Theorem 1.5.7]) that $R^{\prime}_{d}/\langle H\rangle_{d}$ has a $K$-linear basis of the form
+$\{t\in R^{\prime}_{d}:\mbox{$t$ is a monomial and }{t\notin\langle{\rm LM}(\langle H\rangle)\rangle=\langle\mathrm{LM}(G_{H})\rangle}\},$
+which is called the standard monomial basis.
+We set $r:=\dim_{K}R^{\prime}_{d}/\langle H\rangle_{d}$, and write this standard monomial basis as $\{t_{1},\ldots,t_{r}\}$.
+In the following, we prove that $\{t_{1}y^{s},\ldots,t_{r}y^{s}\}$ is a $K$-linear basis of $R^{\prime}_{d+s}/\langle H\rangle_{d+s}$ for any $s$ with $s\geq 1$, from which we have $\dim_{K}R^{\prime}_{d}/\langle H\rangle_{d}=\dim_{K}R^{\prime}_{d+s}/\langle H\rangle_{d+s}$ for any positive integer $s$, so that $d\geq D^{\prime}$ and therefore $d=D^{\prime}$.
+By Remark 3.1.2 and Lemma 2.3.8 below, the multiplication-by-$y^{s}$ map from $R_{d}^{\prime}/\langle H\rangle_{d}$ to $R^{\prime}_{d+s}/\langle H\rangle_{d+s}$ is surjective.
+Therefore $B_{s}:=\{t_{1}y^{s},\ldots,t_{r}y^{s}\}$ generates $R^{\prime}_{d+s}/\langle H\rangle_{d+s}$.
 Suppose to contrary that $B_{s}$ is not a basis.
-In this case, there exists an $i$ such that $t_{i}y^{s}$ is divisible by $\mathrm{LM}(g)$ for some $g\in G_{\rm hom}$, but $t_{i}$ is not divisible by $\mathrm{LM}(g)$, where $G_{\rm hom}$ is the reduced Gröbner basis of $\langle F^{h}\rangle$ with respect to $\prec^{h}$.
-Therefore, we can write $t_{i}y^{s}=s_{i}u_{i}y^{s}$ with $s_{i}u_{i}=t_{i}$ and $u_{i}y^{s^{\prime}}=\mathrm{LM}(g)$ for some $s^{\prime}\leq s$.
-(We note that $\deg s_{i}=d-\deg u_{i}\geq\deg{\rm LM}(g)-\deg u_{i}=s^{\prime}$ and $u_{i}={\rm GCD}(t_{i},{\rm LM}(g))$.) Note that $s^{\prime}\geq 1$ since otherwise $t_{i}$ is divisible by $\mathrm{LM}(g)$.
+In this case, there exists an $i$ such that $t_{i}y^{s}$ is divisible by $\mathrm{LM}(g)$ for some $g\in G_{H}$.
+Putting $u_{i}=\mathrm{GCD}(t_{i},\mathrm{LM}(g))$ and $s_{i}=t_{i}/u_{i}$, we have $u_{i}s_{i}y^{s}=t_{i}y^{s}$.
+Since $t_{i}=s_{i}u_{i}$ is not divisible by $\mathrm{LM}(g)$, we can write $u_{i}y^{s^{\prime}}=\mathrm{LM}(g)$ for some $s^{\prime}\leq s$.
+(We note that $\deg s_{i}=d-\deg u_{i}\geq\deg{\rm LM}(g)-\deg u_{i}=s^{\prime}$.) Note that $s^{\prime}\geq 1$ since otherwise $t_{i}$ is divisible by $\mathrm{LM}(g)$.
 Take an arbitrary monomial $s^{\prime}_{i}$ such that $\deg s^{\prime}_{i}=s^{\prime}$ and $s_{i}^{\prime}$ divides $s_{i}$.
-Then, by the weakly reverse lexicographicness, as ${\rm LM}(g)=u_{i}y^{s^{\prime}}\preceq u_{i}s^{\prime}_{i}$, the monomial $s^{\prime}_{i}u_{i}$ should belong to $\langle{\rm LM}(G_{\rm hom})\rangle$.
-Moreover, since $s^{\prime}_{i}u_{i}$ divides $t_{i}=s_{i}u_{i}$, the monomial $t_{i}=s_{i}u_{i}$ also belongs to $\langle{\rm LM}(G_{\rm hom})\rangle$, which is a contradiction.
-Finally, we consider the remaining case, namely we have $d_{\rm reg}(\langle F^{\rm top}\rangle)>\widetilde{d}_{\rm reg}(\langle F^{h}\rangle)$, so that $\max\{{d}_{\rm reg}(\langle F^{\rm top}\rangle),\widetilde{d}_{\rm reg}(\langle F^{h}\rangle)\}=d_{\rm reg}(\langle F^{\rm top}\rangle)$.
-In this case, it follows from Lemma 2.3.5 below that $d\geq d_{\rm reg}(\langle F^{\rm top}\rangle)$, whence $d=d_{\rm reg}(\langle F^{\rm top}\rangle)$.
-∎
+Then, by the weakly reverse lexicographicness, as ${\rm LM}(g)=u_{i}y^{s^{\prime}}\preceq u_{i}s^{\prime}_{i}$, the monomial $s^{\prime}_{i}u_{i}$ should belong to $\langle{\rm LM}(G_{H})\rangle$.
+Moreover, since $s^{\prime}_{i}u_{i}$ divides $t_{i}=s_{i}u_{i}$, the monomial $t_{i}=s_{i}u_{i}$ also belongs to $\langle{\rm LM}(G_{H})\rangle$, which is a contradiction.
+Finally, we consider the remaining case, namely we have $D>D^{\prime}$, so that $\max\{D,D^{\prime}\}=D$. In this case, it follows from Lemma 2.3.8 below that $d\geq D$, whence $d=D$. ∎
 Remark 2.3.7.
@@
 $\cdots<{\rm HF}_{A^{\prime}}(D-2)<{\rm HF}_{A^{\prime}}(D-1)\geq{\rm HF}_{A^{\prime}}(D)\geq\cdots\geq{\rm HF}_{A^{\prime}}(D^{\prime})={\rm HF}_{A^{\prime}}(D^{\prime}+1)=\cdots$
-with $A^{\prime}:=R^{\prime}/\langle F^{h}\rangle$, and the case where $D^{\prime}=D-1$ (i.e., $D^{\prime}+1=D$) means that
+with $A^{\prime}:=R^{\prime}/\langle H\rangle$, and the case where $D^{\prime}=D-1$ (i.e., $D^{\prime}+1=D$) means that
 $\cdots<{\rm HF}_{A^{\prime}}(D-2)<{\rm HF}_{A^{\prime}}(D-1)={\rm HF}_{A^{\prime}}(D)=\cdots.$
 Lemma 2.3.8.
-Let $\bm{F}=(f_{1},\ldots,f_{m})\in(R\smallsetminus K)^{m}$ be a sequence of not necessarily homogeneous polynomials.
-We set $F=\{f_{1},\ldots,f_{m}\}$, and assume that $R/\langle F^{\rm top}\rangle$ is Artinian and that $\bm{F}^{\rm top}$ is cryptographic semi-regular.
-We also suppose that $\langle\mathrm{LM}(\langle F^{h}\rangle)\rangle$ is weakly reverse lexicographic.
-Then we have $\mathrm{max.GB.deg}_{\prec^{h}}(F^{h})\geq d_{\rm reg}(\langle F^{\rm top}\rangle)$ and $\mathrm{max.GB.deg}_{\prec}(F^{\rm top})=d_{\rm reg}(\langle F^{\rm top}\rangle)$.
+Let $H$, $\bm{H}$, and $\prec^{\prime}$ be as in Proposition 2.3.6.
+Assume that $R/\langle H^{\rm top}\rangle$ is Artinian, and that $\bm{H}^{\rm top}$ is cryptographic semi-regular.
+We also suppose that $\prec^{\prime}$ is a DRL ordering and that $\langle\mathrm{LM}(\langle H\rangle)\rangle$ is weakly reverse lexicographic.
+Then we have $\mathrm{max.GB.deg}_{\prec^{\prime}}(H)\geq d_{\rm reg}(\langle H^{\rm top}\rangle_{R})$ and $\mathrm{max.GB.deg}_{\prec}(H^{\rm top})=d_{\rm reg}(\langle H^{\rm top}\rangle_{R})$, where $\prec$ denotes the restriction of $\prec^{\prime}$ to $R$.
 Proof.
+Note that $\prec^{\prime}$ is the homogenization $\prec^{h}$ of $\prec$.
+We set $f_{j}:=(h_{j})|_{y=1}$ for each $j$ with $1\leq j\leq m$, and put $F:=H|_{y=1}=\{f_{1},\ldots,f_{m}\}$ and $\bm{F}:=\bm{H}|_{y=1}=(f_{1},\ldots,f_{m})$, so that $H=F^{h}$, $\bm{H}=\bm{F}^{h}$, $H^{\rm top}=F^{\rm top}$, and $\bm{H}^{\rm top}=\bm{F}^{\rm top}$ by our assumption $(H|_{y=1})^{h}=H$.
+With these notations, the assertion to be proved is that we have $\mathrm{max.GB.deg}_{\prec^{h}}(F^{h})\geq d_{\rm reg}(\langle F^{\rm top}\rangle_{R})$ and $\mathrm{max.GB.deg}_{\prec}(F^{\rm top})=d_{\rm reg}(\langle F^{\rm top}\rangle_{R})$.
+In the following, let $G_{\rm hom}$ (resp.
+$G_{\rm top}$) denote the reduced Gröbner basis of $\langle F^{h}\rangle_{R^{\prime}}$ (resp.
+$\langle F^{\rm top}\rangle_{R}$) with respect to $\prec^{h}$ (resp.
+$\prec$).
 Put $d:=\mathrm{max.GB.deg}_{\prec^{h}}(F^{h})$ and assume for a contradiction that $d<D:=d_{\rm reg}(\langle F^{\rm top}\rangle)$.
 By Lemma 4.1.4 below, we have ${\rm LM}(G_{\rm hom})={\rm LM}(G_{\rm top})_{\leq d}\subset R$.
-Then $\langle{\rm LM}(G_{\rm hom})\rangle_{R}$ contains ${\rm LM}(G_{\rm top})$.
+Also, $\langle{\rm LM}(G_{\rm hom})\rangle_{R}$ contains ${\rm LM}(G_{\rm top})$.
 This can be shown as follows: For any $t\in{\rm LM}(G_{\rm top})$, there are polynomials $a_{1},\ldots,a_{m}$ in $R$ such that $t={\rm LM}(\sum_{i=1}^{m}a_{i}f_{i}^{\rm top})$.
@@
 Then, for any monomial $t\in R_{d^{\prime}}$, as $t\succeq x_{n}^{d^{\prime}}$, it belongs to $\langle{\rm LM}(\langle F^{h}\rangle)\rangle_{R^{\prime}}(=\langle{\rm LM}(G_{\rm hom})\rangle_{R^{\prime}}=\langle{\rm LM}(G_{\rm top})_{\leq d}\rangle_{R^{\prime}})$ by its weakly reverse lexicographicness.
-This implies that $t$ also belongs to $\langle{\rm LM}(G_{\rm top})\rangle_{R}$, and hence $R_{d}/\langle F^{\rm top}\rangle_{d}=0$.
+This implies that $t$ also belongs to $\langle{\rm LM}(G_{\rm top})\rangle_{R}$, and hence $R_{d^{\prime}}/\langle F^{\rm top}\rangle_{d^{\prime}}=0$.
 Thus, we have $d\geq d^{\prime}\geq d_{\rm reg}(\langle F^{\rm top}\rangle)$, a contradiction.
 By the same argument as above, we can show that $\mathrm{max.GB.deg}_{\prec}(F^{\rm top})=d_{\rm reg}(\langle F^{\rm top}\rangle)$ as follows.
-It is obvious that $\mathrm{max.GB.deg}_{\prec}(F^{\rm top})\leq d_{\rm reg}(\langle F^{\rm top}\rangle)$.
+Recall from Lemma 2.1.8 that $\mathrm{max.GB.deg}_{\prec}(F^{\rm top})\leq d_{\rm reg}(\langle F^{\rm top}\rangle)$.
 Thus, we assume to the contrary that $d:=\mathrm{max.GB.deg}_{\prec}(F^{\rm top})<D:=d_{\rm reg}(\langle F^{\rm top}\rangle)$.
@@
 Remark 2.3.9.
-In Lemma 2.3.5, when $R/\langle F^{\rm top}\rangle$ is Artinian, we can easily prove the equality $\mathrm{max.GB.deg}_{\prec}(F^{\rm top})=d_{\rm reg}(\langle F^{\rm top}\rangle)$ if $\langle\mathrm{LM}(\langle F^{\rm top}\rangle)\rangle$ is weakly reverse lexicographic, not assuming that $\bm{F}^{\rm top}$ is cryptographic semi-regular nor that $\langle\mathrm{LM}(\langle F^{h}\rangle)\rangle$ is weakly reverse lexicographic.
+In Lemma 2.3.8, when $R/\langle H^{\rm top}\rangle$ is Artinian, we can easily prove the equality $\mathrm{max.GB.deg}_{\prec}(H^{\rm top})=d_{\rm reg}(\langle H^{\rm top}\rangle)$ if $\langle\mathrm{LM}(\langle H^{\rm top}\rangle)\rangle$ is weakly reverse lexicographic, not assuming that $\bm{H}^{\rm top}$ is cryptographic semi-regular nor that $\langle\mathrm{LM}(\langle H\rangle)\rangle$ is weakly reverse lexicographic.
 3 Proofs of Theorems 1 and 2
@@
 3.1 Bounded regularity of homogenized sequences
-Here we revisit the main results in [30, Section 4].
+Here we revisit the main results in [31, Section 4].
 For the readability, we remain the proofs.
-Also, as additional remarks, we explicitly give an important property of the Hilbert-Poincaré series of $R^{\prime}/\langle F^{h}\rangle$ with $R^{\prime}=R[y]$, and also give an alternative proof for [30, Theoem 7] (Theorem 3.1.1 below).
+Also, as additional remarks, we explicitly give an important property of the Hilbert-Poincaré series of $R^{\prime}/\langle F^{h}\rangle$ with $R^{\prime}=R[y]$, and also give an alternative proof for [31, Theoem 7] (Theorem 3.1.1 below).
 The Hilbert-Poincaré series associated to a (homogeneous) cryptographic semi-regular sequence is given by (2.1.2).
 On the other hand, the Hilbert-Poincaré series associated to the homogenizaton $F^{h}$ cannot be computed without knowing its Gröbner basis in general, but we shall prove that it can be computed up to the degree $d_{\rm reg}(\langle F^{\rm top}\rangle)-1$ if $\bm{F}$ is affine cryptographic semi-regular, namely $\bm{F}^{\rm top}=(f_{1}^{\rm top},\ldots,f_{m}^{\rm top})$ is cryptographic semi-regular.
-Theorem 3.1.1 (Theorem 1 (1); [30, Theoem 7]).
+Theorem 3.1.1 (Theorem 1 (1); [31, Theoem 7]).
 Let $R^{\prime}=R[y]$, and let $\bm{F}=(f_{1},\ldots,f_{m})$ be a sequence of not necessarily homogeneous polynomials in $R$ of positive degrees.
@@
 Then, for each $d$ with $d<D:=d_{\mathrm{reg}}(\langle F^{\rm top}\rangle)$, we have
-${\rm HF}_{R^{\prime}/\langle F^{h}\rangle}(d)={\rm HF}_{R/\langle F^{\rm top}\rangle}(d)+{\rm HF}_{{\color[rgb]{0,0,0}R^{\prime}}/\langle F^{h}\rangle}(d-1),$ (3.1.1)
+${\rm HF}_{R^{\prime}/\langle F^{h}\rangle}(d)={\rm HF}_{R/\langle F^{\rm top}\rangle}(d)+{\rm HF}_{R^{\prime}/\langle F^{h}\rangle}(d-1),$ (3.1.1)
 and hence
@@
 Note that there is an isomorphism
-$K_{i}\otimes_{R^{\prime}}R\cong\bigoplus_{1\leq j_{1}<\cdots<j_{i}\leq m}R(-d_{j_{1}\cdots j_{i}})\mathbf{e}_{j_{1}\cdots{\color[rgb]{0,0,0}{j_{i}}}},$
+$K_{i}\otimes_{R^{\prime}}R\cong\bigoplus_{1\leq j_{1}<\cdots<j_{i}\leq m}R(-d_{j_{1}\cdots j_{i}})\mathbf{e}_{j_{1}\cdots j_{i}},$
 via which we can interpret $\pi_{i}:K_{i}\to K_{i}\otimes_{R^{\prime}}R$ as a homomorphism that projects each entry of a vector in $K_{i}$ modulo $y$. In particular, we have
@@
 by the dimension theorem.
-Since $H_{0}(K_{\bullet})=R^{\prime}/\langle F^{h}\rangle$ and ${\color[rgb]{0,0,0}H_{0}(K_{\bullet}\otimes_{R^{\prime}}R)}=R/\langle F^{\rm top}\rangle$, we have the equality (3.1.1), as desired.
+Since $H_{0}(K_{\bullet})=R^{\prime}/\langle F^{h}\rangle$ and $H_{0}(K_{\bullet}\otimes_{R^{\prime}}R)=R/\langle F^{\rm top}\rangle$, we have the equality (3.1.1), as desired.
 ∎
-Remark 3.1.2 (Theorem 1 (2), (3); [30, Remark 6]).
+Remark 3.1.2 (Theorem 1 (2), (3); [31, Remark 6]).
 Note that, in the proof of Theorem 3.1.1, the multiplication map $H_{0}(K_{\bullet})_{d-1}\to H_{0}(K_{\bullet})_{d}$ by $y$ is injective for all $d<D$, whence ${\rm HF}_{R^{\prime}/\langle F^{h}\rangle}(d)$ is monotonically increasing for $d<D$.
 On the other hand, since $H_{0}(K_{\bullet}\otimes_{R^{\prime}}R)_{d}=(R/\langle F^{\rm top}\rangle)_{d}=0$ for all $d\geq D$ by the definition of the degree of regularity, the multiplication map $H_{0}(K_{\bullet})_{d-1}\to H_{0}(K_{\bullet})_{d}$ by $y$ is surjective for all $d\geq D$, whence ${\rm HF}_{R^{\prime}/\langle F^{h}\rangle}(d)$ is monotonically decreasing for $d\geq D-1$.
-By this together with [10, Theorem 3.3.4], the homogeneous ideal $\langle F^{h}\rangle$ is zero-dimensional or trivial, i.e., there are at most a finite number of projective zeros of $F^{h}$ (and thus there are at most a finite number of affine zeros of $F$).
-Remark 3.1.3.
-We note that, for each $d\geq D$, the condition $\dim_{K}H_{0}(K_{\bullet})_{d-1}\!=\!\dim_{K}H_{0}(K_{\bullet})_{d}$ is equivalent to that the multiplication map $H_{0}(K_{\bullet})_{d-1}\to H_{0}(K_{\bullet})_{d}$ by $y$ is injective (and thus bijective).
-By this together with Lemma 2.2.2, letting $d_{0}$ be the smallest number with $d_{0}\geq D$ such that $\dim_{K}H_{0}(K_{\bullet})_{d_{0}}=\dim_{K}H_{0}(K_{\bullet})_{d_{0}+1}$, the maximal Gröbner basis degree of $\langle F^{h}\rangle$ is upper-bounded by $d_{0}$.
+By this together with [11, Theorem 3.3.4], the homogeneous ideal $\langle F^{h}\rangle$ is zero-dimensional or trivial, i.e., there are at most a finite number of projective zeros of $F^{h}$ (and thus there are at most a finite number of affine zeros of $F$).
 Remark 3.1.3.
@@
 ${\rm HF}_{R^{\prime}/\langle F^{h}\rangle}(d)={\rm HF}_{R^{\prime}/(\langle F^{h}\rangle+\langle y\rangle)}(d)+{\rm HF}_{R^{\prime}/(\langle F^{h}\rangle:\langle y\rangle)}(d-1),$
-see [25, Lemmas 5.2.1 and 5.2.2].
+see [26, Lemmas 5.2.1 and 5.2.2].
 Note that ${\rm HF}_{R^{\prime}/(\langle F^{h}\rangle+\langle y\rangle)}(d)={\rm HF}_{R/\langle F^{\rm top}\rangle}(d)$ for any positive integer $d$.
@@
 The Hilbert-Poincaré series of $R^{\prime}/\langle F^{h}\rangle$ satisfies the following equality (3.1.6):
-Corollary 3.1.5 (Theorem 1 (3); [30, Corollary 1]).
+Corollary 3.1.4 (Theorem 1 (3); [31, Corollary 1]).
 Let $D=d_{\rm reg}(\langle F^{\rm top}\rangle)$. Then we have
 ${\rm HS}_{R^{\prime}/\langle F^{h}\rangle}(z)\equiv\frac{\prod_{i=1}^{m}(1-z^{d_{i}})}{(1-z)^{n+1}}\pmod{z^{D}}.$ (3.1.6)
-Therefore, by Theorem 2.1.4 ([14, Theorem 1]), the sequence $\bm{F}^{h}$ is $D$-regular.
+Therefore, by Theorem 2.1.4 ([15, Theorem 1]), the sequence $\bm{F}^{h}$ is $D$-regular.
 Here, we note that $D=\deg({\rm HS}_{R/\langle F^{\rm top}\rangle}{)}+1={\deg\left(\left[\frac{\prod_{i=1}^{m}(1-z^{d_{i}})}{(1-z)^{n}}\right]\right)}+1$.
@@
 From now on, we assume that $m\geq n$, and set $\bm{F}_{k}:=(f_{1},\ldots,f_{n+k})$ and $D_{k}:=d_{\rm reg}(\langle F_{k}^{\rm top}\rangle)$ for each $k\geq 0$.
-As $\bm{F}_{0}^{\rm top}$ is regular and $\bm{F}_{1}^{\rm top}$ is semi-regular, we have $D_{0}=d_{1}+\cdots+d_{n}-n+1$ and $D_{1}=\lfloor\frac{d_{1}+\cdots+d_{n+1}-n-1}{2}\rfloor+1$, see [4, Theorem 4.1].
+As $\bm{F}_{0}^{\rm top}$ is regular and $\bm{F}_{1}^{\rm top}$ is semi-regular, we have $D_{0}=d_{1}+\cdots+d_{n}-n+1$ and $D_{1}=\lfloor\frac{d_{1}+\cdots+d_{n+1}-n-1}{2}\rfloor+1$, see [5, Theorem 4.1].
 Thus, by setting $d_{1}\leq d_{2}\leq\cdots\leq d_{m}$, we can minimize the values $D_{0}$ and $D_{1}$.
@@
 Finally in this subsection, under the assumption that $\bm{F}^{\rm top}$ is cryptographic semi-regular, we show that the generalized degree of regularity of $\langle F^{h}\rangle$ (and thus solving degree of $F^{h}$) can be bounded by $D$ plus the saturation exponent, say $S_{0}$ here, that is, the minimal integer $k$ such that $\langle F\rangle^{h}=(\langle F^{h}\rangle:y^{\infty})=(\langle F^{h}\rangle:y^{k})$.
-See [25, p.
+See [26, p.
 81] for the definition of saturation exponent.
 Proposition 3.2.5.
-The generalized degree of regularity $\langle F^{h}\rangle$ and the solving degree of $F^{h}$ are both bounded by $D+S_{0}$.
+The generalized degree of regularity of $\langle F^{h}\rangle$ and the solving degree of $F^{h}$ are both bounded by $D+S_{0}$.
 Proof.
@@
 Then, it follows from Macaulay’s basis theorem (cf.
-[28, Theorem 1.5.7]) that
+[29, Theorem 1.5.7]) that
 $LB_{d}:=\{t\in R^{\prime}_{d}:\mbox{$t$ is a monomial and }{t\notin\langle{\rm LM}(\langle F^{h}\rangle+\langle y^{S_{0}}\rangle)\rangle}\}$
@@
 4 Behaviors of Gröbner bases computation
+We use the same notation as in Section 3.
 Here we show certain correspondences in the Gröbner basis computations among inputs $F^{h}$, $F^{\rm top}$, and $F$.
-First we revisit the correspondence among the computation of the Gröbner basis of $F^{h}$ and that of $F^{\rm top}$ given in [30, Section 5.1].
+First we revisit the correspondence among the computation of the Gröbner basis of $F^{h}$ and that of $F^{\rm top}$ given in [31, Section 5.1].
 Then, we explicitly give an important correspondence between the computation of the Gröbner basis of $F^{h}$ and that of $F$, which brings an upper-bound (Lemma 4.2.4 below) on the solving degree of $F$ related to Samaev-Tenti’s bound [41].
+Subsequently, we observe actual behavior of Gröbner basis computation from a part of our experimental results.
+From our observation, we estimate the complexity of the Gröbner basis computation for $F^{h}$ (and $F$), and will raise a conjecture and some questions related to the generalized cryptographic semi-regularity of $\bm{F}^{h}$.
 Here we use the same notation as in the previous section, and unless otherwise noted, assume that $\bm{F}$ is cryptographic semi-regular.
-Let $G$, $G_{\rm hom}$, and $G_{\rm top}$ be the reduced Gröbner bases of $\langle F\rangle$, $\langle F^{h}\rangle$, and $\langle F^{\rm top}\rangle$, respectively, where their monomial orderings are DRL $\prec$ or its extension $\prec^{h}$.
+Let $G$, $G_{\rm hom}$, and $G_{\rm top}$ be the reduced Gröbner bases of $\langle F\rangle$, $\langle F^{h}\rangle$, and $\langle F^{\rm top}\rangle$, respectively, where their monomial orderings are DRL $\prec$ or its homogenization $\prec^{h}$.
 Also we let $D=d_{\rm reg}(\langle F^{\rm top}\rangle)$, and assume $D<\infty$.
-Moreover, we extend the notion of top part to a homogeneous polynomial $h$ in $R^{\prime}=R[y]$ as follows.
-We call $h|_{y=0}$ the top part of $h$ and denote it by $h^{\rm top}$.
-Thus, if $h^{\rm top}$ is not zero, it coincides with the top part $(h|_{y=1})^{\rm top}$ of the dehomogenization $h|_{y=1}$ of $h$.
-We remark that $g^{\rm top}=(g^{h})^{\rm top}$ for a polynomial $g$ in $R$.
+Moreover, we use the notion of top part to a homogeneous polynomial $h$ in $R^{\prime}=R[y]$, see Definition 2.3.4.
 4.1 Correspondence between $G_{\rm hom}$ and $G_{\rm top}$
-Here we revisit the results in [30, Section 5.1].
-Corollary 4.1.1 ([30, Corollary 2]).
+Here we revisit the results in [31, Section 5.1].
+Corollary 4.1.1 ([31, Corollary 2]).
 With notation as above, assume that $\bm{F}=(f_{1},\ldots,f_{m})\in R^{m}$ is affine cryptographic semi-regular.
@@
 Then, we have $(\langle\mathrm{LM}(\tilde{I})\rangle_{R^{\prime}})_{d}=(\langle\mathrm{LM}(\overline{I})\rangle_{R^{\prime}})_{d}$ for each $d$ with $d<D:=d_{\rm reg}(\overline{I})$.
-Since $\bm{F}^{\rm top}$ is cryptographic semi-regular and since $\bm{F}^{h}$ is $D$-regular by Corollary 3.1.5, we obtain $H_{1}(K_{\bullet}(F^{\rm top}))_{{\color[rgb]{0,0,0}<D}}=H_{1}(K_{\bullet}(F^{h}))_{{\color[rgb]{0,0,0}<D}}=0$.
+Since $\bm{F}^{\rm top}$ is cryptographic semi-regular and since $\bm{F}^{h}$ is $D$-regular by Corollary 3.1.4, we obtain $H_{1}(K_{\bullet}(F^{\rm top}))_{<D}=H_{1}(K_{\bullet}(F^{h}))_{<D}=0$.
 Moreover, as $H_{1}(K_{\bullet}(F^{h}))={\rm syz}(F^{h})/{\rm tsyz}(F^{h})$ and $H_{1}(K_{\bullet}(F^{\rm top}))={\rm syz}(F^{\rm top})/{\rm tsyz}(F^{h})$ (see (A.1.2)), we have the following corollary, where ${\rm tsyz}$ denotes the module of trivial syzygies (see Definition A.1.1).
-Corollary 4.1.2 ([14, Theorem 1]).
+Corollary 4.1.2 ([15, Theorem 1]).
 With notation as above, we have ${\rm syz}(F^{\rm top})_{<D}={\rm tsyz}(F^{\rm top})_{<D}$ and ${\rm syz}(F^{h})_{<D}={\rm tsyz}(F^{h})_{<D}$.
 Remark 4.1.3.
-[…]
+Corollary 4.1.2 implies that, in the Gröbner basis computation $G_{\rm hom}$ with respect to a graded ordering $\prec^{h}$, if an S-polynomial $S(g_{1},g_{2})=t_{1}g_{1}-t_{2}g_{2}$ of degree less than $D$ is reduced to 00, it comes from some trivial syzyzy, that is, $\sum_{i=1}^{m}(t_{1}a^{(1)}_{i}-t_{2}a^{(2)}_{i}-b_{i})\mathbf{e}_{i}$ belongs to ${\rm tsyz}(F^{h})_{<D}$, where $g_{1}=\sum_{i=1}^{m}a^{(1)}_{i}f_{i}^{h}$, $g_{2}=\sum_{i=1}^{m}a^{(2)}_{i}f_{i}^{h}$, and $S(g_{1},g_{2})=\sum_{i=1}^{m}b_{i}f_{i}^{h}$ is obtained by $\Sigma$-reduction in the $F_{5}$ algorithm (or its variant such as the matrix-$F_{5}$ algorithm) with the Schreyer ordering.
+Thus, since the $F_{5}$ algorithm (or its variant) automatically discards an S-polynomial whose signature is the LM of some trivial syzygy, we can avoid unnecessary S-polynomials.
+See [18] for the $F_{5}$ algorithm and its variant, and also for the syzygy criterion.
 In addition to the above facts, as mentioned (somehow implicitly) in [1, Section 3.5] and [4], when we compute a Gröbner basis of $\langle F^{h}\rangle$ for the degree less than $D$ by the $F_{5}$ algorithm with respect to $\prec^{h}$, for each computed non-zero polynomial $g$ from an S-polynomial, say $S(g_{1},g_{2})$, of degree less than $D$, its signature does not come from any trivial syzygy and so the reductions of $S(g_{1},g_{2})$ are done only at its top part.
@@
 (We also discuss similar correspondences among the Gröbner basis computation of $\langle F^{h}\rangle$ and that of $\langle F\rangle$ in the next subsection.) Especially, the following lemma holds.
-Lemma 4.1.4 ([30, Lemma 2]).
+Lemma 4.1.4 ([31, Lemma 2]).
 With notation as above, assume that $\bm{F}=(f_{1},\ldots,f_{m})\in R^{m}$ is affine cryptographic semi-regular. For each degree $d<D$, we have
 ${\rm LM}(G_{\rm hom})_{d}={\rm LM}(G_{\rm top})_{d}.$ (4.1.1)
-We also note that the argument and the proof of Lemma 4.1.4 can be considered as a corrected version of [38, Theorem 4].
+We also note that the argument and the proof of Lemma 4.1.4 can be considered as a corrected version of [39, Theorem 4].
 Next we consider $(G_{\rm hom})_{D}$. The following lemma holds, not assuming that $\bm{F}$ is affine cryptographic semi-regular:
-Lemma 4.1.5 ([30, Lemma 3]).
+Lemma 4.1.5 ([31, Lemma 3]).
 Assume that $D=d_{\rm reg}(\langle F^{\rm top}\rangle)<\infty$ (the assumption that $\bm{F}$ is affine cryptographic semi-regular is not necessary).
@@
 4.2 Correspondence between the computations of $G_{\rm hom}$ and $G$
-In this subsection, we show that, at early stages, there is a strong correspondence between the computation of $G_{\rm hom}$ and that of $G$, from which we shall extend the upper bound on solving degree given in [40, Theorem 2.1] to our case.
+In this subsection, we show that, at early stages, there is a strong correspondence between the computation of $G_{\rm hom}$ and that of $G$, from which we shall extend the upper bound on solving degree given in [41, Theorem 2.1] to our case.
 Remark 4.2.1.
 In [41], polynomial ideals over $R={\mathbb{F}}_{q}[x_{1},\ldots,x_{n}]$ are considered.
-Under the condition where the generating set $F$ contains the field equations $x_{i}^{q}-x_{i}$ for $1\leq i\leq n$, recall from Theorem 2.2.5 ([41, Theorem 6.5 & Corollary 3.67]) that the solving degree $\mathrm{sd}_{\prec}^{\mathcal{A}}(F)$ in the strict sense (see the definition (I) of Subsection 2.2 for the definition) with respect to a Buchberger-like algorithm $\mathcal{A}$ for $\langle F\rangle$ is upper-bounded by $22D-$ 2, where $D=d_{\rm deg}(\langle F^{\rm top}\rangle)$.
-In the proofs of [41, Theorem 6.5 & Corollary 3.67], the property $\langle F^{\rm top}\rangle_{D}=R_{D}$ was essentially used for obtaining the upper-bound.
+Under the condition where the generating set $F$ contains the field equations $x_{i}^{q}-x_{i}$ for $1\leq i\leq n$, recall from Theorem 2.2.5 ([42, Theorem 6.5 & Corollary 3.67]) that the solving degree $\mathrm{sd}_{\prec}^{\mathcal{A}}(F)$ in the strict sense (see the definition (I) of Subsection 2.2 for the definition) with respect to a Buchberger-like algorithm $\mathcal{A}$ for $\langle F\rangle$ is upper-bounded by $22D-$ 2, where $D=d_{\rm deg}(\langle F^{\rm top}\rangle)$.
+In the proofs of [42, Theorem 6.5 & Corollary 3.67], the property $\langle F^{\rm top}\rangle_{D}=R_{D}$ was essentially used for obtaining the upper-bound.
 As the property also holds in our case, we may apply their arguments.
-Also in [4, Section 3.2], the case where $F^{h}$ is cryptographic semi-regular is considered.
+Also in [5, Section 3.2], the case where $F^{h}$ is cryptographic semi-regular is considered.
 The results on the solving degree and the maximal degree of the Gröbner basis are heavily related to our results in this subsection.
@@
 Here we show an upper-bound on the highest step degree appeared in the computation of $G$ with respect to the DRL ordering by a Buchberger-like algorithm $\mathcal{A}$ based on S-polynomials with the normal strategy and the sugar strategy.
-We note that, in [30, Lemma 4.2.4], we restart the computation of the Gröbner basis of $F$ from $H=\{g|_{y=1}:g\in(G_{\rm hom})_{\leq D}\}$.
+We note that, in [31, Lemma 4.2.4], we restart the computation of the Gröbner basis of $F$ from $H=\{g|_{y=1}:g\in(G_{\rm hom})_{\leq D}\}$.
 However, here we do not need $(G_{\rm hom})_{\leq D}$.
-We refer to [8, Remark 15] for another proof of ${\rm max.GB.}\deg_{\prec}(F)\leq D$.
-Lemma 4.2.4 (cf. [30, Lemma 4]).
+We refer to [9, Remark 15] for another proof of ${\rm max.GB.}\deg_{\prec}(F)\leq D$.
+Lemma 4.2.4 (cf. [31, Lemma 4]).
 Assume that $D\geq{\rm max}\{\deg(f):f\in F\}$, and that $\prec$ is a DRL ordering on the set of monomials in $R$.
@@
 Moreover, there exists a Buchberger-like algorithm $\mathcal{A}$ with normal strategy such that
-${\rm sd}_{\prec}^{\mathcal{A}}(F)\leq{\color[rgb]{0,0,0}2D-1},$
+${\rm sd}_{\prec}^{\mathcal{A}}(F)\leq 2D-1,$
 and
-${\rm sd}_{\prec}^{\mathcal{A}}(F)\leq{\color[rgb]{0,0,0}2D-2}.$
+${\rm sd}_{\prec}^{\mathcal{A}}(F)\leq 2D-2$
 in the strict sense (see (I) in Subsection 2.2 for details on the definition of these solving degrees).
@@
 Remark 4.2.5.
-We refer to [8, Remark 15] for another proof of ${\rm max.GB.}\deg_{\prec}({F})\leq D$.
+We refer to [9, Remark 15] for another proof of ${\rm max.GB.}\deg_{\prec}({F})\leq D$.
 We also note that, if $D=d_{\rm reg}({F}^{\rm top})<\infty$, Lemma 4.2.3 and Lemma 4.2.4 hold without the assumption that ${F^{\rm top}}$ is cryptographic semi-regular.
-Remark 4.2.6 (cf. [30, Section 5.2]).
+Remark 4.2.6 (cf. [31, Section 5.2]).
 As to the computation of $G_{\rm hom}$, we have a result similar to Lemma 4.2.4.
@@
 By using the $F_{5}$ algorithm or its efficient variant, under an assumption that every unnecessary S-polynomial can be avoided, the number of computed S-polynomials during the computation of $G_{\rm hom}$ coincides with the number #$\#{\mathcal{U}}$ and is upper-bounded by $\binom{n+D}{n}$.
-We review a simple example shown in [30, Example 1] and examine the correspondences discussed in this and the previous subsections.
+We review a simple example shown in [31, Example 1] and examine the correspondences discussed in this and the previous subsections.
 Example 4.2.7.
@@
 In our experiments, given $n$, $m$, $(d_{1},\ldots,d_{m})$, and an odd prime $q$ with $n<m$ and $d_{i}\geq 2$ (for the case where $n=m$, see Remark 4.3.3 below), we generate inhomogeneous polynomials $f_{1},\ldots,f_{m}$ in $R=\mathbb{F}_{q}[x_{1},\ldots,x_{n}]$ of degrees $d_{1},\ldots,d_{m}$ whose constant terms are all zero (as in Example 4.2.7), where coefficients are chosen uniformly at random.
-Note that, for any sequence $\bm{F}=(f_{1},\ldots,f_{m})$ generated in this way, the homogenization $F^{h}$ with $F=\{f_{1},\ldots,f_{m}\}$ always has a non-trivial root $(0,\ldots,0,1)$: Any sequence of $m$ polynomials (in $R$) that has at least one root is transformed by a linear coordinate change into a polynomial sequence $(f_{1},\ldots,f_{m})\in R^{m}$ such that each $f_{i}$ has no constant term.
+Note that, for any sequence $\bm{F}=(f_{1},\ldots,f_{m})$ generated in this way, the homogenization $F^{h}$ with $F=\{f_{1},\ldots,f_{m}\}$ always has a projective zero $(0:\cdots:0:1)$: Any sequence of $m$ polynomials (in $R$) that has at least one affine zero is transformed by a linear coordinate change into a polynomial sequence $(f_{1},\ldots,f_{m})\in R^{m}$ such that each $f_{i}$ has no constant term.
 Moreover, $\bm{F}^{\rm top}=(f_{1}^{\rm top},\ldots,f_{m}^{\rm top})$ is expected to be cryptographic semi-regular (in fact, semi-regular) with high probability.
@@
 As an experimental result, we heuristically find an upper-bound
-${\deg\left(\left[\frac{\prod_{i=1}^{m}(1-z^{d_{i}})}{(1-z)^{n+1}}\right]\right)}+1$ (4.3.1)
+$D_{\rm new}:={\deg\left(\left[\frac{\prod_{i=1}^{m}(1-z^{d_{i}})}{(1-z)^{n+1}}\right]\right)}+1$ (4.3.1)
 on the maximal Gröbner basis degree $\mathrm{max.GB.deg}_{\prec^{h}}(F^{h})$, which is sharper than our upper-bound provided in Theorem 2 (1).
-This sharper bound comes from a property that $\bm{F}^{h}$ is generalized cryptographic semi-regular (i.e., $D^{\prime}$-regular), where we set $D^{\prime}:=\widetilde{d}_{\rm reg}(F^{h})$: We confirmed in our experiments that $\bm{F}^{h}$ satisfies this property in most cases.
-Here we use the following lemma (the proof is straightfoward):
+This sharper bound comes from a property that $\bm{F}^{h}$ is generalized cryptographic semi-regular (i.e., $D^{\prime}$-regular), where we set $D^{\prime}:=\widetilde{d}_{\rm reg}(\langle F^{h}\rangle_{R^{\prime}})$: We confirmed in our experiments that $\bm{F}^{h}$ satisfies this property in most cases.
+Here we use the following lemma (the proof is straightforward):
 Lemma 4.3.1.
-Let $\bm{F}=(f_{1},\ldots,f_{m})\in(R\smallsetminus K)^{m}$ be a sequence of not necessarily homogeneous polynomial. Assume that $n<m$. Then we have the following:
+Let $\bm{F}=(f_{1},\ldots,f_{m})\in(R\smallsetminus K)^{m}$ be a sequence of not necessarily homogeneous polynomials. Assume that $n<m$. Then we have the following:
   1. 1.
-    If ${F}^{h}$ has at least one solution and if $\bm{F}^{h}$ is $D^{\prime}$-regular (and thus $D^{\prime}<\infty$), then $D^{\prime}$ is upper-bounded by (4.3.1).
+    If ${F}^{h}$ has at least one projective zero and if $\bm{F}^{h}$ is $D^{\prime}$-regular (and thus $D^{\prime}<\infty$), then $D^{\prime}$ is upper-bounded by $D_{\rm new}$ given in (4.3.1).
   2. 2.
-    If $\bm{F}^{\rm top}$ is cryptographic semi-regular, then $D:=d_{\rm reg}(F^{\rm top})$ is also upper-bounded by (4.3.1) (this also holds even if $n=m$).
+    If $\bm{F}^{\rm top}$ is cryptographic semi-regular, then $D:=d_{\rm reg}(\langle F^{\rm top}\rangle_{R})$ is also upper-bounded by $D_{\rm new}$ given in (4.3.1) (this also holds even if $n=m$).
 Supposing both the assumptions in this lemma, we have
-$\mathrm{max.GB.deg}_{\prec^{h}}(F^{h})\leq\max\{D,D^{\prime}\}\leq{\deg\left(\left[\frac{\prod_{i=1}^{m}(1-z^{d_{i}})}{(1-z)^{n+1}}\right]\right)}+1$
-by Proposition 2.3.3.
-We compare exact values for several bounds, for $n=9$ and $10$ with some conditions on $m$ and $(d_{1},\ldots,d_{m})$ in Tables 14, where we set $D:=d_{\rm reg}(F^{\rm top})$ for the case where $\bm{F}^{\rm top}$ is cryptographic semi-regular.
-Table 1: Exact values for several upper-bounds on $\mathrm{max.GB.deg}_{\prec^{h}}(F^{h})$ in the case where $n=9$, $n+1\leq m\leq 2n$, and $d_{1}=\cdots=\cdots=d_{m}=2$.
-$m$ $10$ $11$ $12$ $13$ $14$ $15$ $16$ $17$ $18$
+$\mathrm{max.GB.deg}_{\prec^{h}}(F^{h})\leq\max\{D,D^{\prime}\}\leq D_{\rm new}$ (4.3.2)
+by Proposition 2.3.6 (this holds even for an arbitrary graded monomial ordering on $R$).
+Therefore, the complexity of computing $G_{\rm hom}$ (and $G$) with respect to the number of arithmetic operations on $K$ can be estimated as
+$O\left(m\binom{n+D_{\rm new}}{D_{\rm new}}^{\omega}\right)$ (4.3.3)
+by Corollary A.4.2 and Remark A.4.4 below.
+We also compare exact values for several bounds on $\mathrm{max.GB.deg}_{\prec^{h}}(F^{h})$, for $n=9$ and $10$ with some conditions on $m$ and $(d_{1},\ldots,d_{m})$ in Tables 1 and 2, where we set $D:=d_{\rm reg}(\langle F^{\rm top}\rangle_{R})$ for the case where $\bm{F}^{\rm top}$ is cryptographic semi-regular, namely
+$D={\deg\left(\left[\frac{\prod_{i=1}^{m}(1-z^{d_{i}})}{(1-z)^{n}}\right]\right)}+1,$
+which tends to be close to $D_{\rm new}$ for $m\gg n$.
+Table 1: Exact values for several upper-bounds on $\mathrm{max.GB.deg}_{\prec^{h}}(F^{h})$ (which is equal to the solving degree of $F^{h}$ in this case) in the case where $n\in\{9,10\}$, $n+1\leq m\leq 2n$, and $d_{1}=\cdots=\cdots=d_{m}=2$.
+The first table is for the case $n=9$, and the second one is for the case $n=10$.
+The number $m$ ($>n=9$) of polynomials $10$ $11$ $12$ $13$ $14$ $15$ $16$ $17$ $18$
 Lazard’s bound (Theorem 2.2.1) $11$ $11$ $11$ $11$ $11$ $11$ $11$ $11$ $11$
 Our bound in Theorem 2 (1) $11$ $11$ $11$ $11$ $11$ $11$ $11$ $11$ $11$
-Heuristic bound (4.3.1) (Conjecture 4.3.4) $11$ 66{6}6 66{6}6 55{5}5 55{5}5 44{4}4 44{4}4 44{4}4 44{4}4
+$D_{\rm new}$ given in (4.3.1) (Conjecture 4.3.4) $11$ 66{6}6 66{6}6 55{5}5 55{5}5 44{4}4 44{4}4 44{4}4 44{4}4
 $D$ for semi-regular ${F}^{\rm top}$ $6$ $5$ $5$ $4$ $4$ $4$ $4$ $4$ $4$
 $12D-$ 1 (Theorem 3 (4)) $1${11}11 $9$ $9$ $7$ $7$ $7$ $7$ $7$ $7$
-Table 2: Exact values for several upper-bounds on $\mathrm{max.GB.deg}_{\prec^{h}}(F^{h})$ in the case where $n=10$, $n+1\leq m\leq 2n$, and $d_{1}=\cdots=d_{m}=2$.
-$m$ $11$ $12$ $13$ $14$ $15$ $16$ $17$ $18$ $19$ $20$
+The number $m$ ($>n=10$) of polynomials $11$ $12$ $13$ $14$ $15$ $16$ $17$ $18$ $19$ $20$
 Lazard’s bound (Theorem 2.2.1) $12$ $12$ $12$ $12$ $12$ $12$ $12$ $12$ $12$ $12$
 Our bound in Theorem 2 (1) $12$ $12$ $12$ $12$ $12$ $12$ $12$ $12$ $12$ $12$
-Heuristic bound (4.3.1) (Conjecture 4.3.4) $12$ 77{7}7 66{6}6 55{5}5 55{5}5 55{5}5 55{5}5 44{4}4 44{4}4 44{4}4
+$D_{\rm new}$ given in (4.3.1) (Conjecture 4.3.4) $12$ 77{7}7 66{6}6 55{5}5 55{5}5 55{5}5 55{5}5 44{4}4 44{4}4 44{4}4
 $D$ for semi-regular $\bm{F}^{\rm top}$ $6$ $6$ $5$ $5$ $4$ $4$ $4$ $4$ $4$ $4$
 $12D-$ 1 (Theorem 3 (4)) $1${11}11 $11$ $9$ $9$ $7$ $7$ $7$ $7$ $7$ $7$
-Table 3: Exact values for several upper-bounds on $\mathrm{max.GB.deg}_{\prec^{h}}(F^{h})$ in the case where $n=9$, $n+1\leq m\leq 2n$, $d_{1}=\cdots=d_{n}=3$, and $d_{n+1}=\cdots=d_{m}=2$.
-$m$ $10$ $11$ $12$ $13$ $14$ $15$ $16$ $17$ $18$
+Table 2: Exact values for several upper-bounds on $\mathrm{max.GB.deg}_{\prec^{h}}(F^{h})$ (which is equal to the solving degree of $F^{h}$ in this case) in the case where $n\in\{9,10\}$, $n+1\leq m\leq 2n$, $d_{1}=\cdots=d_{n}=3$, and $d_{n+1}=\cdots=d_{m}=2$.
+The first table is for the case $n=9$, and the second one is for the case $n=10$.
+The number $m$ ($>n=9$) of polynomials $10$ $11$ $12$ $13$ $14$ $15$ $16$ $17$ $18$
 Lazard’s bound (Theorem 2.2.1) $20$ $20$ $20$ $20$ $20$ $20$ $20$ $20$ $20$
 Our bound in Theorem 2 (1) $20$ $19$ $18$ $17$ $16$ $15$ $14$ $13$ $12$
-Heuristic bound (4.3.1) (Conjecture 4.3.4) $20$ $1${11}11 99{9}9 88{8}8 77{7}7 77{7}7 66{6}6 66{6}6 55{5}5
+$D_{\rm new}$ given in (4.3.1) (Conjecture 4.3.4) $20$ $1${11}11 99{9}9 88{8}8 77{7}7 77{7}7 66{6}6 66{6}6 55{5}5
 $D$ for semi-regular ${F}^{\rm top}$ $10$ $9$ $8$ $7$ $6$ $6$ $6$ $5$ $5$
 $12D-$ 1 (Theorem 3 (4)) 1919{19}19 $17$ $15$ $13$ $11$ $11$ $9$ $9$ $9$
-Table 4: Exact values for several upper-bounds on $\mathrm{max.GB.deg}_{\prec^{h}}(F^{h})$ in the case where $n=10$, $n+1\leq m\leq 2n$, $d_{1}=\cdots=d_{n}=3$, and $d_{n+1}=\cdots=d_{m}=2$.
-$m$ $11$ $12$ $13$ $14$ $15$ $16$ $17$ $18$ $19$ $20$
+The number $m$ ($>n=10$) of polynomials $11$ $12$ $13$ $14$ $15$ $16$ $17$ $18$ $19$ $20$
 Lazard’s bound (Theorem 2.2.1) $22$ $22$ $22$ $22$ $22$ $22$ $22$ $22$ $22$ $22$
 Our bound in Theorem 2 (1) $22$ $21$ $20$ $19$ $18$ $17$ $16$ $15$ $14$ $13$
-Heuristic bound (4.3.1) (Conjecture 4.3.4) $22$ 1212{12}12 1010{10}10 99{9}9 88{8}8 77{7}7 77{7}7 66{6}6 66{6}6 66{6}6
+$D_{\rm new}$ given in (4.3.1) (Conjecture 4.3.4) $22$ 1212{12}12 1010{10}10 99{9}9 88{8}8 77{7}7 77{7}7 66{6}6 66{6}6 66{6}6
 $D$ for semi-regular $\bm{F}^{\rm top}$ $11$ $10$ $9$ $8$ $7$ $6$ $6$ $6$ $5$ $5$
@@
 As a more particular case, if $d_{i}=2$ for all $i$, then it is equal to $n+2$.
-On the other hand, recall from [4, Theorem 4.1] that $D=\lfloor(n+1)/2\rfloor+1$, and thus $12D-$ 1 is equal to $n+2$ if $n$ is odd and to $n+1$ if $n$ is even.
+On the other hand, recall from [5, Theorem 4.1] that $D=\lfloor(n+1)/2\rfloor+1$, and thus $12D-$ 1 is equal to $n+2$ if $n$ is odd and to $n+1$ if $n$ is even.
+More precisely, assuming the $D^{\prime}$-regularity of $\bm{F}^{h}$, we have
+${\rm HS}_{R^{\prime}/\langle F^{h}\rangle}(z)\equiv(1+z)^{n+1}\pmod{z^{D^{\prime}}},$
+where $(1+z)^{n+1}=\sum_{i=0}^{n+1}\binom{n+1}{i}z^{i}$.
+In the expansion of $(1+z)^{n+1}$, the coefficient of $z^{(n+1)/2}$ is (resp.
+the coefficients of $z^{n/2}$ and $z^{(n+2)/2}$ are) maximal among the non-zero coefficients for an odd (resp.
+even) $n$, see Figure 1 for some specific $n$.
+In particular, if $n$ is even, then we have $D=n/2+1$ and $\dim_{K}(R^{\prime}/\langle F^{h}\rangle)_{D-1}=\dim_{K}(R^{\prime}/\langle F^{h}\rangle)_{D}$, which means that multiplication by-$y$ map $(R^{\prime}/\langle F^{h}\rangle)_{D-1}\longrightarrow(R^{\prime}/\langle F^{h}\rangle)_{D}$ is bijective.
+Thus, in the Gröbner basis computation of $F$, there is no degree fall at degree $D$ (in fact, up to $D$, see Subsections 4.1 and 4.2).
 Remark 4.3.3.
-If $m=n$, then it follows from $\frac{\prod_{i=1}^{m}(1-z^{d_{i}})}{(1-z)^{n}}=\prod_{i=1}^{n}(1+z+\cdots+z^{d_{i}-1})$ that $D=\sum_{j=1}^{n}(d_{j}-1)+1$, and we have
-$\frac{\prod_{i=1}^{m}(1-z^{d_{i}})}{(1-z)^{n+1}}=\frac{\prod_{i=1}^{n}(1-z^{d_{i}})}{(1-z)^{n}}\cdot(1+z+z^{2}+\cdots),$
-whose degree-$d$ coefficients are equal to a constant for $d\geq D-1$.
-In this case, the $D^{\prime}$-regularity of $F^{h}$ implies that $D^{\prime}\leq D$, whence $D^{\prime}\in\{D-1,D\}$ and $\mathrm{max.GB.deg}_{\prec^{h}}(F^{h})\leq D$ by Proposition 2.3.3.
-Note that the equality $\mathrm{max.GB.deg}_{\prec^{h}}(F^{h})=D$ holds if $\langle\mathrm{LM}(\bm{F}^{h})\rangle$ is a weakly reverse lexicographic ideal.
+If $m=n$, then we can prove $D^{\prime}=D-1$ and $\mathrm{max.GB.deg}_{\prec^{h}}(F^{h})\leq D$ only assuming that $\bm{F}^{\rm top}$ is cryptographic semi-regular, and consequently $\bm{F}^{h}$ is generalized cryptographic semi-regular.
+Indeed, by [22, b) on page 121], we have the following coefficient-wise inequality:
+${\rm HS}_{R^{\prime}/\langle F^{h}\rangle}(z)\geq\frac{\prod_{i=1}^{m}(1-z^{d_{i}})}{(1-z)^{n+1}}=\frac{\prod_{i=1}^{n}(1-z^{d_{i}})}{(1-z)^{n}}\cdot(1+z+z^{2}+\cdots).$ (4.3.4)
+The degree-$d$ coefficients of the right hand side are equal to a constant for $d\geq D-1$, where $D=\sum_{j=1}^{n}(d_{j}-1)+1$ by $\frac{\prod_{i=1}^{m}(1-z^{d_{i}})}{(1-z)^{n}}=\prod_{i=1}^{n}(1+z+\cdots+z^{d_{i}-1})$.
+The $D$-regularity of $\bm{F}^{\rm top}$ implies that $\bm{F}^{h}$ is also $D$-regular, whence the degree $d$-coefficients of the both hand sides of (4.3.4) are equal to one another for each $d$ with $d\leq D-1$, so that we can easily check $D^{\prime}\geq D-1$ (see also the first assertion of Proposition 2.3.6).
+Moreover, recall from Remark 3.1.2 that ${\rm HF}_{R^{\prime}/\langle F^{h}\rangle}(D-1)\geq{\rm HF}_{R^{\prime}/\langle F^{h}\rangle}(d)$ for any $d\geq D-1$.
+Therefore, the inequality of (4.3.4) is in fact the equality, and hence $D^{\prime}=D-1$.
+Therefore, we obtain $\mathrm{max.GB.deg}_{\prec^{h}}(F^{h})\leq D$ by Proposition 2.3.6, and also $\bm{F}^{h}$ is $D^{\prime}$-regular.
+Note that the equality $\mathrm{max.GB.deg}_{\prec^{h}}(F^{h})=D$ holds if $\prec$ is a DRL ordering and if $\langle\mathrm{LM}(\langle{F}^{h}\rangle)\rangle$ is a weakly reverse lexicographic ideal, see Proposition 2.3.6.
+Anyway, in the case where $n=m$, the complexity of computing $G_{\rm hom}$ (and $G$) with respect to the number of arithmetic operations on $K$ can be estimated as
+$O\left(n\binom{n+D}{D}^{\omega}\right)$
+with $D=\sum_{j=1}^{n}(d_{j}-1)+1$, by Corollary A.4.2 and Remark A.4.4 below.
+Figure 1: The values of coefficients in $(1+z)^{n+1}$ for $(n,m)=(9,10)$ (the left figure) and $(n,m)=(10,11)$ (the right figure).
+The horizontal axis indicates the degree $i$ of $z^{i}$, and the vertical axis indicates the value of the coefficient of $z^{i}$.
+Note that $D=(n+3)/2$ for an odd $n$ and $D=(n+2)/2$ for even $n$, and thus $D-1=5$ for $n\in\{9,10\}$.
+See Remark 4.3.2 for a description.
 Based on our experiments, we here raise the following conjecture:
@@
 Let $d_{1},\ldots,d_{m}$ are integers larger than $1$, and let $f_{1},\ldots,f_{m}$ be polynomials in $R$ of degrees $d_{1},\ldots,d_{m}$ such that each $f_{i}$ has no constant term.
-Then, for given $K$, $n$, $m$, and $(d_{1},\ldots,d_{m})$, the property that $\bm{F}^{h}=(f_{1}^{h},\ldots,f_{m}^{h})$ is generalized cryptographic semi-regular (i.e., $D^{\prime}$-regular) is generic, where $D^{\prime}=\widetilde{d}_{\rm reg}(F^{h})$ is the generalized degree of regularity of $\bm{F}^{h}$ defined in Definition 2.3.1.
-Conjecture 4.3.4 can be viewed as Fröberg’s conjecture [21]: A generic sequence of homogeneous polynomials is $D$-regular, where $D$ is the degree of regularity of the sequence.
+Then, for given $K$, $n$, $m$, and $(d_{1},\ldots,d_{m})$, the property that $\bm{F}^{h}=(f_{1}^{h},\ldots,f_{m}^{h})$ is generalized cryptographic semi-regular (i.e., $D^{\prime}$-regular) is generic, where $D^{\prime}=\widetilde{d}_{\rm reg}(\langle F^{h}\rangle_{R^{\prime}})$ is the generalized degree of regularity of $\langle{F}^{h}\rangle_{R^{\prime}}$ defined in Definition 2.3.1.
+Conjecture 4.3.4 can be viewed as a variant of Fröberg’s conjecture [22]: A generic sequence of homogeneous polynomials is $D$-regular, where $D$ is the degree of regularity of the ideal generated by the sequence.
+Remark 4.3.5.
+In Conjecture 4.3.4, the reason why we take all the constant terms of $f_{i}$’s to be zero (equivalently the ideal $\langle F\rangle_{R}$ vanishes at $(0,\ldots,0)$) is the following: For any inhomogeneous polynomial system having at least one affine zero $(a_{1},\ldots,a_{n})$, we can convert a system of inhomogeneous polynomials of the form $f_{1},\ldots,f_{m}$ as in the conjecture with an affine linear transformation sending $(a_{1},\ldots,a_{n})$ to $(0,\ldots,0)$.
 As a consequence of Conjecture 4.3.4, if $F$ is generated in the way described in the beginning of this subsection, we may expect the following properties:
   1. 1.
-    As we described as above, the solving degree $\mathrm{sd}^{\rm mac}_{\prec^{h}}(F^{h})$ (which is equal to the solving degree $\mathrm{sd}^{\rm mac}_{\prec}(F)$ if $\prec$ is a DRL order, see Subsection 2.2) is also upper-bounded by $D^{\prime}$, which can be quite smaller than $12D-$ 1.
+    As we described as above, the solving degree $\mathrm{sd}^{\rm mac}_{\prec^{h}}(F^{h})$ (which is equal to the solving degree $\mathrm{sd}^{\rm mac}_{\prec}(F)$ if $\prec$ is a DRL order, see Subsection 2.2) is also upper-bounded by $D_{\rm new}$ given in (4.3.1), which can be quite smaller than $12D-$ 1.
   2. 2.
@@
     ${\rm HS}_{A^{\prime}}(z)=\left(\frac{\prod_{j=1}^{m}(1-z^{d_{j}})}{(1-z)^{n}}\bmod{z^{D^{\prime}}}\right)+\sum_{d=D^{\prime}}^{\infty}N_{F^{h}}z^{d},$ (4.3.5)
-    where $N_{F^{h}}$ is the number of projective roots of $F^{h}$ counted with multiplicity.
+    where $N_{F^{h}}$ is the number of projective zeros of $\langle F^{h}\rangle$ counted with multiplicity.
 This implies that the Hilbert driven algorithm can be effectively applied to the Gröbner basis computation of $F^{h}$, from which a Gröbner basis of $F$ is easily obtained.
@@
 Here, we also raise interesting questions:
-Question 4.3.5.
+Question 4.3.6.
   1. 1.
@@
     • (B)
-      $\langle\mathrm{LM}(\bm{F}^{h})\rangle$ is a weakly reverse lexicographic ideal. (Cf. Moreno-Socías conjecture [35].)
+      $\langle\mathrm{LM}(\langle{F}^{h}\rangle)\rangle$ is a weakly reverse lexicographic ideal. (Cf. Moreno-Socías conjecture [36].)
   3. 3.
@@
 Final remark for security analysis in cryptography
-The assumption that $\bm{F}^{h}$ is generalized cryptographic semi-regular (i.e., $D^{\prime}$-regular) could be useful to estimate the security of multivariate cryptosystems and signature schemes (or algebraic attacks based on Gröbner basis computation), see e.g., [22, Subsection 2.3], where the authors of [22] assume that the bound (4.3.1) gives a degree bound of the XL algorithm [11].
-In fact, the bound (4.3.1) has been sometimes used in the cryptographic community without assuming the $D^{\prime}$-regularity of $\bm{F}^{h}$, see [22, Subsection 2.3] (in particular [22, Remark 1]) for details.
-Supposing the $D^{\prime}$-regularity of the homogenization of a target polynomial system, one could give a mathematically rigid explanation about the complexity of computing the Gröbner basis of the system.
+The assumption that $\bm{F}^{h}$ is generalized cryptographic semi-regular (i.e., $D^{\prime}$-regular) could be useful to estimate the security of multivariate cryptosystems (or algebraic attacks based on Gröbner basis computation), see e.g., [23, Subsection 2.3], where the authors of [23] assume that the bound (4.3.1) gives a degree bound of the XL algorithm [12].
+In fact, the bound (4.3.1) has been sometimes used in the cryptographic community without assuming the $D^{\prime}$-regularity of $\bm{F}^{h}$, see [23, Subsection 2.3] (in particular [23, Remark 1]) for details.
+Supposing the $D^{\prime}$-regularity of the homogenization of a target polynomial system, one could give a mathematically rigid explanation about the complexity of computing the Gröbner basis of the system, and one can apply our estimation (4.3.3) for the complexity.
 Acknowledgement
@@
 This work was supported by JSPS Grant-in-Aid for Young Scientists 20K14301 and 23K12949, JSPS Grant-in-Aid for Scientific Research (C) 21K03377, and JST CREST Grant Number JPMJCR2113.
-Appendix A Koszul complex and homogenization
+Appendix A Supplemental definitions and results
 A.1 Koszul complex
@@
 In the following, we denote by $H_{i}(K_{\bullet})_{d}$ the degree-$d$ homogeneous part of $H_{i}(K_{\bullet})$.
-Note that $\mathrm{Ker}(\varphi_{1})=\mathrm{syz}(f_{1},\ldots,f_{m})$ (the right hand side is the module of syzygies), and that $\mathrm{Im}(\varphi_{2})\subset K_{1}=\bigoplus_{j=1}^{m}R(-d_{j})\mathbf{e}_{j}$ is generated by
+Here, the module of syzygies $\mathrm{syz}(f_{1},\ldots,f_{m})$ is defined as $\mathrm{Ker}(\varphi_{1})$, and its element is called a syzygy for $(f_{1},\ldots,f_{m})$.
+We also note that $\mathrm{Im}(\varphi_{2})\subset K_{1}=\bigoplus_{j=1}^{m}R(-d_{j})\mathbf{e}_{j}$ is generated by
 $\{\mathbf{t}_{i,j}:=f_{i}\mathbf{e}_{j}-f_{j}\mathbf{e}_{i}:1\leq i<j\leq m\}.$
@@
 A.2 Homogenization of polynomials and monomial orders
-We here recall the notion of homogenization; see [29, Chapter 4] for details.
+We here recall the notion of homogenization; see [30, Chapter 4] for details.
 Let $R=K[x_{1},\ldots,x_{n}]$ be the polynomial ring of $n$ variables over a field $K$, and $\mathcal{T}$ the set of all monomials in $x_{1},\ldots,x_{n}$.
@@
   1. (1)
-    For a non-homogeneous and non-zero polynomial $f=\sum_{{t}\in\mathcal{T}}c_{t}{t}$ in $R$ with $c_{t}\in K$, its homogenization $f^{h}$ is defined, by introducing an extra variable $y$, as
-    $f^{h}=\sum_{{t}\in\mathcal{T}}c_{t}{t}{y^{\deg(f)-{\color[rgb]{0,0,0}\deg(t)}}}\in R^{\prime}=R[y].$
+    For an inhomogeneous and non-zero polynomial $f=\sum_{{t}\in\mathcal{T}}c_{t}{t}$ in $R$ with $c_{t}\in K$, its homogenization $f^{h}$ is defined, by introducing an extra variable $y$, as
+    $f^{h}=\sum_{{t}\in\mathcal{T}}c_{t}{t}{y^{\deg(f)-{\deg(t)}}}\in R^{\prime}=R[y].$
     Thus $f^{h}$ is a homogeneous polynomial in $R^{\prime}$ with total degree $d=\deg(f)$.
@@
   5. (5)
-    For a monomial ordering $\prec$ on the set of monomials $\mathcal{T}$ in $X$, its homogenization $\prec_{h}$ on the set of monomials $\mathcal{T}^{h}$ in $x_{1},\ldots,x_{n},y$ is defined as follows: For two monomials $X^{\alpha}y^{a}$ and $X^{\beta}y^{b}$ in $\mathcal{T}^{h}$, we say $X^{\alpha}y^{a}\prec_{h}X^{\beta}y^{b}$ if and only if one of the following holds:
+    For a monomial ordering $\prec$ on the set of monomials $\mathcal{T}$ in $X$, its homogenization $\prec^{h}$ on the set of monomials $\mathcal{T}^{h}$ in $x_{1},\ldots,x_{n},y$ is defined as follows: For two monomials $X^{\alpha}y^{a}$ and $X^{\beta}y^{b}$ in $\mathcal{T}^{h}$, we say $X^{\alpha}y^{a}\prec^{h}X^{\beta}y^{b}$ if and only if one of the following holds:
     • (i)
@@
     where $\alpha=(\alpha_{1},\ldots,\alpha_{n})\in\mathbb{Z}_{\geq 0}^{n}$ and $|\alpha|=\alpha_{1}+\cdots+\alpha_{n}$, and where $X^{\alpha}$ denotes $x_{1}^{\alpha_{1}}\cdots x_{n}^{\alpha_{n}}$.
-Here, for a monomial $X^{\alpha}{\color[rgb]{0,0,0}y^{a}}$, we call $X^{\alpha}$ and ${\color[rgb]{0,0,0}y^{a}}$ the $X$-part and the $y$-part, respectively.
-If a monomial ordering $\prec$ is graded, that is, it first compares the total degrees, the restriction $\prec_{h}|_{\mathcal{T}}$ of $\prec_{h}$ on $\mathcal{T}$ coincides with $\prec$.
+Here, for a monomial $X^{\alpha}{y^{a}}$, we call $X^{\alpha}$ and ${y^{a}}$ the $X$-part and the $y$-part, respectively.
+If a monomial ordering $\prec$ is graded, that is, it first compares the total degrees, the restriction $\prec^{h}|_{\mathcal{T}}$ of $\prec^{h}$ on $\mathcal{T}$ coincides with $\prec$.
 It is well-known that, for a Gröbner basis $H$ of $\langle F^{h}\rangle$ with respect to $\prec^{h}$, its dehomogenization $H^{\rm deh}=\{h^{\rm deh}:h\in H\}$ is also a Gröbner basis of $\langle F\rangle$ with respect to $\prec$ if $\prec$ is graded.
-Moreover, we have $\langle F\rangle^{h}=(\langle F^{h}\rangle:\langle y\rangle^{\infty})=(\langle F^{h}\rangle:\langle y^{k}\rangle)$ for some integer $k$, where $(\langle F^{h}\rangle:\langle y^{k}\rangle)$ is the ideal quotient of $\langle F^{h}\rangle$ by $\langle y^{k}\rangle$, namely $\{f\in R^{\prime}:f\langle y^{k}\rangle\subset\langle F^{h}\rangle\}$ see [29, Corollary 4.3.8].
+Moreover, we have $\langle F\rangle^{h}=(\langle F^{h}\rangle:\langle y\rangle^{\infty})=(\langle F^{h}\rangle:\langle y^{k}\rangle)$ for some integer $k$, where $(\langle F^{h}\rangle:\langle y^{k}\rangle)$ is the ideal quotient of $\langle F^{h}\rangle$ by $\langle y^{k}\rangle$, namely $\{f\in R^{\prime}:f\langle y^{k}\rangle\subset\langle F^{h}\rangle\}$ see [30, Corollary 4.3.8].
+A.3 A generalization of Lemma 2.3.5
+We generalize Lemma 2.3.5.
+Let $R^{\prime}=K[x_{1},\ldots,x_{n},x_{n+1}]$ be the polynomial ring of $n+1$ variables over a field $K$.
+Let $I$ be a proper homogeneous ideal of $R^{\prime}$ that is zero-dimensional (see Terminology in Section 1 for the meaning of zero-dimensional), i.e., the number of its projective zeros over $\overline{K}$ is finite at most.
+Then, if necessary replacing $K$ by its finite extension with enough number of elements, there exist a non-negative integer $d$ and a linear form $\ell\in R^{\prime}$ with $\ell\notin I$ such that the $K$-linear map $(R^{\prime}/I)_{d-1}\to(R^{\prime}/I)_{d}$ defined by the multiplication by $\ell$ is surjective, see [33, Theorem 3.2] (see also [11, Theorem 3.3.4]).
+In fact, from the proof of [11, Theorem 3.3.4], it can be shown that no field extension is necessary, if the cardinality of $K$ is greater than the number of projective zeros of $I$.
+We claim that $R^{\prime}/\langle I,\ell\rangle$ is Artinian.
+Indeed, letting $H=\{h_{1},\ldots,h_{m}\}$ be a set of homogeneous polynomials in $R^{\prime}$ with $I=\langle H\rangle$ and considering a mapping cone of the Koszul complexes $K_{\bullet}$ and $K_{\bullet}^{\prime}$ on $(h_{1},\ldots,h_{m})$ and $(h_{1},\ldots,h_{m},\ell)$, we obtain an exact sequence
+[…]
+so that $H_{0}(K_{\bullet}^{\prime})_{d}=(R^{\prime}/\langle I,\ell\rangle)_{d}=0$.
+For a linear transformation $\sigma$ of variables $x_{1},\ldots,x_{n},x_{n+1}$ over $K$ represented by a $(n+1)\times(n+1)$ matrix $P$ over $K$, we denote by $h^{\sigma}:=h(\bm{x}\cdot P)$ the image of $h\in R^{\prime}$ by $\sigma$, where $\bm{x}:=(x_{1},\ldots,x_{n},x_{n+1})$.
+Lemma A.3.1.
+With notation as above, choose an integer $i\in\{1,\ldots,n+1\}$ and fix it.
+Then, there exists a linear transformation $\sigma$ of variables $x_{1},\ldots,x_{n},x_{n+1}$ over $K$ sending $\ell$ to $x_{i}$ such that, if $\widetilde{d}_{\rm reg}(I)\geq d_{\rm reg}(\langle I,\ell\rangle_{R^{\prime}})$, we have
+$\mathrm{max.GB.deg}_{\prec^{\prime}}(H^{\sigma})\leq\widetilde{d}_{\rm reg}(I)$
+with $H^{\sigma}:=\{h^{\sigma}:h\in H\}$ for any graded monomial ordering $\prec^{\prime}$ on $R^{\prime}$ satisfying $t_{1}\prec^{\prime}t_{2}$ for any two monomials $t_{1}$ and $t_{2}$ in $R^{\prime}$ of the same total degree with $\deg_{x_{i}}(t_{1})>\deg_{x_{i}}(t_{2})$.
+Proof.
+We first note that $\widetilde{d}_{\rm reg}(I)=\widetilde{d}_{\rm reg}(\langle H^{\tau}\rangle_{R^{\prime}})$ and ${d}_{\rm reg}(\langle I,\ell\rangle_{R^{\prime}})={d}_{\rm reg}(\langle H^{\tau},\ell^{\tau}\rangle_{R^{\prime}})$ for any invertible linear transformation $\tau$ of variables.
+By permuting variables, it suffices to consider the case where $i=n+1$, and we may assume that the $x_{n+1}$-coefficient of $\ell$ is not zero, say $\ell=a_{1}x_{1}+\cdots+a_{n}x_{n}+a_{n+1}x_{n+1}$ for $a_{i}\in K$ with $a_{n+1}\neq 0$.
+We may also assume that $a_{n+1}=1$, without loss of generality.
+Then, let $\sigma$ be a linear transformation of variables $x_{1},\ldots,x_{n},x_{n+1}$ represented by the matrix
+$P:=\begin{pmatrix}1&&&-a_{1}\\ &\ddots&&\vdots\\ &&1&-a_{n}\\ &&&1\\ \end{pmatrix},$
+where blank entries are 00’s. Namely, for $h\in R^{\prime}$, we define
+$h^{\sigma}:=h(\bm{x}\cdot P)=h(x_{1},\ldots,x_{n},x_{n+1}-a_{1}x_{1}-\cdots-a_{n}x_{n})$
+with $\bm{x}:=(x_{1},\ldots,x_{n},x_{n+1})$.
+This $\sigma$ sends $\ell$ to $x_{n+1}$.
+Since $P$ is invertible, we have the equalities $D^{\prime}:=\widetilde{d}_{\rm reg}(I)=\widetilde{d}_{\rm reg}(\langle H^{\sigma}\rangle_{R^{\prime}})$ and $D:=d_{\rm reg}(\langle H,\ell\rangle_{R^{\prime}})=d_{\rm reg}(\langle H^{\sigma},x_{n+1}\rangle_{R^{\prime}})$.
+Then, similarly to the proof of Lemma 2.3.5, we obtain $\mathrm{max.GB.deg}_{\prec^{\prime}}(H^{\sigma})\leq D^{\prime}$, as desired.
+∎
+By a discussion similar to the proof of Proposition 2.3.6, we also obtain the following proposition:
+Proposition A.3.2.
+In Lemma A.3.1, if the sequence $(h_{1},\ldots,h_{m},\ell)$ is cryptographic semi-regular, then
+${\rm max.GB.deg}_{\prec^{\prime}}(H^{\sigma})\leq\mathrm{max}\{d_{\rm reg}(\langle I,\ell\rangle_{R^{\prime}}),\widetilde{d}_{\rm reg}(I)\}$
+for any graded monomial ordering $\prec^{\prime}$ on $R^{\prime}$ satisfying $t_{1}\prec^{\prime}t_{2}$ for any two monomials $t_{1}$ and $t_{2}$ in $R^{\prime}$ of the same total degree with $\deg_{x_{i}}(t_{1})>\deg_{x_{i}}(t_{2})$.
+A further discussion on Lemma A.3.1 and Proposition A.3.2 will be provided in our separated paper.
+A.4 Complexity of GB computation in homogeneous case
+Let $F=\{f_{1},\ldots,f_{m}\}\subset R$ be a set of homogeneous polynomials, and let $\prec$ be a fixed graded monomial ordering on $R$.
+In this case, we have $\mathrm{sd}_{\prec}^{\rm mac}(F)=\mathrm{max.GB.deg}_{\prec}(F)$.
+Therefore, once we know the value (or an upper bound) $D$ of $\mathrm{max.GB.deg}_{\prec}(F)$, we can estimate the complexity of the Gröbner basis computation for $F$ as that of computing the reduced row echelon form of the degree-$D$ Macaulay matrix $M_{\leq D}(F)$.
+Note that $D$ does not denote degree of regularity in this subsection.
+Here, we can take $M_{\leq D}(F)$ to be a block matrix of the form
+$M_{\leq D}(F)=\begin{pmatrix}M_{D}(F)&&\\ &\ddots&\\ &&M_{d_{0}}(F)\end{pmatrix}$
+with $d_{0}:=\min\{\deg(f_{j}):1\leq j\leq m\}$, where each $M_{d}(F)$ is a homogeneous Macaulay matrix defined as follows: For each $d$ with $d\geq d_{0}$, the degree-$d$ homogeneous Macaulay matrix $M_{d}(F)$ of $F$ has columns indexed by the terms of $R_{d}$ sorted, from left to right, according to the chosen order $\prec$.
+The rows of $M_{d}(F)$ are indexed by the polynomials $m_{i,j}f_{j}$, where $m_{i,j}\in R$ is a term such that $\deg(m_{i,j}f_{j})=d$.
+Theorem A.4.1.
+Let $F=\{f_{1},\ldots,f_{m}\}\subset R=K[x_{1},\ldots,x_{n}]$ be a set of non-constant homogeneous polynomials, and put $d_{j}=\deg(f_{j})\geq 1$ for each $j$ with $1\leq j\leq m$.
+Let $D$ be a non-negative integer such that $D=O(n)$.
+Then, the reduced row echelon form of $M_{\leq D}(F)$ is computed in
+$O\left(m\binom{n+D-1}{D}^{\omega}\right)$ (A.4.1)
+arithmetic operations in $K$.
+Hence, if $\prec$ is a graded order, and if $\mathrm{max.GB.deg}_{\prec}(F)=D=O(n)$, then the complexity of the Gröbner basis computation for $F$ is upper-bounded by (A.4.1).
+Proof.
+It suffices to estimate the complexity of computing the RREFs of the homogeneous Macaulay matrices $M_{D}(F),\ldots,M_{d_{0}}(F)$, where $d_{0}:=\min\{\deg(f_{i}):1\leq i\leq m\}\geq 1$.
+The number of rows (resp.
+columns) in each $M_{d}(F)$ is $\sum_{1\leq i\leq m,\ d\geq d_{i}}^{m}\binom{n-1+d-d_{i}}{d-d_{i}}$ (resp.
+$\binom{n-1+d}{d}$).
+It follows from $d-d_{i}\leq d-1$ that
+$\sum_{1\leq i\leq m,\ d\geq d_{i}}^{m}\binom{n-1+d-d_{i}}{d-d_{i}}\leq\sum_{i=1}^{m}\binom{n-1+d-1}{d-1}\leq m\binom{n-1+d-1}{d-1},$
+whence the number of rows in $M_{d}(F)$ is upper-bounded by $m\binom{n-1+d-1}{d-1}$. Therefore, the complexity of the row reductions on $M_{D}(F),\ldots,M_{d_{0}}(F)$ is upper-bounded by
+$\sum_{d=d_{0}}^{D}m\binom{n-1+d-1}{d-1}\binom{n-1+d}{d}^{\omega-1}$ $\leq m\sum_{d=1}^{D}\binom{n-1+d-1}{d-1}\binom{n-1+d}{d}^{\omega-1}$
+$\leq m\sum_{d=1}^{D}\binom{n-1+d}{d}^{\omega},$ (A.4.2)
+where we used a fact that the reduced row echelon form of a $k\times\ell$ matrix $A$ over a field $K$ can be computed in $O(k\ell^{\omega-1})$, see Remark A.4.3 below for its proof.
+Here, for each $d$ with $11\leq d\leq D-$ 1, we have
+$\binom{n-1+d}{d}=\frac{(n-1+d)!}{(n-1)!d!}=\binom{n-1+d+1}{d+1}\times\frac{d+1}{n+d}.$
+Repeating this procedure, we obtain
+$\binom{n-1+d+1}{d+1}\times\frac{d+1}{n+d}$ $=\binom{n-1+d+2}{d+2}\times\frac{d+2}{n+d+1}\times\frac{d+1}{n+d}$
+$=\cdots=\binom{n-1+D}{D}\times\frac{D}{n+D-1}\times\cdots\times\frac{d+2}{n+d+1}\times\frac{d+1}{n+d}.$
+It is clear that $\frac{d^{\prime}+1}{n+d^{\prime}}=1-\frac{n-1}{n+d^{\prime}}$ is monotonically increasing with respect to $d^{\prime}$. Putting $c:=(D-1)/n$, we have $D=cn+1$, so that
+$\frac{d^{\prime}+1}{n+d^{\prime}}\leq\frac{D}{n+D-1}=\frac{cn+1}{(c+1)n}=\frac{c+\frac{1}{n}}{c+1}$
+for any $d^{\prime}$ with $d^{\prime}\leq D-1$. As we consider $n\to\infty$, we may suppose that
+$\frac{d^{\prime}+1}{n+d^{\prime}}\leq\frac{c+\frac{1}{n}}{c+1}<\frac{c+\varepsilon}{c+1}<1$
+for some constant $\varepsilon$ with $10<\varepsilon<$ 1. Therefore, we obtain
+$\binom{n-1+D}{D}\times\frac{D}{n+D-1}\times\cdots\times\frac{d+2}{n+d+1}\times\frac{d+1}{n+d}<\binom{n-1+D}{D}\left(\frac{c+\varepsilon}{c+1}\right)^{D-d},$
+whence
+$\sum_{d=1}^{D}\binom{n-1+d}{d}^{\omega}$ $<\binom{n-1+D}{D}^{\omega}\sum_{k=0}^{D-1}\left(\frac{c+\varepsilon}{c+1}\right)^{k\omega}$
+$=\binom{n-1+D}{D}^{\omega}\times\frac{1-\left(\frac{c+\varepsilon}{c+1}\right)^{D\omega}}{1-\left(\frac{c+\varepsilon}{c+1}\right)^{\omega}}<\binom{n-1+D}{D}^{\omega}\times\frac{1}{1-\left(\frac{c+\varepsilon}{c+1}\right)^{\omega}},$
+where we put $k:=D-d$.
+It follows from $D=O(n)$ that $c$ is upper-bounded by a constant.
+By $10<\varepsilon<$ 1, it is clear that $\frac{c+\varepsilon}{c+1}$ is also upper-bounded by a positive constant smaller than $1$, whence
+$\frac{1}{1-\left(\frac{c+\varepsilon}{c+1}\right)^{\omega}}=O(1)$
+as $n\to\infty$. We have proved the assertion. ∎
+Corollary A.4.2.
+For an inhomogeneous $F=\{f_{1},\ldots,f_{m}\}\subset R$, and for a graded monomial ordering $\prec$ on $R$, we put $D_{\rm max}=\mathrm{max.GB.deg}_{\prec}(F^{h})$.
+If $D_{\rm max}=O(n)$, then the complexity of computing a Gröbner basis $G$ of $\langle F^{h}\rangle$ with respect to $\prec^{h}$ is
+$O\left(m\binom{n+D_{\rm max}}{D_{\rm max}}^{\omega}\right).$
+Hence, a Gröbner basis of $\langle F\rangle$ with respect to $\prec$ can be computed with the same complexity, if the complexity of substituting $y=1$ to $G$ is negligible.
+Remark A.4.3.
+In the proof of Theorem A.4.1, we used a fact that the reduced row echelon form of a $k\times\ell$ matrix $A$ over a field $K$ can be computed in $O(k\ell^{\omega-1})$, where $k$ can be greater than $\ell$.
+This fact is proved by considering the following procedures to compute the reduced row echelon form of $A$:
+  1. (0)
+    Write $k=\ell q+r$ for $q,r\in\mathbb{Z}$ with $0\leq r<\ell$.
+  2. (1)
+    Choose and remove $2\ell$ rows from $A$, and let $A^{\prime}$ be a $2\ell\times\ell$ matrix whose rows are the chosen $2\ell$ rows.
+  3. (2)
+    Compute the reduced row echelon form of $A^{\prime}$: Inserting $\ell$ zeros to the end of each row of $A^{\prime}$, construct a $2\ell\times 2\ell$ square matrix, and compute its reduced row echelon form.
+  4. (3)
+    As a result of (2), we obtain more than or equal to $\ell$ zero row vectors. We add the nonzero (linearly independent) row vectors obtained to $A$.
+  5. (4)
+    Go back to (1).
+Repeating the procedures from (1) to (4) at most $q$ times, we obtain the reduced row echelon form $B$ of the initial $A$.
+Since the complexity of (2) is $(2\ell)^{\omega}$, the total complexity to obtain $B$ is $q\cdot(2\ell)^{\omega}\leq(k/\ell)\cdot 2^{\omega}\ell^{\omega}=O(k\ell^{\omega-1})$, as desired.
+Remark A.4.4.
+In Corollary A.4.2, the assumption that $D_{\rm max}=O(n)$ holds in our setting in Subsection 4.3, when $d_{1},\ldots,d_{m}$ are fixed.
+Indeed, by the assumption that $\bm{F}^{h}$ is $\widetilde{d}(\langle F^{h}\rangle_{R^{\prime}})$-regular, recall that $D_{\rm max}$ is upper-bounded by the value of (4.3.1), which is maximized at $m=n+1$.
+The value at $m=n+1$ is equal to the Lazard’s bound $\sum_{j=1}^{n+1}(d_{j}-1)+1$ (cf.
+Remark 4.3.2), and this bound is $O(n)$ for fixed $d_{1},\ldots,d_{m}$.
+For example, Corollary A.4.2 can be applied to the MQ problem (cf.
+[44]), where all the total degrees of input polynomials are equal to two.
+If we do not assume $D=O(n)$ in Theorem A.4.1, a formula on the complexity is given by $O(mD\binom{n+D-1}{D}^{\omega})$, see [3, Proposition 1].
+This formula is obtained by simply upper-bounding the sum in (A.4.2) by $D\binom{n+D-1}{D}^{\omega}$.
+A.5 A proof of the complexity estimation given in Remark 1
+We here prove the the complexity estimation given in Remark 1, by a way similar to the proof of [42, Theorem 3.65] together with Theorem A.4.1.
+Let $F$ be a set of non-constant polynomials in $R=K[x_{1},\ldots,x_{n}]$, and assume that $F$ is inhomogeneous.
+As in Theorem 3, we also suppose that the sequence $\bm{F}=(f_{1},\ldots,f_{m})$ is cryptographic semi-regular, and let $\mathcal{A}$ be as in (4) of the theorem.
+By the cryptographic semi-regularity of $\bm{F}$, the procedures of $\mathcal{A}$ with input $F$ correspond completely to those with input $F^{h}$, until the step degree reaches $D:=d_{\rm reg}(\langle F^{\rm top}\rangle_{R})$.
+This means that the complexity of the procedures of $\mathcal{A}$ to obtain the intermediate set $\mathcal{H}:=\mathcal{G}^{(e)}\cup\mathcal{G}^{(e),D}$ given in Lemma 4.2.3 is estimated as that of computing the (reduced) $D$-Gröbner basis of $\langle F^{h}\rangle_{R^{\prime}}$.
+Therefore, it follows from Theorem A.4.1 that $\mathcal{H}$ is computed in $O(m\binom{n+D}{D}^{\omega})$.
+After this first half computation, we obtain a set $\mathcal{H}$ for generating $\langle F\rangle_{R}$ with $R_{D}=\langle\mathrm{LM}(\mathcal{H})\rangle_{D}$ such that the top part each element of $\mathcal{H}$ of total degree $D$ is a single term.
+Next, we bound the complexity of the latter half part of $\mathcal{A}$ as that of the classical Buchberger algorithm $\mathcal{B}$ (with Buchberger’s first criterion) for the input $\mathcal{H}$.
+During the execution of $\mathcal{B}$, the total degrees of the $S$-polynomials generated are upper-bounded by $22D-$ 2 (cf.
+Lemma 4.2.4), and those of the polynomials that are added to $\mathcal{H}$ are upper-bounded by $D-1$.
+In the following, we denote by $L(n,d)$ the number of monomials in $R_{\leq d}$ (cf.
+Tenti [42] uses $L_{q}(n,d)$ to denote the number of monomials in $(R/\langle x_{1}^{q},\ldots,x_{n}^{q}\rangle)_{\leq d}$.
+In the algorithm $\mathcal{B}$, we first initialize $\mathcal{H}^{\prime}$ and $B$ to be $\mathcal{H}$ and $\{\{f,g\}\mid f,g\in\mathcal{H}\mbox{ with $f\neq g$}\}$ respectively.
+Then, until $B$ becomes empty, we repeat the following procedures, where $i$ and $N_{i}$ with initial values $i=0$ and $N_{0}=0$ are used to count the number of loops:
+  1. 1.
+    Select and remove $\{f,g\}$ from $B$. Then compute the normal form $r$ of $S(f,g)$ with respect to $\mathcal{H}^{\prime}$, and replace $N_{i}$ by $N_{i}+1$.
+  2. 2.
+    If $r\neq 0$, add pairs of the form $(h,r)$ with $h\in\mathcal{H}^{\prime}$ to $B$, where we avoid to add unnecessary pairs by Buchberger’s first criterion.
+We also add $r$ to $\mathcal{H}$, and replace $i$ by $i+1$.
+Initialize $N_{i}:=0$, and then go back to the first step.
+We count the number of loops in the worst case.
+Let $k$ be the number of obtaining $r\neq 0$ through the loops; clearly $k\leq L(n,D-1)$.
+After we obtain such $r$’s $k$ times (namely $r\neq 0$ occurs $k$ times), the number of elements in $B$ decreases one by one for each loop, and finally it becomes zero (this means the loops end).
+Since, #$\#B\leq\frac{1}{2}L(n,D)^{2}$, we have $N_{i}\leq\frac{1}{2}L(n,D)^{2}$ for each $i$ with $0\leq i\leq k$, and hence the total number of loops is
+$N_{0}+N_{1}+\cdots+N_{k}\leq(k+1)\frac{1}{2}L(n,D)^{2}=O(L(n,D-1)L(n,D)^{2}).$
+If we avoid 0-reduction completely, we may assume $N_{i}=1$ for all $i$, so that the total number of loops is $O(L(n,D-1))$.
+We estimate the complexity of obtaining the normal form $r$ of each $S(f,g)$.
+Here $\mathcal{H}^{\prime}$ consists of polynomials of total degree $\leq D-1$ or ones of total degree $D$ whose top parts are single terms.
+By this together with $\deg S(f,g)\leq 2D-2$, the computation of the normal form $r$ is done in $O(L(n,2D-2)L(n,D-1))$.
+Therefore, the complexity of $\mathcal{B}$ with input $\mathcal{H}$ is
+$O(L(n,D-1)^{2}L(n,D)^{2}L(n,2D-2)),$
+which can be reduced to $O(L(n,D-1)^{2}L(n,2D-2))$ if we can avoid every 0-reduction completely.
+Considering this together with the complexity of the first half computation, we obtain the estimation in Remark 1.